### v3

# Semiformal method I. Pólya’s theorem for the complete graph $K_{\mathbb{N}}$ with complex edge weights

Abstract
Using the semiformal method in combinatorics we generalize Pólya’s theorem. This theorem determines the limiting probability that walks in the grid graph $\mathbb{Z}^{d}$ visit the given vertex. We generalize it to the countable complete graph $K_{\mathbb{N}}$ with edge weights in $\mathbb{C}$. In part II we treat edge weights in $\mathbb{C}[[x_{1},\dots,x_{k}]]$.
1 Introduction
Let $d,n\in\mathbb{N}$ ($\{1,2,\dots\}$) and $\overline{v}\in\mathbb{Z}^{d}$, where $\mathbb{Z}=\{\dots,-1,0,1,\dots\}$ are the integers. Let $P_{d}(\overline{v},n)$ be the probability that a walk of length $n$ in the grid graph $\mathbb{Z}^{d}$ starting at $\overline{0}$ visits in a step $i>0$ the vertex $\overline{v}$. In 1921, G. Pólya proved in [24] a theorem, which we generalize, asserting that for every vertex $\overline{v}$,
$\lim_{n\to\infty}P_{d}(\overline{v},\,n)\left\{\begin{array}[]{lll}=1&\dots&\text{if $d\leq 2$ and}\\ <1&\dots&\text{if $d\geq 3$}\,.\end{array}\right$.
Pólya’s theorem is almost always cited in the special case with $\overline{v}=\overline{0}$, although he proved it for any vertex $\overline{v}$ (see [16, Section 1] for a translated quote from [24]). A selection of articles and books discussing Pólya’s theorem is [3, 4, 6, 9, 11, 13, 18, 19, 20, 21, 25, 26, 29, 30].
We generalize it as follows. Let
$\mathbb{N}_{2}=\binom{\mathbb{N}}{2}=\{e\colon\;e\subset\mathbb{N}\wedge|e|=2\}$}
be edges of the countable complete graph $K_{\mathbb{N}}=\langle\mathbb{N},\mathbb{N}_{2}\rangle$ on the vertex set $\mathbb{N}$. For a finite set $A$ we denote by $|A|$ ($\in\mathbb{N}_{0}=\{0,1.\dots\}$) its cardinality. Let
$h\colon\mathbb{N}_{2}\to\mathbb{C}$
be a complex edge weight. $\mathbb{C}$ denotes the field of complex numbers, which is endowed with the usual Archimedean absolute value $|z|=\sqrt{z\cdot\overline{z}}$. For $n\in\mathbb{N}_{0}$ let $W(n)$ be the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at the vertex $1$. Thus $W(0)=\{\langle 1\rangle\}$ and for $n>0$ the set $W(n)$ is infinite and countable. For $v\in\mathbb{N}$ let $W(v,n)$ be the subset of $W(n)$ of the walks that visit the vertex $v$ in a step $i>0$. We extend the weight in the usual multiplicative way to
$h\colon\bigcup_{n=0}^{\infty}W(n)\to\mathbb{C}\,.$}
Let $W\subset\bigcup_{n=0}^{\infty}W(n)$. If the series
$\sum_{w\in W}h(w)$}
absolutely converges, we denote its sum by $h(W)$. (Results on series are reviewed in the next section.) Suppose that for every $n\in\mathbb{N}_{0}$ the series $\sum_{w\in W(n)}h(w)$ absolutely converges. Then so does $\sum_{w\in W(v,n)}h(w)$ and we consider the generating function
$A_{h,v}(x)=\sum_{n\geq 0}h(W(v,\,n))x^{n}\ \ (\in\mathbb{C}[[x]])\,.$}
Our generalizations of Pólya’s theorem are formulas for the quantities
$A_{h,v}(1)=\sum_{n\geq 0}h(W(v,\,n))\,\text{ and }\,\lim_{n\to\infty}h(W(v,\,n))\,.$}
The formulas appear in fourteen theorems: Theorems 3.23.7 in Section 3 and Theorems 4.24.9 in Section 4. Pólya’s theorem is subsumed in five of them. Theorems 3.5 and 3.7 extend cases $d\geq 3$ and $d\leq 2$, respectively, when $\overline{v}=\overline{0}$. Theorems 4.6 (4.7) and 4.9 extend cases $d\geq 3$ and $d\leq 2$, respectively, when $\overline{v}\neq\overline{0}$.
We have fourteen generalizations of one theorem for three reasons. First, in Pólya’s theorem weights are nonnegative real numbers and convergence of infinite series is automatically absolute. For complex weights we consider both kinds of convergence, absolute and conditional one. We also consider, for nonnegative real weights, series with sum $\infty$. Second, Pólya’s theorem lives in the probabilistic setup where weights of edges incident to a fixed vertex sum up to $1$. We consider the case when this condition is kept (convex weights) and the case when it is dropped. Third, already Pólya’s theorem is in reality two theorems because the cases $\overline{v}=\overline{0}$ and $\overline{v}\neq\overline{0}$ differ in complexity of proofs. In our generalization we distinguish cases $v=1$ and $v\neq 1$ as well. Now $3\cdot 2\cdot 2=12$; two more generalizations arise by variations of assumptions.
We explain how our approach includes Pólya’s original problem. If $f\colon A\to B$ is a map and $C$ is any set (not necessarily a subset of $A$), we call the set
$f[C]=\{f(x)\colon\;x\in A\cap C\}\ \ (\subset B)$
the image of $C$ by $f$. Let $d\in\mathbb{N}$ and $\overline{v}\in\mathbb{Z}^{d}$. We take any bijection $f\colon\mathbb{Z}^{d}\to\mathbb{N}$ such that $f(\overline{0})=1$, and set $v=f(\overline{v})$. The grid graph on $\mathbb{Z}^{d}$, which is the graph $\langle\mathbb{Z}^{d},E_{d}\rangle$ with the edges
$\{\overline{a},\,\overline{b}\}\in E_{d}\iff\sum_{i=1}^{d}|a_{i}-b_{i}|=1\,$,}
determines the weight $h\colon\mathbb{N}_{2}\to\{\frac{1}{2d},0\}$ by the relation $h(e)\neq 0$ iff $e=f[e^{\prime}]$ for some $e^{\prime}\in E_{d}$. Then for every $n\in\mathbb{N}_{0}$ we have
$P_{d}(\overline{v},\,n)=h(W(v,\,n))\,.$
We structure our article as follows. In Section 2 we review infinite series with complex summands and generating functions. Sections 3 and 4 contain our main results, the mentioned fourteen generalizations of Pólya’s theorem. Section 5 contains concluding remarks. We continue our investigation in [17] (see Section 5 for more details).
In Sections 3 and 4 we use what we term the semiformal method in (enumerative) combinatorics. It extends the symbolic method in combinatorics [2, 8, 10, 12, 22, 23, 27] from finite to countable sets. In our situation such extension is natural and inevitable, coefficients in generating functions $A_{h,v}(x)$ arise as sums of infinite series $\sum_{w\in W(v,n)}h(w)$. We develop the semiformal method more in [14]. Why “semiformal”? Generating “functions” in the symbolic method are formal power series with coefficients determined by finite operations. Our extension allows operations based on limit transitions. See [15] more applications of the semiformal method.
2 Series and generating functions in $\mathbb{C}$
We review series and generating functions in the domain $\mathbb{C}$. A series (in $\mathbb{C}$) is a map $h\colon X\to\mathbb{C}$, defined on an at most countable set $X$. We write it as $\sum_{x\in X}h(x)$. We say that $\sum_{x\in X}h(x)$ absolutely converges if two equivalent conditions hold.
  1. 1.
    For some $c>0$ and every finite set $Y\subset X$ we have $\sum_{x\in Y}|h(x)|\leq c$.
  2. 2.
    If $X$ is countable, then for every bijection $f\colon\mathbb{N}\to X$, the limit
    $s=\lim_{n\to\infty}{\textstyle\sum_{i=1}^{n}h(f(i))}$
    exists and does not depend on $f$.
We call $s$ the sum of the series and denote it again by $\sum_{x\in X}h(x)$. If the set $X=\{x_{1},x_{2},\dots,x_{n}\}\neq\emptyset$ is finite, then every series $\sum_{x\in X}h(x)$ absolutely converges and has the sum
$h(x_{1})+h(x_{2})+\dots+h(x_{n})\,.$
For $X=\emptyset$ we define the sum as 0. If
$U(x)=\sum_{n\geq 0}u_{n}x^{n}\ \ (\in\mathbb{C}[[x]])$}
is a formal power series, so that $x$ is a formal variable, we denote for any $x\in\mathbb{C}$ by $F_{U}(x)$ the sum of the power series $\sum_{n\geq 0}u_{n}x^{n}$, if it absolutely converges.
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$. A walk $w$ in $K_{\mathbb{N}}$ is any $(n+1)$-tuple $w$ of vertices $v_{i}\in\mathbb{N}$,
$w=\langle v_{0},\,v_{1},\,\dots,\,v_{n}\rangle\,$,
such that $n\in\mathbb{N}_{0}$ and $v_{i-1}\neq v_{i}$ for every $i\in[n]$ ($\{1,2,\dots,n\}$, $0]=\emptyset$). The length $n$ of $w$ is is $n$ ($\in\mathbb{N}_{0}$). We extend $h$ to walks. For $n>0$ we set
$h(w)=\prod_{i=1}^{n}h(\{v_{i-1},\,v_{i}\})\,$,}
and for $n=0$ we define $h(w)=1$. Recall that $W(n)$ is the set of walks $w$ in $K_{\mathbb{N}}$ with length $n$ such that $v_{0}=1$.
Definition 2.1 (lightness)
A weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is light if for every $n\in\mathbb{N}_{0}$ the series
$\sum_{w\in W(n)}h(w)$}
absolutely converges.
Let $S=\sum_{x\in X}h(x)$ be a series. If $Y\subset X$, we say that $\sum_{x\in Y}h(x)$ is a subseries of $S$. Any subseries of an absolutely convergent series absolutely converges, but we cannot compare sizes of their sums
$\big|\sum_{x\in Y}h(x)\big|\,\text{ and }\,\big|\sum_{x\in X}h(x)\big|\,.$}
Except when $h\colon X\to\mathbb{R}_{\geq 0}$: if $\sum_{x\in X}h(x)$ absolutely converges and $Y\subset X$, then the sums satisfy
$0\leq\sum_{x\in Y}h(x)\leq\sum_{x\in X}h(x)\,.$}
If $R=\sum_{x\in X}g(x)$ and $S=\sum_{x\in X}h(x)$ are series and $\alpha,\beta\in\mathbb{C}$, the series
$\alpha R+\beta S=\sum_{x\in X}(\alpha g(x)+\beta h(x))$}
is their linear combination. We omit the easy proof of the next result.
Proposition 2.2
If $R$ and $S$ absolutely converge and have respective sums $r$ and $s$, then series $\alpha R+\beta S$ absolutely converges and has the sum $\alpha r+\beta s$.
The next approximation lemma is a key tool.
Lemma 2.3
If $S=\sum_{x\in X}h(x)$ absolutely converges and has sum $s$, then for every $\varepsilon>0$ there exists a finite set $Y\subset X$, denoted by $Y(S,\,\varepsilon)$, such that for every finite set $Z$ with $Y\subset Z\subset X$ we have
$\big|s-\sum_{x\in Z}h(x)\big|\leq\varepsilon\,.$}
Proof. Let an $\varepsilon>0$ be given. For finite $X$ we set $Y=X$. For countable $X$ we take any bijection $f\colon\mathbb{N}\to X$, take an $N\in\mathbb{N}$ such that
$|\sum_{n=1}^{N}h(f(n))-s|\leq\frac{\varepsilon}{2}\,\text{ and }\,\sum_{n>N}|h(f(n))|\leq\frac{\varepsilon}{2}$}
and set Y=f[[N]]Y=f[\,[N]\,]. Then for every finite set $Z$ with $Y\subset Z\subset X$ we have
$|s-\sum_{x\in Z}h(x)|\leq|s-\sum_{x\in Y}h(x)|+\sum_{x\in Z\setminus Y}|h(x)|\leq\frac{\varepsilon}{2}+\frac{\varepsilon}{2}=\varepsilon\,.$}
$\Box$
Two more operations with series are grouping and product. A partition of a set $X$ is a set $P$ of nonempty and disjoint sets such that $\bigcup P=X$. If $X$ is at most countable, then so is $P$ and every set $Z\in P$. If $S=\sum_{x\in X}h(x)$ is a series and $P$ is a partition of $X$ such that for every set $Z\in P$ the subseries $\sum_{x\in Z}h(x)$ absolutely converges and has sum $s_{Z}$, then the series
$S_{P}=\sum_{Z\in P}s_{Z}$}
is called the grouping of $S$.
Proposition 2.4
If $S=\sum_{x\in X}h(x)$ absolutely converges and has sum $s$, and if $P$ is a partition of $X$, then $S_{P}$ is correctly defined, absolutely converges and has the same sum $s$.
Proof. For every $Z\in P$ the series $S_{Z}=\sum_{x\in Z}h(x)$ is a subseries of $S$ and absolutely converges. Thus the series $S_{P}$ is correctly defined. We show that it absolutely converges. Let
$c=\sup(\{\sum_{x\in Z}|h(x)|\colon\;\text{$Z\subset X$ and $Z$ is finite}\})\ \ (<+\infty)\,.$}
Let $\{Z_{1},Z_{2},\dots,Z_{n}\}\subset P$ be a finite set. We use Lemma 2.3 and take for $i\in[n]$ finite sets
$Z_{i}^{\prime}=Y(S_{Z_{i}},\,2^{-i})\ \ (\subset Z_{i})\,.$
We form the disjoint union $Z_{0}=Z_{1}^{\prime}\cup\dots\cup Z_{n}^{\prime}$ ($\subset X$). Then
${\textstyle\sum_{i=1}^{n}\big|\sum_{x\in Z_{i}}h(x)\big|}$ $\leq$ ${\textstyle\sum_{i=1}^{n}\big|\sum_{x\in Z_{i}}h(x)-\sum_{x\in Z_{i}^{\prime}}h(x)\big|\,+}$
$+$ ${\textstyle\sum_{x\in Z_{0}}|h(x)|\leq\sum_{i=1}^{n}2^{-i}+c\leq 1+c}$
and $S_{P}$ absolutely converges.
Let $t$ be the sum of $S_{P}$. We show that $|t-s|\leq\varepsilon$ for every $\varepsilon>0$; then $t=s$. Let an $\varepsilon>0$ be given. We use Lemma 2.3 and take finite sets
$X^{\prime}=Y(S,\,\frac{\varepsilon}{3})\ \ (\subset X)\,\text{ and }\,P^{\prime}=Y(S_{P},\,\frac{\varepsilon}{3})\ \ (\subset P)\,.$}
We take a finite set $\{Z_{1},\dots,Z_{n}\}\subset P$ such that
$P^{\prime}\subset\{Z_{1},\,\dots,\,Z_{n}\}\,\text{ and }\,X^{\prime}\subset\bigcup_{i=1}^{n}Z_{i}\,.$}
We again use Lemma 2.3 and take for $i\in[n]$ finite sets
$Z_{i}^{\prime}=Y(S_{Z_{i}},\,2^{-i}\cdot\frac{\varepsilon}{3})\ \ (\subset Z_{i})\,.$}
For $i\in[n]$ let
$Z_{i}^{\prime\prime}=Z_{i}^{\prime}\cup(X^{\prime}\cap Z_{i})\ \ (\subset Z_{i})\,.$
We form the disjoint union
$X_{0}=\bigcup_{i=1}^{n}Z_{i}^{\prime\prime}\ \ (\subset X)\,.$}
Then $X_{0}$ is finite and $X^{\prime}\subset X_{0}$. Also, $Z_{i}^{\prime}\subset Z_{i}^{\prime\prime}$ for every $i\in[n]$. We have
$|s-t|$ $\leq$ ${\textstyle|s-\sum_{x\in X_{0}}h(x)|+\sum_{i=1}^{n}|\sum_{x\in Z_{i}^{\prime\prime}}h(x)-\sum_{x\in Z_{i}}h(x)|\,+}$
$+$ ${\textstyle|\sum_{i=1}^{n}\sum_{x\in Z_{i}}h(x)-t|\leq\frac{\varepsilon}{3}+\sum_{i=1}^{n}2^{-i}\cdot\frac{\varepsilon}{3}+\frac{\varepsilon}{3}\leq\varepsilon\,.}$
$\Box$
Let $A$ and $B$ be sets. We denote their ordered pair by $\langle A,\,B\rangle$. If $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i=1,2$, are series, then the series
$S_{1}\cdot S_{2}=\sum_{\langle x,\,y\rangle\in X_{1}\times X_{2}}h_{1}(x)h_{2}(y)$}
is the product of $S_{1}$ and $S_{2}$.
Proposition 2.5
If $S_{1}$ and $S_{2}$ absolutely converge and have sums $s_{1}$ and $s_{2}$, respectively, then $S_{1}\cdot S_{2}$ absolutely converges and has the sum $s_{1}s_{2}$.
Proof. We show that $S_{1}\cdot S_{2}$ absolutely converges. Let $c\geq 0$ be such that for any finite sets $Y_{i}\subset X_{i}$, $i=1,2$, we have $\sum_{x\in Y_{i}}|h_{i}(x)|\leq c$. Let $A\subset X_{1}\times X_{2}$ be a finite set. We take any finite sets $Y_{i}\subset X_{i}$ such that $A\subset Y_{1}\times Y_{2}$ and get the desired bound:
$\sum_{\langle x,\,y\rangle\in A}|h_{1}(x)h_{2}(y)|\leq\sum_{x\in Y_{1}}|h_{1}(x)|\cdot\sum_{y\in Y_{2}}|h_{2}(y)|\leq c\cdot c=c^{2}\,.$}
We show that the sum $s$ of $S_{1}\cdot S_{1}$ equals $s_{1}s_{2}$. It suffices to prove that for any $\varepsilon\in(0,1)$ we have $|s-s_{1}s_{2}|\leq\varepsilon$. So let $\varepsilon\in(0,1)$. Using Lemma 2.3 we take finite sets
$X^{\prime}_{1}=Y(S_{1},\,\frac{\varepsilon}{4(1+|s_{2}|)})\ \ (\subset X_{1}),\ \ X^{\prime}_{2}=Y(S_{2},\,\frac{\varepsilon}{4(1+|s_{1}|)})\ \ (\subset X_{2})\$}
and
$Z=Y(S_{1}\cdot S_{2},\frac{\varepsilon}{4})\ \ (\subset X_{1}\times X_{2})\,.$}
We take finite sets $X_{1}^{\prime\prime}$ and $X_{2}^{\prime\prime}$ such that $X_{1}^{\prime}\subset X_{1}^{\prime\prime}\subset X_{1}$, $X_{2}^{\prime}\subset X_{2}^{\prime\prime}\subset X_{2}$ and $Z\subset X_{1}^{\prime\prime}\times X_{2}^{\prime\prime}$. Then $|s-s_{1}s_{2}|$ is at most
${\textstyle\big|s-\sum_{\langle x,\,y\rangle\in X_{1}^{\prime\prime}\times X_{2}^{\prime\prime}}h_{1}(x)h_{2}(y)\big|+\big|\sum_{x\in X_{1}^{\prime\prime}}h_{1}(x)\sum_{y\in X_{2}^{\prime\prime}}h_{2}(y)-s_{1}s_{2}\big|}$
${\textstyle\leq\frac{\varepsilon}{4}+|(s_{1}+t_{1})(s_{2}+t_{2})-s_{1}s_{2}|\,,}$
where $|t_{1}|\leq\frac{\varepsilon}{4(1+|s_{2}|)}$ and $|t_{2}|\leq\frac{\varepsilon}{4(1+|s_{1}|)}$. Hence $|s-s_{1}s_{2}|\leq\frac{\varepsilon}{4}+|s_{1}t_{2}|+|s_{2}t_{1}|+|t_{1}t_{2}|\leq\frac{\varepsilon}{4}+\frac{\varepsilon}{4}+\frac{\varepsilon}{4}+\frac{\varepsilon}{4}=\varepsilon$. $\Box$
Let $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i\in[k]$, be $k\in\mathbb{N}$ series. We define more generally their product as the series
$\prod_{i=1}^{k}S_{i}=\sum_{\overline{x}\in\overline{X}}h_{1}(x_{1})h_{2}(x_{2})\dots h_{k}(x_{k})\,$,}
where $\overline{x}=\langle x_{1},\dots,x_{k}\rangle$ is an ordered $k$-tuple of $x_{1}$, $\dots$, $x_{k}$ and $\overline{X}=X_{1}\times\dots\times X_{k}$. Iterating Proposition 2.5 we generalize it as follows.
Proposition 2.6
If the series $S_{1}$, $\dots$, $S_{k}$ absolutely converge and have sums $s_{1}$, $\dots$, $s_{k}$, respectively, then $\prod_{i=1}^{k}S_{i}$ absolutely converges and has sum $s_{1}s_{2}\dots s_{k}$.
Interestingly, a reversal of Proposition 2.5 holds.
Corollary 2.7
If nonempty series $S_{1}$ and $S_{2}$ are not identically zero and if their product $S_{1}\cdot S_{2}$ absolutely converges and has the sum $s$, then $S_{1}$ and $S_{2}$ absolutely converge and their respective sums $s_{1}$ and $s_{2}$ satisfy $s_{1}s_{2}=s$.
Proof. This follows from Proposition 2.5 if we show that $S_{1}$ and $S_{2}$ absolutely converge. Let $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i=1,2$, and let $c>0$ be such that for every finite set $X\subset X_{1}\times X_{2}$ we have
$\sum_{\langle x,y\rangle\in X}|h_{1}(x)h_{2}(y)|\leq c\,.$}
We prove absolute convergence of $S_{1}$. We take any $y\in X_{2}$ with $h_{2}(y)\neq 0$. Then for every finite set $X\subset X_{1}$ we have
$\sum_{x\in X}|h_{1}(x)|=\frac{1}{|h_{2}(y)|}\sum_{\langle x,y\rangle\in X\times\{y\}}|h_{1}(x)h_{2}(y)|\leq\frac{c}{|h_{2}(y)|}\,.$}
Thus $S_{1}$ absolutely converges. For $S_{2}$ we argue similarly. $\Box$
It is clear that the assumptions on $S_{1}$ and $S_{2}$ cannot be omitted.
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and $n\in\mathbb{N}_{0}$. We obtain a sufficient condition for the lightness of the weight $h$ in terms of its values. We denote by $V(h,n)$ the set of vertices in $K_{\mathbb{N}}$ reachable from $1$ by a walk of length at most $n$ and with nonzero weight.
Definition 2.8 (slimness)
A weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is slim if for every $n\in\mathbb{N}_{0}$ there is a constant $c=c(n)>0$ such that for every vertex $u\in V(h,n)$ and every finite set $X\subset\mathbb{N}\setminus\{u\}$ we have
$\sum_{v\in X}|h(\{u,\,v\})|\leq c\,.$}
Proposition 2.9
Every slim weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is light.
Proof. Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be slim. We show by induction on $n\in\mathbb{N}_{0}$ that the series
$\sum_{w\in W(n)}h(w)$}
absolutely converges. For $n=0$ it is trivial because $W(0)=\{\langle 1\rangle\}$. Let $n>0$ and let $c>0$ be a constant such that for every finite set $X\subset W(n-1)$ we have
$\sum_{w\in X}|h(w)|\leq c$}
and that for every $u\in V(h,n-1)$ and every finite set $X\subset\mathbb{N}\setminus\{u\}$ we have
$\sum_{v\in X}|h(\{u,\,v\})|\leq c\,.$}
Let $X\subset W(n)$ be finite. We may assume that walks in $X$ have only edges with nonzero weights. We decompose every walk $w\in X$ as $w=w^{\prime}\,\ell(w)$, where $w^{\prime}\in W(n-1)$ and $\ell(w)$ is the last vertex. Let $Y$ be the set of walks $w^{\prime}$. For $w^{\prime}\in Y$ let
$X(w^{\prime})=\{v\in\mathbb{N}\setminus\{\ell(w^{\prime})\}\colon\;w^{\prime}v\in X\}\,.$
We get the bound
${\textstyle\sum_{w\in X}|h(w)|}$ $=$ ${\textstyle\sum_{w^{\prime}\in Y}|h(w^{\prime})|\cdot\sum_{v\in X(w^{\prime})}|h(\{\ell(w^{\prime}),\,v\})|}$
$\leq$ ${\textstyle\sum_{w^{\prime}\in Y}|h(w^{\prime})|\cdot c\leq c^{2}\,.}$
$\Box$
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$. We define $V(h)=\bigcup_{n\geq 0}V(h,n)$. It is the set of vertices in $K_{\mathbb{N}}$ reachable from $1$ by walks with nonzero weights. We define the graph
$G(h)=\langle V(h),\,E(h)\rangle$
as the connected component of $1$ in the subgraph of $K_{\mathbb{N}}$ formed by edges with nonzero weight. It follows that if $G(h)$ has finite degrees then $h$ is slim and hence light.
Definition 2.10 (convexity)
A light weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is convex if for every vertex $u\in V(h)$ the sum
$\sum_{e\in E(h),\,e\ni u}h(e)=\sum_{v\in\mathbb{N}\setminus\{u\}}h(\{u,\,v\})=1\,.$}
We show that these sums exist.
Proposition 2.11
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be light and $u\in V(h)$. Then the series
$\sum_{v\in\mathbb{N}\setminus\{u\}}h(\{u,v\})$}
absolutely converges.
Proof. There is a walk $w_{0}=\langle u_{0},\dots,u_{n}\rangle$ in $G(h)$ such that $u_{0}=1$ and $u_{n}=u$. Let $c=h(w_{0})$ ($\neq 0$) and $W$ be the set of walks $\langle u_{0},\dots,u_{n},u_{n+1}\rangle$ in $K_{\mathbb{N}}$. Thus vertices $u_{0}$, $\dots$, $u_{n}$ are fixed and $u_{n+1}$ runs in $\mathbb{N}\setminus\{u_{n}\}=\mathbb{N}\setminus\{u\}$. The series $\sum_{w\in W}h(w)$ is a subseries of $\sum_{w\in W(n+1)}h(w)$ and therefore absolutely converges. By Proposition 2.2,
$\sum_{v\in\mathbb{N}\setminus\{u\}}h(\{u,v\})=\frac{1}{c}\sum_{w\in W}h(w)$}
absolutely converges. $\Box$
We survey some results on generating functions. We work with univariate formal power series
[…]
They form an integral domain denoted by $\mathbb{C}[[x]]$. For $U(x)$ and $V(x)$ in $\mathbb{C}[[x]]$,
$U(x)+V(x)=\sum_{n\geq 0}(u_{n}+v_{n})x^{n},\ U(x)\cdot V(x)=\sum_{n\geq 0}\big(\sum_{j=0}^{n}u_{j}v_{n-j}\big)x^{n}\,.$}
Neutral elements are $0=0+0x+0x^{2}+\dots$ and $1=1+0x+0x^{2}+\cdots$. Units in $\mathbb{C}[[x]]$ are exactly the formal power series $\sum_{n\geq 0}a_{n}x^{n}$ with $a_{0}\neq 0$.
Definition 2.12 ($U(1)_{1}$ and $U(1)_{2}$)
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$ be in $\mathbb{C}[[x]]$. We introduce the following terminology and objects.
  1. 1.
    If the limit
    $U(1)_{1}=\lim_{n\to\infty}{\textstyle\sum_{j=0}^{n}u_{j}\ \ (\in\mathbb{C})}$
    exists, we say that $U(1)_{1}$ exists.
  2. 2.
    If the series
    $\sum_{n\geq 0}u_{n}1^{n}=\sum_{n\geq 0}u_{n}$}
    absolutely converges, we denote the sum by $U(1)_{2}$ and say that $U(1)_{2}$ exists.
If $U(1)_{2}$ exists, then $U(1)_{1}$ exists and $U(1)_{1}=U(1)_{2}$, but the opposite implication is, in general, not true. For example, if
$U(x)=\sum_{n\geq 1}(-1)^{n-1}\frac{1}{n}x^{n}$}
then $U(1)_{1}$ exists and $U(1)_{1}=\log 2$, but $U(1)_{2}$ does not exist. But if
$U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{R}_{\geq 0}[[x]]$}
and $U(1)_{1}$ exists, then $U(1)_{2}$ exists and $U(1)_{1}=U(1)_{2}$. The following result is well known.
Proposition 2.13
If $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{C}[[x]]$ and $U(1)_{1}$ exists, then the power series $\sum_{n\geq 0}u_{n}x^{n}$ absolutely converges for every $x\in\mathbb{C}$ with $|x|<1$.
In fact, for the absolute convergence it suffices (and is necessary) by Hadamard’s formula that $\limsup_{n\to\infty}|u_{n}|^{1/n}\leq 1$.
Definition 2.14 ($U(1)_{1}=+\infty$)
If $U(x)=\sum_{n\geq 0}u_{n}x^{n}$ is in $\mathbb{R}[[x]]$ and
$\lim_{n\to\infty}{\textstyle\sum_{j=0}^{n}u_{j}=+\infty\,,}$
we say that $U(1)_{1}=+\infty$.
Proposition 2.15
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$ and $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ be in $\mathbb{C}[[x]]$. If $V(1)_{1}$ exists and $U(x)=V(x)\frac{1}{1-x}$, then
$\lim_{n\to\infty}u_{n}=V(1)_{1}\,.$
Proof. Since $\frac{1}{1-x}=1+x+x^{2}+\dots$, we have $u_{n}=v_{0}+v_{1}+\dots+v_{n}$. $\Box$
Quantities $U(1)_{1}$ are handled by means of the classical theorem due to N. Abel, see [1] and [28, Chapter II.7]. We state it for the radius of convergence $R\geq 1$.
Theorem 2.16 (Abel’s 1)
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{C}[[x]]$. If $U(1)_{1}$ exists then the function
$F_{U}(x)=\sum_{n\geq 0}u_{n}x^{n}\colon\{x\in\mathbb{C}\colon\;|x|<1\}\to\mathbb{C}\,$,}
see Proposition 2.13, has the limit
$\lim_{\begin{subarray}{c}x\to 1\\ x\in[0,\,1)\end{subarray}}F_{U}(x)=U(1)_{1}\,.$
The following folklore result is in fact another variant of Abel’s theorem, incomparable with Theorem 2.16. We omit the easy proof.
Theorem 2.17 (Abel’s 2)
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{R}_{\geq 0}[[x]]$. If $U(x)$ absolutely converges for every $x\in[0,1)$, then the equality
$\lim_{x\to 1}F_{U}(x)=\lim_{n\to\infty}\sum_{j=0}^{n}u_{j}\ \ (\in[0,\,+\infty)\cup\{+\infty\})$}
always holds.
Compared to Theorem 2.16, the new case is when both limits equal $\infty$.
Proposition 2.18
Let $\alpha,\beta\in\mathbb{C}$ and $U(x),V(x),W(x)\in\mathbb{C}[[x]]$. If $U(x)=\alpha V(x)+\beta W(x)$ and if $V(1)_{1}$ and $W(1)_{1}$ exist, then $U(1)_{1}$ exists and
$U(1)_{1}=\alpha V(1)_{1}+\beta W(1)_{1}\,.$
Proof. Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$, $V(x)=\sum_{n\geq 0}u_{n}x^{n}$ and $W(x)=\sum_{n\geq 0}u_{n}x^{n}$. Then
$\sum_{j=0}^{n}u_{j}=\sum_{j=0}^{n}(\alpha v_{j}+\beta w_{j})=\alpha\sum_{j=0}^{n}v_{j}+\beta\sum_{j=0}^{n}w_{j}\,.$}
By the limit transition $n\to\infty$ we get that that $U(1)_{1}$ exists and $U(1)_{1}=\alpha V(1)_{1}+\beta W(1)_{1}$. $\Box$
Iterating this result we easily generalize it to $k$-term linear combinations.
Proposition 2.19
Let $U(x),V_{1}(x),\dots,V_{n}(x)\in\mathbb{C}[[x]]$, $n\in\mathbb{N}$. If $U(x)=\prod_{i=1}^{n}V_{i}(x)$ and $U(1)_{1}$, $V_{1}(1)_{1}$, $\dots$, $V_{n}(1)_{1}$ exist, then
$U(1)_{1}=\prod_{i=1}^{n}V_{i}(1)_{1}\,.$}
Proof. By Propositions 2.4, 2.6 and 2.13, and by the assumptions,
$F_{U}(x)=\prod_{i=1}^{n}F_{V_{i}}(x)$ for every $x\in\mathbb{C}$ with $|x|<1$
(see proofs of Propositions 2.25 and 2.27 for details of similar arguments). Theorem 2.16 and properties of limits of functions give
$U(1)_{1}$ $=$ $\lim_{\begin{subarray}{c}x\to 1\\ x\in[0,\,1)\end{subarray}}F_{U}(x)=\lim_{\begin{subarray}{c}x\to 1\\ x\in[0,\,1)\end{subarray}}{\textstyle\prod_{i=1}^{n}F_{V_{i}}(x)}$
$=$ ${\textstyle\prod_{i=1}^{n}\lim_{\begin{subarray}{c}x\to 1\\ x\in[0,\,1)\end{subarray}}F_{V_{i}}(x)=\prod_{i=1}^{n}V_{i}(1)_{1}}\,.$
$\Box$
Since we have to assume that $U(1)_{1}$ exists, the case $n>2$ does not follow by iterating the case $n=2$. The example in [7] shows that the assumption of existence of $U(1)_{1}$ cannot be omitted.
Proposition 2.20
Let $U(x),V(x)\in\mathbb{C}[[x]]$ and $V(0)\neq 0$. If $U(x)=\frac{1}{V(x)}$ and if $U(1)_{1}$ and $V(1)_{1}$ exist, then $V(1)_{1}\neq 0$ and
$U(1)_{1}=\frac{1}{V(1)_{1}}\,.$}
Proof. In $\mathbb{C}[[x]]$ we have the identity $U(x)V(x)=1$. Proposition 2.19 gives that $U(1)_{1}V(1)_{1}=1$. Thus $V(1)_{1}\neq 0$ and division by $V(1)_{1}$ yields the stated formula. $\Box$
Proposition 2.21
Let $U(x),V(x)\in\mathbb{R}_{\geq 0}[[x]]$ and $V(0)\neq 0$. If $\frac{1}{V(x)}, U(x)U(x)$ and $V(x)$ absolutely converge for $x\in[0,1)$, and if $V(1)_{1}=+\infty$, then
$U(1)_{1}=0\,.$
Proof. In $\mathbb{C}[[x]]$ we have the identity $U(x)V(x)=1$. By Propositions 2.4 and 2.5, $F_{U}(x)F_{V}(x)=1$ for $x\in[0,1)$. Thus $F_{V}(x)\neq 0$ and $F_{U}(x)=\frac{1}{F_{V}(x)}$ for every $x\in[0,\,1)$. Two applications of Theorem 2.17 and properties of limits of functions give
$U(1)_{1}$ $=$ ${\textstyle\lim_{x\to 1}F_{U}(x)=\lim_{x\to 1}\frac{1}{F_{V}(x)}=\frac{1}{\lim_{x\to 1}F_{V}(x)}}$
$=$ ${\textstyle\frac{1}{\lim_{n\to\infty\sum_{j=0}^{n}v_{j}}}=\frac{1}{+\infty}=0\,.}$
$\Box$
Proposition 2.22
Let $U(x),V(x),W(x)\in\mathbb{C}[[x]]$ and $W(0)\neq 0$. If $U(x)^{2}=1-\frac{V(x)}{W(x)}$, if $U(1)_{1}$, $V(1)_{1}$ and $W(1)_{1}$ exist, and if $W(1)_{1}\neq 0$, then
$\big(U(1)_{1}\big)^{2}=1-\frac{V(1)_{1}}{W(1)_{1}}\,.$}
Proof. In $\mathbb{C}[[x]]$ we have the identity $U(x)^{2}\cdot W(x)=W(x)-V(x)$. Let T(x)=W(x)V(x)T(x)=W(x)-V(x). By Proposition 2.18, $T(1)_{1}$ exists and $T(1)_{1}=W(1)_{1}-V(1)_{1}$. By Proposition 2.19 we have $(U(1)_{1})^{2}\cdot W(1)_{1}=T(1)_{1}=W(1)_{1}-V(1)_{1}$. Division by $W(1)_{1}$ yields the stated formula. $\Box$
Proposition 2.23
Let $U(x),V(x),W(x)\in\mathbb{R}_{\geq 0}[[x]]$ and $V(0),W(0)>0$. If $U(x)^{2}=1-\frac{V(x)}{W(x)}$, if $U(x)$, $V(x)$ and $W(x)$ absolutely converge for $x\in[0,1)$, if $F_{V}(x)$ is bounded on $[0,1)$ and if $W(1)_{1}=+\infty$, then $U(1)_{1}$ exists and
$U(1)_{1}=1\,.$
Proof. In $\mathbb{C}[[x]]$ we have the identity $U(x)^{2}\cdot W(x)=W(x)-V(x)$. Using the assumptions and Propositions 2.2, 2.4 and 2.6 we get that $F_{U}(x)^{2}\cdot F_{W}(x)=F_{W}(x)-F_{V}(x)$ for every $x\in[0,\,1)$. Since $F_{W}(0)>0$ and $F_{W}(x)$ is non-decreasing on $[0,1)$, we can divide by $F_{W}(x)$ and get that $F_{U}(x)^{2}=1-\frac{F_{V}(x)}{F_{W}(x)}$ for every $x\in[0,\,1)$. Note that $\lim_{x\to 1}F_{V}(x)=a>0$ because $F_{V}(0)>0$ and $F_{V}(x)$ is non-decreasing and bounded from above. Theorem 2.17 and properties of limits of functions give
$(U(1)_{1})^{2}$ $=$ ${\textstyle\big(\lim_{x\to 1}F_{U}(x)\big)^{2}=\lim_{x\to 1}F_{U}(x)^{2}=\lim_{x\to 1}\big(1-\frac{F_{V}(x)}{F_{W}(x)}\big)}$
$=$ ${\textstyle 1-\frac{\lim_{x\to 1}F_{V}(x)}{\lim_{x\to 1}F_{W}(x)}=1-\frac{a}{\lim_{n\to\infty}\sum_{j=0}^{n}w_{j}}=1-\frac{a}{+\infty}=1}\,.$
Hence $U(1)_{1}=1$ because $U(1)_{1}\geq 0$. $\Box$
We obtain analogs of the previous results in absolute convergence.
Proposition 2.24
Let $\alpha,\beta\in\mathbb{C}$ and $U(x),V(x),W(x)\in\mathbb{C}[[x]]$. If $U(x)=\alpha V(x)+\beta W(x)$ and if $V(1)_{2}$ and $W(1)_{2}$ exist, then $U(1)_{2}$ exists and
$U(1)_{2}=\alpha V(1)_{2}+\beta W(1)_{2}\,.$
Proof. Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$, $V(x)=\sum_{n\geq 0}u_{n}x^{n}$ and $W(x)=\sum_{n\geq 0}u_{n}x^{n}$. Since $V(1)_{2}$ and $W(1)_{2}$ are sums of absolutely convergent series $\sum_{n\in\mathbb{N}}v_{n}$ and $\sum_{n\in\mathbb{N}}w_{n}$ and $u_{n}=\alpha v_{n}+\beta w_{n}$, the result follows from Proposition 2.2. $\Box$
Proposition 2.25
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$, $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ and $W(x)=\sum_{n\geq 0}w_{n}x^{n}$ be in $\mathbb{C}[[x]]$. If $U(x)=V(x)W(x)$ and if $V(1)_{2}$ and $W(1)_{2}$ exist, then the following holds.
  1. 1.
    $U(1)_{2}$ exists and $U(1)_{2}=V(1)_{2}W(1)_{2}$.
  2. 2.
    $\sum_{n\geq 0}|u_{n}|\leq\sum_{n\geq 0}|v_{n}|\cdot\sum_{n\geq 0}|w_{n}|$.
Proof. 1. We consider the series
$S=\sum_{\langle j,\,k\rangle\in\mathbb{N}_{0}^{2}}v_{j}w_{k}=\sum_{j\in\mathbb{N}_{0}}v_{j}\cdot\sum_{k\in\mathbb{N}_{0}}w_{k}\,.$}
By Proposition 2.5 it absolutely converges and has sum $V(1)_{2}W(1)_{2}$. On the other hand, $S$ has the grouping $\sum_{n\geq 0}\sum_{j=0}^{n}v_{j}w_{n-j}$ (the inner $\sum$ means sum and the outer $\sum$ means series) and $\sum_{j=0}^{n}v_{j}w_{n-j}=u_{n}$. By Proposition 2.4, $S$ has the sum $u_{0}+u_{1}+\dots=U(1)_{2}$. Thus $U(1)_{2}=V(1)_{2}W(1)_{2}$.
2. This follows from the inequality ($n\in\mathbb{N}_{0}$)
$\sum_{m=0}^{n}|u_{m}|=\sum_{m=0}^{n}\big|\sum_{j=0}^{m}v_{j}w_{m-j}\big|\leq\sum_{j=0}^{n}|v_{j}|\cdot\sum_{j=0}^{n}|w_{j}|\,.$}
$\Box$
Proposition 2.19 for $n=2$ and item 1 of Proposition 2.25 are incomparable. The former assumes only conditional convergence, but for all three series. The latter assumes stronger absolute convergence, but only for two series. Another variant of the proposition is due to Franz (Franciszek) Mertens (1840–1927).
Theorem 2.26 (Mertens’s)
Suppose that $U(x)$, $V(x)$ and $W(x)$ are in $\mathbb{C}[[x]]$. If $U(x)=V(x)W(x)$, and if $V(1)_{1}$ and $W(1)_{2}$ exist, then $U(1)_{1}$ exists and
$U(1)_{1}=V(1)_{1}W(1)_{2}\,.$
See [7] for a proof.
Proposition 2.27
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$ and $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ be in $\mathbb{C}[[x]]$ and $v_{0}=1$. If $U(x)=\frac{1}{V(x)}$ and $c=|v_{1}|+|v_{2}|+\dots<1$, then the following holds.
  1. 1.
    $V(1)_{2}$ and $U(1)_{2}$ exist, $V(1)_{2}\neq 0$ and $U(1)_{2}=\frac{1}{V(1)_{2}}$.
  2. 2.
    $\sum_{n\geq 0}|u_{n}|\leq\frac{1}{1-c}$.
Proof. 1. It is clear that $V(1)_{2}$ exists. Since $|V(1)_{2}-1|\leq c<1$, we get $V(1)_{2}\neq 0$. We consider the series
$S=1+\sum_{\begin{subarray}{c}m\in\mathbb{N}\\ j_{m,1},\,j_{m,2},\,\dots,\,j_{m,m}\in\mathbb{N}\end{subarray}}(-1)^{m}v_{j_{m,1}}v_{j_{m,2}}\dots v_{j_{m,m}}\,.$}
We show that $S$ absolutely converges. It follows from the bound that the sum of absolute values of summands in any finite subseries of $S$ is at most $1+\sum_{m\geq 1}(|v_{1}|+|v_{2}|+\dots)^{m}=\frac{1}{1-c}$.
For $m\in\mathbb{N}$ we consider the series
$S_{m}=\prod_{i=1}^{m}(v_{1}+v_{2}+\dots)=\sum_{\langle j_{1},\,\dots,\,j_{m}\rangle\in\mathbb{N}^{m}}v_{j_{1}}v_{j_{2}}\dots v_{j_{m}}\,.$}
By Proposition 2.6 it absolutely converges and has sum $(V(1)_{2}-1)^{m}$. Thus by Propositions 2.2 and 2.4, $S$ has sum $1+\sum_{m\geq 1}(-1)^{m}(V(1)_{2}-1)^{m}=\frac{1}{V(1)_{2}}$.
On the other hand, in $\mathbb{C}[[x]]$ we have the identity
$U(x)=1+\sum_{m\geq 1}(-1)^{m}(v_{1}x+v_{2}x^{2}+\dots)^{m}\,.$}
Thus $u_{0}=1$ and for every $n\in\mathbb{N}$,
$u_{n}=\sum_{m\geq 1}(-1)^{m}\sum_{\begin{subarray}{c}j_{m,1},\,j_{m,2},\,\dots,\,j_{m,m}\in\mathbb{N}\\ j_{m,1}+j_{m,2}+\dots+j_{m,m}=n\end{subarray}}v_{j_{m,1}}v_{j_{m,2}}\dots v_{j_{m,m}}\,.$}
By Proposition 2.4, $S$ has also sum $u_{0}+u_{1}+u_{2}+\cdots=U(1)_{2}$ and the stated formula follows.
2. Indeed,
${\textstyle\sum_{k=0}^{n}|u_{k}|=}$
${\textstyle=1+\sum_{k=1}^{n}\big|\sum_{m\geq 1}(-1)^{m}\sum_{\begin{subarray}{c}j_{m,1},\,j_{m,2},\,\dots,\,j_{m,m}\in\mathbb{N}\\ j_{m,1}+j_{m,2}+\dots+j_{m,m}=k\end{subarray}}v_{j_{m,1}}v_{j_{m,2}}\dots v_{j_{m,m}}\big|}$
${\textstyle\leq\sum_{m\geq 0}(|v_{1}|+|v_{2}|+\dots)^{m}=\frac{1}{1-c}\,.}$
$\Box$
Again, Proposition 2.20 and item 1 of Proposition 2.27 are incomparable.
The analog of Proposition 2.22 for absolute convergence requires two lemmas.
Lemma 2.28
Let $V(x)\in\mathbb{C}[[x]]$ and $V(x)^{2}=tx^{2l}(1+t_{1}x+t_{2}x^{2}+\dots)$ where $l\in\mathbb{N}_{0}$ and $t,t_{j}\in\mathbb{C}$ (every square in $\mathbb{C}[[x]]$ has this form). Then there exists $v\in\mathbb{C}$ such that $v^{2}=t$ and
$V(x)=vx^{l}\sum_{n\geq 0}\binom{1/2}{n}(t_{1}x+t_{2}x^{2}+\dots)^{n}\,.$}
Proof. Let $n\in\mathbb{N}_{0}$ and a,ba,\,b be formal variables. By the Vandermonde identity,
$\sum_{j=0}^{n}\binom{a}{j}\binom{b}{n-j}=\binom{a+b}{n}\ \ (\in\mathbb{Q}[a,\,b])\,.$}
Let $V_{0}(x)=\sum_{n\geq 0}\binom{1/2}{n}(t_{1}x+t_{2}x^{2}+\dots)^{n}$. For any $v_{0}\in\mathbb{C}$ such that $(v_{0})^{2}=t$ the instance $a=b=\frac{1}{2}$ of the identity gives
$\big(v_{0}x^{l}V_{0}(x)\big)^{2}$ $=$ ${\textstyle tx^{2l}\sum_{n\geq 0}\big(\sum_{j=0}^{n}\binom{1/2}{j}\binom{1/2}{n-j}\big)(t_{1}x+t_{2}x^{2}+\dots)^{n}}$
$=$ ${\textstyle tx^{2l}\sum_{n\geq 0}\binom{1}{n}(t_{1}x+t_{2}x^{2}+\dots)^{n}}$
$=$ $tx^{2l}(1+t_{1}x+t_{2}x^{2}+\dots)=V(x)^{2}\,.$
Hence $V(x)=\pm v_{0}x^{l}V_{0}(x)$. $\Box$
Lemma 2.29
Let $U(x),V(x)\in\mathbb{C}[[x]]$ and $U(x)=V(x)^{2}=tx^{2l}(1+t_{1}x+t_{2}x^{2}+\dots)$ where $l\in\mathbb{N}_{0}$ and $t,t_{j}\in\mathbb{C}$. If $c=|t_{1}|+|t_{2}|+\dots<1$, then $U(1)_{2}$ and $V(1)_{2}$ exist, and
$U(1)_{2}=\big(V(1)_{2}\big)^{2}\,.$
Proof. Clearly, $U(1)_{2}$ exists. By Lemma 2.28 there is $v\in\mathbb{C}$ such that $v^{2}=t$ and
$V(x)=\sum_{n\geq 0}v_{n}x^{n}=vx^{l}\sum_{n\geq 0}\binom{1/2}{n}(t_{1}x+t_{2}x^{2}+\dots)^{n}\,.$}
Thus $v_{n}=0$ for $n<l$, $v_{l}=v$ and for $n>l$,
$v_{n}=v\sum_{m\geq 1}\binom{1/2}{m}\sum_{\begin{subarray}{c}j_{m,1},\,j_{m,2},\,\dots,\,j_{m,m}\in\mathbb{N}\\ j_{m,1}+j_{m,2}+\dots+j_{m,m}=n-l\end{subarray}}t_{j_{m,1}}t_{j_{m,2}}\dots t_{j_{m,m}}\,.$}
We consider the series
$S=0+0+\dots+0+v+\sum_{\begin{subarray}{c}m\in\mathbb{N}\\ j_{m,1},\,j_{m,2},\,\dots,\,j_{m,m}\in\mathbb{N}\end{subarray}}v\binom{1/2}{m}t_{j_{m,1}}t_{j_{m,2}}\dots t_{j_{m,m}}$}
with $l$ zero summands. We show that it absolutely converges. Let $T$ be any finite subseries of $S$. Indeed, since $\big|\binom{1/2}{m}\big|\leq 1$ for every $m\in\mathbb{N}$, the sum of absolute values of summands in $T$ is at most
$|v|+|v|\sum_{m\geq 1}(|t_{1}|+|t_{2}|+\dots)^{m}=\frac{|v|}{1-c}\,.$}
By Proposition 2.4, $S$ has the sum $v_{0}+v_{1}+\dots=V(1)_{2}$. By item 1 of Proposition 2.25, $U(1)_{2}=\big(V(1)_{2}\big)^{2}$. $\Box$
Proposition 2.30
Let $U(x),V(x)=\sum_{n\geq 0}v_{n}x^{n}$ and $W(x)=\sum_{n\geq 0}w_{n}x^{n}$ be in $\mathbb{C}[[x]]$. Suppose that $V(1)_{2}$ exists, $w_{0}=1$, $U(x)^{2}=1-\frac{V(x)}{W(x)}$ and $c=|w_{1}|+|w_{2}|+\dots<1$. If $V(x)=W(x)$, then $U(x)=0$. Else, if $k\in\mathbb{N}_{0}$ is minimum such that $v_{k}\neq w_{k}$ and
$\big(|w_{k}-v_{k}|\cdot(1-c)\big)^{-1}\sum_{n>k}|w_{n}-v_{n}|<1\,$,}
then $U(1)_{2}$ and $W(1)_{2}$ exist, $W(1)_{2}\neq 0$ and
$\big(U(1)_{2}\big)^{2}=1-\frac{V(1)_{2}}{W(1)_{2}}\,.$}
Proof. If $V(x)=W(x)$ then $U(x)^{2}=0$ and $U(x)=0$. Let $V(x)\neq W(x)$ and $k\in\mathbb{N}_{0}$ be as stated. We write
$U(x)^{2}=\frac{W(x)-V(x)}{W(x)}=tx^{2l}(1+t_{1}x+t_{2}x^{2}+\dots)$}
where $l\in\mathbb{N}_{0}$, $t=w_{k}-v_{k}\in\mathbb{C}^{*}$ and $t_{j}\in\mathbb{C}$. Let
$U_{0}(x)=\frac{1}{W(x)}=\sum_{n\geq 0}u_{0,n}x^{n},\ U_{1}(x)=W(x)-V(x)$}
and $U_{2}(x)=U_{0}(x)U_{1}(x)=\sum_{n\geq 0}u_{2,n}x^{n}$. By item 1 of Proposition 2.27, $U_{0}(1)_{2}$ and $W(1)_{2}$ exist, $W(1)_{2}\neq 0$ and $U_{0}(1)_{2}=\frac{1}{W(1)_{2}}$. By Proposition 2.24, $U_{1}(1)_{2}$ exists and $U_{1}(1)_{2}=W(1)_{2}-V(1)_{2}$. By item 1 of Proposition 2.25, $U_{2}(1)_{2}$ exists and $U_{2}(1)_{2}=U_{0}(1)_{2}\cdot U_{1}(1)_{2}$. By item 2 of Proposition 2.27, $\sum_{n\geq 0}|u_{0,n}|\leq\frac{1}{1-c}$. By item 2 of Proposition 2.25,
$\sum_{n\geq 0}|u_{2,n}|\leq\sum_{n\geq 0}|u_{0,n}|\cdot\sum_{n>k}|w_{n}-v_{n}|\leq\frac{1}{1-c}\sum_{n>k}|w_{n}-v_{n}|\,.$}
Since $U(x)^{2}=U_{2}(x)$, we have bound
$|t_{1}|+|t_{2}|+\dots\leq\frac{1}{|t|(1-c)}\sum_{n>k}|w_{n}-v_{n}|<1\,.$}
By Lemma 2.29, $U(1)_{2}$ exists and
$\big(U(1)_{2}\big)^{2}=U_{2}(1)_{2}=U_{0}(1)_{2}\cdot U_{1}(1)_{2}=1-\frac{V(1)_{2}}{W(1)_{2}}\,.$}
$\Box$
Finally, we consider convex light weights.
Proposition 2.31
For every convex light weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ we have
$\sum_{n=0}^{\infty}h(W(n))\cdot x^{n}=\frac{1}{1-x}$}\,.
Proof. Let $h$ be a convex light weight. We prove by induction on $n\in\mathbb{N}_{0}$ that $h(W(n))=1$. For $n=0$ this holds as $W(0)=\{\langle 1\rangle\}$. Let $n>0$. Recall that $\ell(w)$ denotes the last vertex of a walk $w$. Using Propositions 2.2 and 2.4, the convexity assumption and induction, we get
$h(W(n))$ $=$ ${\textstyle\sum_{w\in W(n)}h(w)}$
$=$ ${\textstyle\sum_{\begin{subarray}{c}w^{\prime}\in W(n-1)\\ \ell(w^{\prime})\in V(h)\end{subarray}}h(w^{\prime})\sum_{v\in\mathbb{N}\setminus\{\ell(w^{\prime})\}}h(\{\ell(w^{\prime}),\,v\})}$
$=$ ${\textstyle\sum_{w^{\prime}\in W(n-1)}h(w^{\prime})\cdot 1=1\,.}$
We can add and remove the condition $\ell(w^{\prime})\in V(h)$ without affecting the sum because walks $w^{\prime}$ not satisfying it have zero weight. $\Box$
3 Extensions of Pólya’s theorem with $v=1$
We establish the first group of generalizations of Pólya’s theorem in Theorems 3.23.7. We consider the case $v=1$ when the vertex visited by walks coincides with their starting vertex.
We introduce some generating functions. Let $n\in\mathbb{N}_{0}$, $v\in\mathbb{N}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. Recall that $W(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, and that $W(v,n)\subset W(n)$ is the set of walks that visit $v$ in step $i>0$. We are interested in the generating function
$A_{h}(x)=\sum_{n\geq 0}a_{n}^{h}x^{n}=\sum_{n\geq 0}h(W(1,\,n))x^{n}$}
and define
$B_{h}(x)$ $=$ ${\textstyle\sum_{n\geq 0}b_{n}^{h}x^{n}=\sum_{n\geq 0}h(W_{b}(n))x^{n}\,,}$
$C_{h}(x)$ $=$ ${\textstyle\sum_{n\geq 0}c_{n}^{h}x^{n}=\sum_{n\geq 0}h(W_{c}(n))x^{n}\,\text{ and}}$
$D_{h}(x)$ $=$ ${\textstyle\sum_{n\geq 0}d_{n}^{h}x^{n}=\sum_{n\geq 0}h(W(n))x^{n}\,,}$
all in $\mathbb{C}[[x]]$. Here $W_{b}(n)$ are the walks of length $n$ in $K_{\mathbb{N}}$ that start and end at $1$, and $W_{c}(n)$ are the walks of length $n$ in $K_{\mathbb{N}}$ that start and end at $1$, and avoid $1$ between. The coefficients $a_{n}^{h}$, $\dots$, $d_{n}^{h}$ are correctly defined because they are sums of subseries of absolutely convergents series. We have $a_{0}^{h}=c_{0}^{h}=0$ and $b_{0}^{h}=d_{0}^{h}=1$.
Proposition 3.1
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight and generating functions $A_{h}(x)$, $B_{h}(x)$, $C_{h}(x)$ and $D_{h}(x)$ be as above. The following relations hold between them.
  1. 1.
    $A_{h}(x)=C_{h}(x)D_{h}(x)$.
  2. 2.
    $B_{h}(x)=\frac{1}{1-C_{h}(x)}$, equivalently, $C_{h}(x)=1-\frac{1}{B_{h}(x)}$.
Proof. 1. We show that $a_{n}^{h}=\sum_{j=0}^{n}c_{j}^{h}\cdot d_{n-j}^{h}$ for every $n\in\mathbb{N}_{0}$. For $n\leq 1$ it holds, $a_{n}^{h}=c_{n}^{h}=0$. For $n\geq 2$ the set $W(1,n)$ is countable and we split every walk in it at the first revisit of $1$. We get a map
$F\colon W(1,\,n)\to\bigcup_{j=0}^{n}W_{c}(j)\times W(n-j)$}
that is defined as follows. For any walk
$w=\langle u_{0},\,u_{1},\,\dots,\,u_{j},\,\dots,\,u_{n}\rangle\in W(1,\,n)\,$,
where $u_{0}=u_{j}=1$, $j>0$ and $u_{i}\neq 1$ for $i\in[j-1]$, we set
$F(w)=\langle w_{1},\,w_{2}\rangle\ \ (\in W_{c}(j)\times W(n-j))\,$,
where $w_{1}=\langle u_{0},u_{1},\dots,u_{j}\rangle$ and $w_{2}=\langle u_{j},u_{j+1},\dots,u_{n}\rangle$. It is easy to see that the map $F$ is a bijection. $F$ is weight-preserving (WP): for every weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and value $F(w)=\langle w_{1},w_{2}\rangle$ we have $h(w)=h(w_{1})h(w_{2})$. If $X$ is a set of pairs of walks in $K_{\mathbb{N}}$, we write $h(X)$ for the sum of the series
$\sum_{\langle w,\,w^{\prime}\rangle\in X}h(w)h(w^{\prime})\,$,}
provided that it absolutely converges.
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight and $n\geq 2$. Using $F$ and above propositions, we get
$a_{n}^{h}$ $=$ $h(W(1,\,n))\stackrel{{\scriptstyle\mathrm{WP}}}{{=}}h(F[W(1,\,n)])$
[…]
2. We show that
$B_{h}(x)=\frac{1}{1-C_{h}(x)}=1+\sum_{j\geq 1}(C_{h}(x))^{j}\,.$}
We recall that $b_{0}^{h}=1$ and $c_{0}^{h}=0$ and show that for every $n\in\mathbb{N}$,
$b_{n}^{h}=\sum_{j=1}^{\infty}\sum_{\begin{subarray}{c}n_{1},\,\dots,\,n_{j}\in\mathbb{N}\\ n_{1}+\dots+n_{j}=n\end{subarray}}c_{n_{1}}^{h}c_{n_{2}}^{h}\dots c_{n_{j}}^{h}\,.$}
For odd $n$ it is true because then $b^{h}_{n}=c^{h}_{n}=0$ as $W_{b}(n)=W_{c}(n)=\emptyset$, and in every inner summand at least one index $n_{i}$ is odd. For even $n\geq 2$ the set $W_{b}(n)$ is countable. We split every walk in it at the visits of $1$ and get a map
$F\colon W_{b}(n)\to\bigcup_{j=1}^{\infty}\bigcup_{\begin{subarray}{c}n_{1},\,\dots,\,n_{j}\in\mathbb{N}\\ n_{1}+\dots+n_{j}=n\end{subarray}}W_{c}(n_{1})\times\cdots\times W_{c}(n_{j})$}
defined as follows. For any walk
$w=\langle u_{0},\,\dots,\,u_{m_{1}},\,\dots,\,\dots,\,u_{m_{j-1}},\,\dots,\,u_{n}\rangle\in W_{b}(n)\,$,
where $0=m_{0}<m_{1}<\dots<m_{j}=n$ and $u_{k}=1$ iff $k=0$ or $k=m_{i}$ for some $i\in[j]$, we set
$F(w)=\langle w_{1},\,w_{2},\,\dots,\,w_{j}\rangle\ \ (\in W_{c}(n_{1})\times W_{c}(n_{2})\times\dots\times W_{c}(n_{j}),\,n_{i}=m_{i}-m_{i-1})$
where $w_{i}=\langle u_{m_{i-1}},u_{m_{i-1}+1},\dots,u_{m_{i}}\rangle$, $i\in[j]$. It is easy to see that $F$ is a bijection and that it is weight-preserving similarly as above. If $X$ is a set of tuples of walks, we write $h(X)$ for the sum of the series
$\sum_{\langle w_{1},\,\dots,\,w_{k})\in X}h(w_{1})\dots h(w_{k})$}
if it absolutely converges.
Let $h$ be a light weight and $n\geq 2$ be even. Using $F$ and above propositions, we get
$b_{n}^{h}$ $=$ $h(W_{b}(n))\stackrel{{\scriptstyle\mathrm{WP}}}{{=}}h(F[W_{b}(n)])$
[…]
$\Box$
For better orientation we label the fourteen theorems by the triples
$\langle x,\,y,\,z\rangle\in\{v=1,\,v\neq 1\}\times\{\mathrm{gen},\,\mathrm{con}\}\times\{U(1)_{1},\,U(1)_{2},\,U(1)_{1}=+\infty\}\,$,
where “gen” refers to general edge weights $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and “con” to convex ones. The generating functions $A_{h}(x)$, $\dots$, $D_{h}(x)$ are as above.
Theorem 3.2 (1: $v=1$, gen, $U(1)_{1}$)
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. Suppose that $D_{h}(1)_{1}$, $C_{h}(1)_{1}$, $B_{h}(1)_{1}$ and $A_{h}(1)_{1}$ exist. Then $B_{h}(1)_{1}\neq 0$ and
$A_{h}(1)_{1}=\big(1-\frac{1}{B_{h}(1)_{1}}\big)D_{h}(1)_{1}\,.$}
Proof. By item 1 of Proposition 3.1 we have $A_{h}(x)=C_{h}(x)D_{h}(x)$. By Proposition 2.19, $A_{h}(1)_{1}=C_{h}(1)_{1}D_{h}(1)_{1}$. By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$. By Proposition 2.20 we have $B_{h}(1)_{1}\neq 0$ and $C_{h}(1)_{1}=1-\frac{1}{B_{h}(1)_{1}}$. The stated formula follows. $\Box$
Theorem 3.3 (2: $v=1$, gen, $U(1)_{2}$)
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. Suppose that $D_{h}(1)_{2}$ exists and $\sum_{n\geq 1}|b_{n}^{h}|<1$. Then $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $A_{h}(1)_{2}$ exists and
$A_{h}(1)_{2}=\big(1-\frac{1}{B_{h}(1)_{2}}\big)D_{h}(1)_{2}\,.$}
Proof. By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$. By item 1 of Proposition 2.27, $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $C_{h}(1)_{2}$ exists and $C_{h}(1)_{2}=1-\frac{1}{B_{h}(1)_{2}}$. By item 1 of Proposition 3.1 we have $A_{h}(x)=C_{h}(x)D_{h}(x)$. By item 1 of Proposition 2.25, $A_{h}(1)_{2}$ exists and $A_{h}(1)_{2}=C_{h}(1)_{2}D_{h}(1)_{2}$ . The stated formula follows. $\Box$
Theorem 3.4 (3: $v=1$, gen, $U(1)_{1}=+\infty$)
Suppose that $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$ is a light weight such that $D_{h}(1)_{1}=+\infty$. Then
$A_{h}(1)_{1}=+\infty\,.$
Proof. Thus $d_{n}^{h}>0$ for some $n>0$ and there is a vertex $u\in\mathbb{N}\setminus\{1\}$ with $h(\{1,u\})>0$. Then the walk $\langle 1,u,1\rangle$ shows that $c_{2}^{h}>0$. Since by item 1 of Proposition 3.1 we have $a_{n}^{h}=\sum_{j=0}^{n}c_{j}^{h}d_{n-j}^{h}$, we get $a_{n+2}^{h}\geq c_{2}^{h}d_{n}^{h}$ for every $n\in\mathbb{N}_{0}$. Hence
$\sum_{j=0}^{n}a_{j+2}^{h}\geq c_{2}^{h}\cdot\sum_{j=0}^{n}d_{j}^{h}\to+\infty,\ n\to\infty\,$,}
and $A_{h}(1)_{1}=+\infty$. $\Box$
We proceed to convex weights.
Theorem 3.5 (4: $v=1$, con, $U(1)_{1}$)
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a convex light weight. Suppose that $B_{h}(1)_{1}$ and $C_{h}(1)_{1}$ exist. Then $B_{h}(1)_{1}\neq 0$ and
$\lim_{n\to\infty}{\textstyle a_{n}^{h}=1-\frac{1}{B_{h}(1)_{1}}\,.}$
Proof. By item 1 of Proposition 3.1 and by Proposition 2.31 we have $A_{h}(x)=C_{h}(x)D_{h}(x)=C_{h}(x)\frac{1}{1-x}$. By Proposition 2.15, $\lim_{n\to\infty}a_{n}^{h}=C_{h}(1)_{1}$. By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$. By Proposition 2.20 we have $B_{h}(1)_{1}\neq 0$ and $C_{h}(1)_{1}=1-\frac{1}{B_{h}(1)_{1}}$. The stated formula follows. $\Box$
This theorem generalizes the case $d\geq 3$ of Pólya’s theorem with $\overline{v}=\overline{0}$. The case $d\leq 2$ is generalized in Theorem 3.7.
Theorem 3.6 (5: $v=1$, con, $U(1)_{2}$)
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a convex light weight. Suppose that $\sum_{n\geq 1}|b_{n}^{h}|<1$. Then $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$ and
$\lim_{n\to\infty}{\textstyle a_{n}^{h}=1-\frac{1}{B_{h}(1)_{2}}\,.}$
Proof. By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$. By item 1 of Proposition 2.27, $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $C_{h}(1)_{2}$ exists and $C_{h}(1)_{2}=1-\frac{1}{B_{h}(1)_{2}}$. By item 1 of Proposition 3.1 and by Proposition 2.31 we have $A_{h}(x)=C_{h}(x)D_{h}(x)=C_{h}(x)\frac{1}{1-x}$. By Proposition 2.15,
$\lim_{n\to\infty}a_{n}^{h}=C_{h}(1)_{1}=C_{h}(1)_{2}=1-\frac{1}{B_{h}(1)_{2}}\,.$}
$\Box$
Theorem 3.7 (6: $v=1$, con, $U(1)_{1}=+\infty$)
Let $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$ be a convex light weight such that $B_{h}(1)=+\infty$. Then
$\lim_{n\to\infty}a_{n}^{h}=1\,.$
Proof. By item 1 of Proposition 3.1 and Proposition 2.31 we have $A_{h}(x)=C_{h}(x)D_{h}(x)=C_{h}(x)\frac{1}{1-x}$. By Proposition 2.15, $\lim_{n\to\infty}a_{n}^{h}=C_{h}(1)_{1}$. By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$. But now by Proposition 2.31 we have $0\leq b_{n}^{h},c_{n}^{h}\leq d_{n}^{h}=1$ for every $n\in\mathbb{N}_{0}$. Hence $b_{n}^{h},c_{n}^{h}\in[0,1]$ and the generating functions $B_{h}(x)$ and $C_{h}(x)$ absolutely converge for $x\in[0,1)$. By Proposition 2.21, $C_{h}(1)_{1}=1$. The stated formula follows. $\Box$
4 Extensions of Pólya’s theorem with $v\neq 1$
The case when the vertex visited by walks differs from their starting vertex is more complicated. The grid graph $\mathbb{Z}^{d}$ is vertex-transitive, but so far, this has been irrelevant. For $\overline{v}\neq\overline{0}$ vertex-transitivity becomes important, and we generalize it to $K_{\mathbb{N}}$ as follows. Let $v\in\mathbb{N}\setminus\{1\}$ with $v\neq 1$. We say that a weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is $v$-transitive if there is a bijection $f\colon\mathbb{N}\to\mathbb{N}$ such that $f(1)=v$ and $h(f[e])=h(e)$ for every $e\in\mathbb{N}_{2}$.
We again introduce some generating functions. Let $n\in\mathbb{N}_{0}$, $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. Recall that $W(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, and that $W(v,n)\subset W(n)$ are the walks visiting the vertex $v$ ($\neq 1$); necessarily in a step $i>0$. We are interested in the generating function
$A_{h,v}(x)=\sum_{n\geq 0}a_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W(v,\,n))x^{n}\,.$}
The generating functions $B_{h}(x)$, $C_{h}(x)$ and $D_{h}(x)$ are as before. We define
$B_{h,v}(x)$ $=$ ${\textstyle\sum_{n\geq 0}b_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W_{b,v}(n))x^{n}\,,}$
$C_{h,v}(x)$ $=$ ${\textstyle\sum_{n\geq 0}c_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W_{c,v}(n))x^{n}\,\text{ and}}$
$E_{h,v}(x)$ $=$ ${\textstyle\sum_{n\geq 0}e_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W_{e,v}(n))x^{n}\,,}$
all in $\mathbb{C}[[x]]$. Here $W_{b,v}(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting and ending at $1$ and avoiding $v$. $W_{c,v}(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, ending at $v$ and between avoiding $v$. Finally, $W_{e,v}(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, ending at $v$ and between avoiding both $1$ and $v$. We have $a_{0}^{h,v}=c_{0}^{h,v}=c_{0}^{h}=e_{0}^{h,v}=0$ and $b_{0}^{h}=b_{0}^{h,v}=d_{0}^{h}=1$.
We obtain an analog of Proposition 3.1. The proof is similar. We omit details and argue only on the high level of semiformal method.
Proposition 4.1
Let $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight and generating functions $A_{h,v}(x)$, $B_{h}(x)$, $B_{h,v}(x)$, $C_{h,v}(x)$, $D_{h}(x)$ and $E_{h,v}(x)$ be as above. The following relations hold between them.
  1. 1.
    $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)$.
  2. 2.
    $B_{h}(x)=B_{h,v}(x)+C_{h,v}(x)^{2}B_{h}(x)$, equivalently, $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$.
  3. 3.
    $C_{h,v}(x)=B_{h,v}(x)E_{h,v}(x)$.
Proof. 1. Every walk $w$ weight-counted by $A_{h,v}(x)$ splits at the first visit of $v$ as $w=w_{1}w_{2}$. Walks $w_{1}$ are weight-counted by $C_{h,v}(x)$ and $w_{2}$ are arbitrary walks starting at $v$. By the $v$-transitivity of $h$, walks $w_{2}$ are weight-counted by $D_{h}(x)$.
2. It suffices to prove the first equality. $B_{h}(x)$ weight-counts walks $w$ starting and ending at $1$. Those avoiding $v$ are weight-counted by $B_{h,v}(x)$. If $w$ visits $v$, then it uniquely splits at the first and last visits of $v$ as $w=w_{1}w_{2}w_{3}$. Walks $w_{1}$ start at $1$, end at $v$ and between avoid $v$. Walks $w_{2}$ start and end at $v$. Walks $w_{3}$ start at $v$, end at $1$ and between avoid $v$. We reverse walks $w_{3}$ and see that both $w_{1}$ and $w_{3}$ are weight-counted by $C_{h,v}(x)$. Due to the $v$-transitivity of $h$, the middle walks $w_{2}$ are weight-counted by $B_{h}(x)$.
3. We consider walks $w$ weight-counted by $C_{h,v}(x)$. They start at $1$, end at $v$, and between avoid $v$. These walks uniquely split at the last visit of $1$ as $w=w_{1}w_{2}$. Walks $w_{1}$ are weight-counted by $B_{h,v}(x)$, and $w_{2}$ by $E_{h,v}(x)$. $\Box$
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and $v\in\mathbb{N}\setminus\{1\}$. If $v\not\in V(h)$ then $a_{n}^{h,v}=\sum_{w\in W(v,n)}h(w)=\sum_{w\in W(v,n)}0=0$ for every $n\in\mathbb{N}_{0}$ because every walk starting at $1$ and visiting $v$ contains an edge with zero weight. Henceforth, we therefore assume that $v\in V(h)$. For $w,z\in\mathbb{C}$ we define the set
$w\cdot\mathrm{sqrt}(z)=\{w\alpha\colon\;\alpha\in\mathbb{C},\,\alpha^{2}=z\}\ \ (\subset\mathbb{C})\,.$
For $w=0$ or $z=0$ it equals $\{0\}$. Else it has two non-zero elements differing by sign. If $w=1$, we write just $\mathrm{sqrt}(z)$.
Theorem 4.2 (7: $v\neq 1$, gen, $U(1)_{1}$)
Let $v\in\mathbb{N}\setminus\{1\}$ and let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight such that $v\in V(h)$. Suppose that $D_{h}(1)_{1}$, $C_{h,v}(1)_{1}$, $C_{h}(1)_{1}$, $B_{h,v}(1)_{1}$, $B_{h}(1)_{1}$ and $A_{h,v}(1)_{1}$ exist. Then $B_{h}(1)_{1}\neq 0$ and
$A_{h,\,v}(1)_{1}\in{\textstyle D_{h}(1)_{1}\cdot\mathrm{sqrt}\big(1-\frac{B_{h,\,v}(1)_{1}}{B_{h}(1)_{1}}}\big)\,.$
Proof. By item 1 of Proposition 4.1 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)$. By Proposition 2.19, $A_{h,v}(1)_{1}=C_{h,v}(1)_{1}D_{h}(1)_{1}$. By item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$. By item 2 of Proposition 3.1 and by Proposition 2.20, $B_{h}(1)_{1}\neq 0$. Thus by Proposition 2.22, $C_{h,v}(1)_{1}^{2}=1-\frac{B_{h,v}(1)_{1}}{B_{h}(1)_{1}}$. The stated formula follows. $\Box$
We assume the existence of $E_{h,v}(1)_{1}$ instead of $C_{h}(1)_{1}$ and get the following variant of Theorem 4.2.
Theorem 4.3 (7: $v\neq 1$, gen, $U(1)_{1}$)
Let $v\in\mathbb{N}\setminus\{1\}$ and let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight such that $v\in V(h)$. Suppose that $E_{h,v}(1)_{1}$, $D_{h}(1)_{1}$, $C_{h,v}(1)_{1}$, $B_{h,v}(1)_{1}$, $B_{h}(1)_{1}$ and $A_{h,v}(1)_{1}$ exist. Then
$A_{h,\,v}(1)_{1}\in{\textstyle D_{h}(1)_{1}\cdot\mathrm{sqrt}\big(1-\frac{B_{h,\,v}(1)_{1}}{B_{h}(1)_{1}}}\big)\,$,
where if $B_{h}(1)_{1}=0$ then $B_{h,v}(1)_{1}=0$ and we interpret $\frac{0}{0}$ as $1$, so that $A_{h,v}(1)_{1}=0$.
Proof. For $B_{h}(1)_{1}\neq 0$ we use the previous proof. If $B_{h}(1)_{1}=0$ then it follows from item 2 of Proposition 4.1 and Propositions 2.19 and 2.18 that also $B_{h,v}(1)_{1}=0$. Item 3 of Proposition 4.1 and Proposition 2.19 give $C_{h,v}(1)_{1}=0$. Finally, item 1 of Proposition 4.1 and Proposition 2.19 give $A_{h,v}(1)_{1}=0$. $\Box$
Theorem 4.4 (8: $v\neq 1$, gen, $U(1)_{2}$)
Let $v\in\mathbb{N}\setminus\{1\}$ and let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight such that $v\in V(h)$. Suppose that $D_{h}(1)_{2}$ and $B_{h,v}(1)_{2}$ exist, and that $c=\sum_{n\geq 1}|b_{n}^{h}|<1$. If $B_{h,v}(x)=B_{h}(x)$ then $A_{h,v}(x)=0$. Let $B_{h,v}(x)\neq B_{h}(x)$ and $k\in\mathbb{N}_{0}$ be minimum such that $b_{k}^{h,v}\neq b_{k}^{h}$. If
$\big(|b_{k}^{h,v}-b_{k}^{h}|\cdot(1-c)\big)^{-1}\sum_{n>k}|b_{n}^{h,v}-b_{n}^{h}|<1$}
then $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $A_{h,v}(1)_{2}$ exists and
$A_{h,\,v}(1)_{2}\in{\textstyle D_{h}(1)_{2}\cdot\mathrm{sqrt}\big(1-\frac{B_{h,\,v}(1)_{2}}{B_{h}(1)_{2}}}\big)\,.$
Proof. If $B_{h,v}(x)=B_{h}(x)$, by item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}=1-1=0$ and $C_{h,v}(x)=0$. Item 1 of Proposition 4.1 gives $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=0$.
Let $B_{h,v}(x)\neq B_{h}(x)$, $k$ be as stated and let the displayed bound hold. By item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$. By Proposition 2.30, $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $C_{h,v}(1)_{2}$ exists and $(C_{h,v}(1)_{2})^{2}=1-\frac{B_{h,v}(1)_{2}}{B_{h}(1)_{2}}$. By item 1 of Proposition 4.1 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)$. By item 1 of Proposition 2.25, $A_{h,v}(1)_{2}$ exists and $A_{h,v}(1)_{2}=C_{h,v}(1)_{2}D_{h}(1)_{2}$. The stated formula follows. $\Box$
Theorem 4.5 (9: $v\neq 1$, gen, $U(1)_{1}=+\infty$)
Let $v\in\mathbb{N}\setminus\{1\}$ and let $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$ be a $v$-transitive light weight such that $v\in V(h)$. If $D_{h}(1)_{1}=+\infty$ then
$A_{h,v}(1)_{1}=+\infty\,.$
Proof. Let $m\in\mathbb{N}$ be the minimum length of a walk in $G(h)$ joining $1$ and $v$. The inner vertices of this shortest walk differ from $v$ and we get that $c^{h,v}_{m}>0$. Since by item 1 of Proposition 4.1 we have $a_{n}^{h,v}=\sum_{j=0}^{n}c_{j}^{h,v}d_{n-j}^{h}$, we conclude that $a_{n+m}^{h,v}\geq c_{m}^{h,v}d_{n}^{h}$ for every $n\in\mathbb{N}_{0}$. Hence
$\sum_{j=0}^{n}a_{j+m}^{h,v}\geq c_{m}^{h,v}\cdot\sum_{j=0}^{n}d_{j}^{h}$}
for every $n\in\mathbb{N}_{0}$ and $A_{h,v}(1)_{1}=+\infty$. $\Box$
We proceed to convex weights.
Theorem 4.6 (10: $v\neq 1$, con, $U(1)_{1}$)
Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive convex light weight such that $v\in V(h)$. Suppose that $C_{h,v}(1)_{1}$, $C_{h}(1)_{1}$, $B_{h,v}(1)_{1}$ and $B_{h}(1)_{1}$ exist. Then $B_{h}(1)_{1}\neq 0$ and
$\lim_{n\to\infty}a_{n}^{h,v}\in{\textstyle\mathrm{sqrt}\big(1-\frac{B_{h,v}(1)_{1}}{B_{h}(1)_{1}}}\big)\,.$
Proof. By item 1 of Proposition 4.1 and by Proposition 2.31 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=C_{h,v}(x)\frac{1}{1-x}$. By Proposition 2.15, $\lim_{n\to\infty}a_{n}^{h,v}=C_{h,v}(1)_{1}$. By item 2 of Proposition 3.1 and by Proposition 2.20, $B_{h}(1)_{1}\neq 0$. By item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$. By Proposition 2.22, $C_{h,v}(1)_{1}^{2}=1-\frac{B_{h,v}(1)_{1}}{B_{h}(1)_{1}}$. The stated formula follows. $\Box$
We again assume the existence of $E_{h,v}(1)_{1}$ instead of $C_{h}(1)_{1}$ and get the following variant of the theorem.
Theorem 4.7 (10: $v\neq 1$, con, $U(1)_{1}$)
Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive convex light weight such that $v\in V(h)$. Suppose that $E_{h,v}(1)_{1}$, $C_{h,v}(1)_{1}$, $B_{h,v}(1)_{1}$ and $B_{h}(1)_{1}$ exist. Then
$\lim_{n\to\infty}a_{n}^{h,v}\in{\textstyle\mathrm{sqrt}\big(1-\frac{B_{h,v}(1)_{1}}{B_{h}(1)_{1}}}\big)\ \ (\subset\mathbb{C})\,$,
where if $B_{h}(1)_{1}=0$ then $B_{h,v}(1)_{1}=0$ and we interpret $\frac{0}{0}$ as $1$, so that the limit is zero.
Proof. If $B_{h}(1)_{1}\neq 0$, the previous proof works. If $B_{h}(1)_{1}=0$ then it follows from item 2 of Proposition 4.1 and Propositions 2.19 and 2.18 that also $B_{h,v}(1)_{1}=0$. Item 3 of Proposition 4.1 and Proposition 2.19 give $C_{h,v}(1)_{1}=0$. Thus, as in the previous proof, $\lim_{n\to\infty}a_{n}^{h,v}=C_{h,v}(1)_{1}=0$. $\Box$
These two theorems generalize the case $d\geq 3$ of Pólya’s theorem with $\overline{v}\neq\overline{0}$. The case $d\leq 2$ is generalized in Theorem 4.9.
Theorem 4.8 (11: $v\neq 1$, con, $U(1)_{2}$)
Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive convex light weight such that $v\in V(h)$. Suppose that $B_{h,v}(1)_{2}$ exists and $c=\sum_{n\geq 1}|b_{n}^{h}|<1$. If $B_{h,v}(x)=B_{h}(x)$ then $A_{h,v}(x)=0$. Let $B_{h,v}(x)\neq B_{h}(x)$ and $k\in\mathbb{N}_{0}$ be minimum such that $b_{k}^{h,v}\neq b_{k}^{h}$. If
$\big(|b_{k}^{h,v}-b_{k}^{h}|\cdot(1-c)\big)^{-1}\sum_{n>k}|b_{n}^{h,v}-b_{n}^{h}|<1$}
then $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$ and
$\lim_{n\to\infty}a_{n}^{h,v}\in{\textstyle\mathrm{sqrt}\big(1-\frac{B_{h,\,v}(1)_{2}}{B_{h}(1)_{2}}}\big)\,.$
Proof. If $B_{h,v}(x)\neq B_{h}(x)$, by item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}=1-1=0$ and $C_{h,v}(x)=0$. Item 1 of Proposition 4.1 gives $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=0$.
Let $B_{h,v}(x)\neq B_{h}(x)$, $k$ be as stated and let the displayed bound hold. By item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$. By Proposition 2.30, $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $C_{h,v}(1)_{2}$ exists and $C_{h,v}(1)_{2}^{2}=1-\frac{B_{h,v}(1)_{2}}{B_{h}(1)_{2}}$. By item 1 of Proposition 4.1 and by Proposition 2.31 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=C_{h,v}(x)\frac{1}{1-x}$. Proposition 2.15 gives the stated formula:
$\lim_{n\to\infty}a_{n}^{h,v}={\textstyle C_{h,\,v}(1)_{1}=C_{h,\,v}(1)_{2}\in\mathrm{sqrt}\big(1-\frac{B_{h,v}(1)_{2}}{B_{h}(1)_{2}}\big)\,.}$
$\Box$
Theorem 4.9 (12: $v\neq 1$, con, $U(1)_{1}=+\infty$)
Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$ be a $v$-transitive convex light weight such that $v\in V(h)$. If $B_{h}(1)_{1}=+\infty$ then
$\lim_{n\to\infty}a_{n}^{h,v}=1\,.$
Proof. By Proposition 2.31, $D_{h}(x)=\frac{1}{1-x}$. Hence $0\leq b_{n}^{h},b_{n}^{h,v},c_{n}^{h,v},e_{n}^{h,v}\leq d_{n}^{h}=1$ for every $n\in\mathbb{N}_{0}$ and we can use continuous and non-decreasing functions
$F_{B_{h}},\,F_{B_{h,v}},\,F_{C_{h,v}},\,F_{E_{h,v}}\colon[0,\,1)\to[0,\,+\infty)\,.$
Any shortest walk in $G(h)$ joining $1$ and $v$ shows that for some $m\in\mathbb{N}$ we have $e_{m}^{h,v}>0$. Thus the function $F_{E_{h,v}}(x)$ is on the interval $(0,1)$ positive and increasing. We deduce from it that $F_{B_{h,v}}(x)$ is bounded on $[0,1)$.
Suppose for the contrary that $\lim_{x\to 1}F_{B_{h,v}}(x)=+\infty$. By item 3 of Proposition 4.1 and by Propositions 2.4 and 2.5 we have for every $x\in[0,1)$ the equality $F_{C_{h,v}}(x)=F_{B_{h,v}}(x)F_{E_{h,v}}(x)$. Thus $\lim_{x\to 1}F_{C_{h,v}}(x)=+\infty$. Since $F_{C_{h,v}}(0)=0$, $F_{C_{h,v}}(x)$ increases and is continuous and $\lim_{x\to 1}F_{C_{h,v}}(x)=+\infty$, we obtain a (unique) number $x_{0}\in(0,1)$ such that $F_{C_{h,v}}(x)<1$ for $x\in[0,x_{0})$ and $\lim_{x\to x_{0}^{-}}F_{C_{h,v}}(x)=1$. By item 2 of Proposition 4.1 and by Propositions 2.4 and 2.5 we have for every $x\in[0,1)$ the equality
$F_{B_{h}}(x)=F_{B_{h,v}}(x)+F_{C_{h,v}}(x)^{2}F_{B_{h}}(x)\,\,.$
Hence for every $x\in[0,x_{0})$ we have $F_{B_{h}}(x)=\frac{F_{B_{h,v}}(x)}{1-F_{C_{h,v}}(x)^{2}}$. Since $F_{B_{h,v}}(0)=1$, it follows that $\lim_{x\to x_{0}^{-}}F_{B_{h}}(x)=+\infty$. But this contradicts the continuity of $F_{B_{h}}(x)$ at $x_{0}$.
Using item 2 of Proposition 4.1, the boundedness of $F_{B_{h,v}}(x)$ and Proposition 2.23, we get $C_{h,v}(1)_{1}=1$. By item 1 of Proposition 4.1 and by Proposition 2.31 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=C_{h,v}(x)\frac{1}{1-x}$. By Proposition 2.15, $\lim_{n\to\infty}a_{n}^{h,v}=C_{h,v}(1)_{1}=1$. $\Box$
5 Concluding remarks
In [17] we consider the same problem of determination of formulas for the quantities
$A_{h,v}(1)=\sum_{n\geq 0}h(W(v,\,n))\,\text{ and }\,\lim_{n\to\infty}h(W(v,\,n))$}
for formal weights $h$. These are maps
$h\colon\mathbb{N}_{2}\to\mathbb{C}[[x_{1},\,x_{2},\,\dots,\,x_{k}]]\,$,
where the last integral domain of formal power series in $k$ variables is endowed with the usual non-Archimedean norm.
The main novelty of our generalization is this. In an equality like
$P_{d}(\overline{v},\,n)=h(W(v,\,n))$
in the Introduction, the left-hand side involves, in the standard rigorous probabilistic treatment, an implicit probability space supporting the corresponding Markov chain. Construction of this probability space is far from trivial; see, for example, [6]. In our approach, the right-hand side is completely self-contained.