### v6

# Semiformal method I. Pólya’s theorem for the complete graph $K_{\mathbb{N}}$ with complex edge weights

Abstract
Using the semiformal method in combinatorics we generalize Pólya’s theorem. This theorem determines limits of probabilities that walks in the grid graph $\mathbb{Z}^{d}$ starting at the origin visit the given vertex. We generalize it to the countable complete graph $K_{\mathbb{N}}$ with edge weights in $\mathbb{C}$. In part II we consider analogs of Pólya’s theorem for edge weights in $\mathbb{C}[[x_{1},\dots,x_{k}]]$.
Besides the abstract, we want to help the reader understand our article by the following preamble. In it we have more space to explain what we do.
We mention Pólya’s theorem, which is a well known result in the discipline of random walks. However, our article is not really about generalizing Pólya’s theorem. We use extensions of Pólya’s theorem to edge weighted countable complete graphs merely as a test case for the novel generalization of the well known ([3, 8, 10, 12, 22, 23, 28]) symbolic method in enumerative (analytic) combinatorics. This generalization is due to the first author, and we call it the semiformal method. It will be developed in detail in [14]. The present article is an explanation of the functioning of the semiformal method in the form of an extensive example. More applications of it are given in [15]. The idea of the semiformal method is to extend finite sums defining coefficients of generating functions in the symbolic method to infinite (countable) series. Finite sums in a ring (in the symbolic method, one also needs to multiply) are straightforward, but it is less clear what is a sum of an infinite series of real or complex numbers. We use three kinds of sums of infinite series: absolute convergence (which is used in probability where summands are nonnegative), the more general and weaker conditional convergence, and sums equal to $\pm\infty$. The last sums are in fact important — the heart of the combinatorial proof of the case $d\leq 2$ in Pólya’s theorem is the fact that a certain series has sum $\infty$. For details on the uses of these three kinds of sums we refer to the article below. But why confine ourselves to infinite series of real or complex numbers, when non-Archimedean infinite series are also interesting? This is the topic of [17]. To summarize, the main contribution of this article is the demonstration of the working of the semiformal method in the fourteen cases that arise when Pólya’s theorem is generalized to $K_{\mathbb{N}}$ with complex edge weights.
1 Introduction
Let $d,n\in\mathbb{N}$ ($\{1,2,\dots\}$) and $\overline{v}\in\mathbb{Z}^{d}$, where $\mathbb{Z}=\{\dots,-1,0,1,\dots\}$ are the integers. Let $P_{d}(\overline{v},n)$ be the probability that a walk of length $n$ in the grid graph $\mathbb{Z}^{d}$ starting at $\overline{0}$ visits in a step $i>0$ the vertex $\overline{v}$. In 1921, G. Pólya proved in [24] a theorem, which we generalize, that for every vertex $\overline{v}$,
$\lim_{n\to\infty}P_{d}(\overline{v},\,n)\left\{\begin{array}[]{lll}=1&\dots&\text{if $d\leq 2$ and}\\ <1&\dots&\text{if $d\geq 3$}\,.\end{array}\right$.
Pólya’s theorem is almost always mentioned in the special case with $\overline{v}=\overline{0}$, but he proved it for any vertex $\overline{v}$ (see [16, Section 1] for a quote where he states his result). A selection of articles and books discussing Pólya’s theorem is [4, 5, 6, 9, 11, 13, 18, 19, 20, 21, 25, 26, 27, 31, 32].
We generalize it as follows. Let
$\mathbb{N}_{2}=\binom{\mathbb{N}}{2}=\{e\colon\;e\subset\mathbb{N}\text{ with }|e|=2\}$}
be edges of the countable complete graph $K_{\mathbb{N}}=\langle\mathbb{N},\mathbb{N}_{2}\rangle$. Here $|A|$ ($\in\mathbb{N}_{0}=\{0,1.\dots\}$) denotes the cardinality of a finite set $A$. Let
$h\colon\mathbb{N}_{2}\to\mathbb{C}$
be an edge weight, where $\mathbb{C}$ denotes the field of complex numbers, which is endowed with the usual Archimedean absolute value $|z|=\sqrt{z\cdot\overline{z}}$. Let $n\in\mathbb{N}_{0}$. We denote by $W(n)$ the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$. Thus $W(0)=\{\langle 1\rangle\}$ and for $n>0$ the set $W(n)$ is infinite and countable. Let $v\in\mathbb{N}$. We denote by $W(v,n)$ ($\subset W(n)$) the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$ and visiting $v$ in a step $i>0$. We extend the weight multiplicatively to
$h\colon\bigcup_{n=0}^{\infty}W(n)\to\mathbb{C}\,.$}
Let $W\subset\bigcup_{n=0}^{\infty}W(n)$. If the series
$\sum_{w\in W}h(w)$}
absolutely converges, we denote its sum by $h(W)$. (Results on series are reviewed in the next section.) Suppose that for every $n\in\mathbb{N}_{0}$ the series $\sum_{w\in W(n)}h(w)$ absolutely converges. Then so does the series $\sum_{w\in W(v,n)}h(w)$ and we consider the generating function
$A_{h,v}(x)=\sum_{n\geq 0}h(W(v,\,n))x^{n}\ \ (\in\mathbb{C}[[x]])\,.$}
Our generalizations of Pólya’s theorem are formulas for three quantities:
$\lim_{x\nearrow 1}A_{h,v}(x),\ A_{h,v}(1)=\sum_{n\geq 0}h(W(v,\,n))\,\text{ and }\,\lim_{n\to\infty}h(W(v,\,n))\,$,}
where $\lim_{x\nearrow 1}$ means that $x$ goes to $1$ through values $x\in[0,1)$. In some cases, these quantities are twisted by powers $\xi^{n}$ for $\xi\in\mathbb{C}^{*}$ ($\mathbb{C}\setminus\{0\}$). The formulas appear in fourteen theorems: Theorems 3.23.7 in Section 3 and Theorems 4.34.10 in Section 4. Pólya’s theorem is subsumed in five of them. Theorems 3.5 and 3.7 extend cases $d\geq 3$ and $d\leq 2$, respectively, when $\overline{v}=\overline{0}$. Theorems 4.7 (4.8) and 4.10 extend cases $d\geq 3$ and $d\leq 2$, respectively, when $\overline{v}\neq\overline{0}$.
The fourteen generalizations come from three sources. First, weights in Pólya’s theorem are nonnegative real numbers, and the convergence of infinite series is absolute. For our complex weights, we consider two kinds of convergence: absolute and conditional. We also consider real series with sums $\pm\infty$. Second, Pólya’s theorem is on probabilities, which means that the weights of edges incident to a fixed vertex sum up to $\xi=1$. We generalize it to $\xi$-convexity for any $\xi\in\mathbb{C}$; the case of $\xi\neq 0$ is normalized to $|\xi|=1$ and the case $\xi=0$ is treated separately (it corresponds to Kirchhoff’s first law). Thus we consider two cases, general weights and $\xi$-convex weights. Third, Pólya’s theorem is, in reality, two theorems because the cases $\overline{v}=\overline{0}$ and $\overline{v}\neq\overline{0}$ differ in the complexity of proofs. In our generalization, we correspondingly distinguish between cases $v=1$ and $v\neq 1$. Now $3\cdot 2\cdot 2=12$ and two additional generalizations arise by varying the assumptions.
We explain how our approach includes Pólya’s original problem. If $f\colon A\to B$ is a map and $C$ is any set (not necessarily a subset of $A$), we denote by
$f[C]=\{f(x)\colon\;x\in A\cap C\}\ \ (\subset B)$
the image of $C$ by $f$. Let $d\in\mathbb{N}$ and $\overline{v}\in\mathbb{Z}^{d}$. We take any bijection $f\colon\mathbb{Z}^{d}\to\mathbb{N}$ such that $f(\overline{0})=1$, and set $v=f(\overline{v})$. The grid graph $\mathbb{Z}^{d}$, which is the graph $\langle\mathbb{Z}^{d},E_{d}\rangle$ with the edges
$\{\overline{a},\,\overline{b}\}\in E_{d}\iff\sum_{i=1}^{d}|a_{i}-b_{i}|=1\,$,}
determines a weight $h\colon\mathbb{N}_{2}\to\{\frac{1}{2d},0\}$ by the relation $h(e)\neq 0$ iff $e=f[e^{\prime}]$ for $e^{\prime}\in E_{d}$. Then for every $n\in\mathbb{N}_{0}$ we have
$P_{d}(\overline{v},\,n)=h(W(v,\,n))\,.$
See Section 5 for the importance of (the general form of) this equality.
We structure our article as follows. In Section 2 we review infinite series with complex summands and the domain $\mathbb{C}[[x]]$ of generating functions. Sections 3 and 4 contain our main results, the mentioned fourteen generalizations of Pólya’s theorem. Section 5 contains concluding remarks. We continue our investigation in [17]; see Section 5 for more details on [17].
In Sections 3 and 4 we use what we term the semiformal method in (enumerative) combinatorics. It extends the symbolic method in combinatorics [3, 8, 10, 12, 22, 23, 28] from finite to countable sets. In our situation such extension is natural and inevitable, coefficients in generating functions $A_{h,v}(x)$ arise as sums of infinite series $\sum_{w\in W(v,n)}h(w)$. We develop this method more in [14]. Why “semiformal”? Generating “functions” in the symbolic method are formal power series and their coefficients are determined by finite operations. We also use formal power series, but our extension allows operations based on limit transitions. See [15] for more applications of the semiformal method.
2 Series and generating functions in $\mathbb{C}$
We review series and generating functions in $\mathbb{C}$. A series is a map $h\colon X\to\mathbb{C}$ defined on a finite or countable set $X$; we write it as $\sum_{x\in X}h(x)$. We say that $\sum_{x\in X}h(x)$ absolutely converges if two equivalent conditions hold.
  1. 1.
    For some constant $c\geq 0$ and every finite set $Y\subset X$ we have
    $\sum_{x\in Y}|h(x)|\leq c\,.$}
  2. 2.
    If the set $X$ is countable, then for every bijection $f\colon\mathbb{N}\to X$ the proper limit
    $s=\lim_{n\to\infty}{\textstyle\sum_{i=1}^{n}h(f(i))\ \ (\in\mathbb{C})}$
    exists and does not depend on $f$.
(In fact, the independence of $s$ on $f$ may be omitted from the formulation.) We call $s$ the sum of the series and denote it again by $\sum_{x\in X}h(x)$. If the set $X$ is finite and nonempty, $X=\{x_{1},x_{2},\dots,x_{n}\}$, then every series $\sum_{x\in X}h(x)$ absolutely converges and has the sum
$h(x_{1})+h(x_{2})+\dots+h(x_{n})\,.$
If $X=\emptyset$, we define the sum as 0. If
$U=U(x)=\sum_{n\geq 0}u_{n}x^{n}\ \ (\in\mathbb{C}[[x]])$}
is a formal power series, so that $x$ is a formal variable, we denote for any $c\in\mathbb{C}$ by $F_{U}(c)$ the sum of the power series $\sum_{n\geq 0}u_{n}c^{n}$, if it absolutely converges. Let
$H_{1}={\textstyle\{U(x)\in\mathbb{C}[[x]]\colon\;\text{$U(c)$ absolutely converges for every $c\in\mathbb{C}_{<1}$}\}\,,}$
where $\mathbb{C}_{<1}=\{c\in\mathbb{C}\colon\;|c|<1\}$. Thus for any $U\in H_{1}$ the sums form the function $F_{U}\colon\mathbb{C}_{<1}\to\mathbb{C}$.
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$. A walk $w$ in $K_{\mathbb{N}}$ is any $(n+1)$-tuple $w$ of vertices $v_{i}\in\mathbb{N}$,
$w=\langle v_{0},\,v_{1},\,\dots,\,v_{n}\rangle\,$,
such that $n\in\mathbb{N}_{0}$ and $v_{i-1}\neq v_{i}$ for every $i\in[n]$ ($\{1,2,\dots,n\}$, $0]=\emptyset$). The length of $w$ is $n$ ($\in\mathbb{N}_{0}$). We extend $h$ to walks. For $n>0$ we set
$h(w)=\prod_{i=1}^{n}h(\{v_{i-1},\,v_{i}\})\,$,}
and for $n=0$ we define $h(w)=1$. Recall that $W(n)$ is the set of walks $w$ in $K_{\mathbb{N}}$ with length $n$ such that $v_{0}=1$.
Definition 2.1 (lightness)
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$. We call the weight $h$ light if for every $n\in\mathbb{N}_{0}$, the series
$\sum_{w\in W(n)}h(w)$}
absolutely converges.
Let $S=\sum_{x\in X}h(x)$ be a series. If $Y\subset X$, we say that $\sum_{x\in Y}h(x)$ is a subseries of $S$. Any subseries of an absolutely convergent series absolutely converges. However, we cannot compare the sizes of sums
$\big|\sum_{x\in Y}h(x)\big|\,\text{ and }\,\big|\sum_{x\in X}h(x)\big|\,.$}
Except when $h\colon X\to\mathbb{R}_{\geq 0}$: if $\sum_{x\in X}h(x)$ absolutely converges and $Y\subset X$, then the sums satisfy
$0\leq\sum_{x\in Y}h(x)\leq\sum_{x\in X}h(x)$}
— we call it the subseries bound principle.
If $R=\sum_{x\in X}g(x)$ and $S=\sum_{x\in X}h(x)$ are series and $\alpha,\beta\in\mathbb{C}$, the series
$\alpha R+\beta S=\sum_{x\in X}(\alpha g(x)+\beta h(x))$}
is the linear combination of $R$ and $S$. We omit the easy proof of the next result.
Proposition 2.2
If $R$ and $S$ absolutely converge and have sums $r$ and $s$, respectively, then $\alpha R+\beta S$ absolutely converges and has sum $\alpha r+\beta s$.
The next approximation lemma is useful.
Lemma 2.3
If $S=\sum_{x\in X}h(x)$ absolutely converges and has sum $s$, then for every $\varepsilon>0$ there exists a finite set $Y\subset X$, denoted by $Y(S,\,\varepsilon)$, such that for every finite set $Z$ with $Y\subset Z\subset X$,
$\big|s-\sum_{x\in Z}h(x)\big|\leq\varepsilon\,.$}
Proof. Let an $\varepsilon>0$ be given. For finite $X$ we set $Y=X$. For countable $X$ we take any bijection $f\colon\mathbb{N}\to X$, take an $N\in\mathbb{N}$ such that
$\big|\sum_{n=1}^{N}h(f(n))-s\big|\leq\frac{\varepsilon}{2}\,\text{ and }\,\sum_{n>N}|h(f(n))|\leq\frac{\varepsilon}{2}\,$,}
and set $Y=\{f(1),f(2),\dots,f(N)\}$. Then for every finite set $Z$ with $Y\subset Z\subset X$ we have
$\big|s-\sum_{x\in Z}h(x)\big|\leq\big|s-\sum_{x\in Y}h(x)\big|+\sum_{x\in Z\setminus Y}|h(x)|\leq\frac{\varepsilon}{2}+\frac{\varepsilon}{2}=\varepsilon\,.$}
$\Box$
Two operations on series are grouping and product. A partition of a set $X$ is a set $P$ of nonempty and disjoint sets such that $\bigcup P=X$. If $X$ is at most countable, then so is $P$ and every set $Z\in P$. If $S=\sum_{x\in X}h(x)$ is a series and $P$ is a partition of $X$ such that for every $Z\in P$ the subseries $\sum_{x\in Z}h(x)$ absolutely converges and has sum $s_{Z}$, then the series
$S_{P}=\sum_{Z\in P}s_{Z}$}
is called the grouping of $S$.
Proposition 2.4
If $S=\sum_{x\in X}h(x)$ absolutely converges and has sum $s$, and if $P$ is a partition of $X$, then $S_{P}$ is correctly defined, absolutely converges and has the same sum $s$.
Proof. For every $Z\in P$ the series $S_{Z}=\sum_{x\in Z}h(x)$ is a subseries of $S$ and absolutely converges. Thus the series $S_{P}$ is correctly defined. We show that it absolutely converges. Let
$c=\sup(\{\sum_{x\in Z}|h(x)|\colon\;\text{$Z\subset X$ and $Z$ is finite}\})\ \ (<+\infty)\,.$}
Let $\{Z_{1},Z_{2},\dots,Z_{n}\}\subset P$ be a finite set. We use Lemma 2.3 and take for $i\in[n]$ finite sets
$Z_{i}^{\prime}=Y(S_{Z_{i}},\,2^{-i})\ \ (\subset Z_{i})\,.$
We form the disjoint union $Z_{0}=Z_{1}^{\prime}\cup\dots\cup Z_{n}^{\prime}$ ($\subset X$). Then
${\textstyle\sum_{i=1}^{n}\big|\sum_{x\in Z_{i}}h(x)\big|}$ $\leq$ ${\textstyle\sum_{i=1}^{n}\big|\sum_{x\in Z_{i}}h(x)-\sum_{x\in Z_{i}^{\prime}}h(x)\big|\,+}$
$+$ ${\textstyle\sum_{x\in Z_{0}}|h(x)|\leq\sum_{i=1}^{n}2^{-i}+c\leq 1+c}$
and $S_{P}$ absolutely converges.
Let $t$ be the sum of $S_{P}$. We show that $|t-s|\leq\varepsilon$ for every $\varepsilon>0$; then $t=s$. Let an $\varepsilon>0$ be given. We use Lemma 2.3 and take finite sets
$X^{\prime}=Y(S,\,\frac{\varepsilon}{3})\ \ (\subset X)\,\text{ and }\,P^{\prime}=Y(S_{P},\,\frac{\varepsilon}{3})\ \ (\subset P)\,.$}
We take a finite set $\{Z_{1},\dots,Z_{n}\}\subset P$ such that
$P^{\prime}\subset\{Z_{1},\,\dots,\,Z_{n}\}\,\text{ and }\,X^{\prime}\subset\bigcup_{i=1}^{n}Z_{i}\,.$}
We again use Lemma 2.3 and take for $i\in[n]$ finite sets
$Z_{i}^{\prime}=Y(S_{Z_{i}},\,2^{-i}\cdot\frac{\varepsilon}{3})\ \ (\subset Z_{i})\,.$}
For $i\in[n]$ let
$Z_{i}^{\prime\prime}=Z_{i}^{\prime}\cup(X^{\prime}\cap Z_{i})\ \ (\subset Z_{i})\,.$
We form the disjoint union
$X_{0}=\bigcup_{i=1}^{n}Z_{i}^{\prime\prime}\ \ (\subset X)\,.$}
Then $X_{0}$ is finite and $X^{\prime}\subset X_{0}$. Also, $Z_{i}^{\prime}\subset Z_{i}^{\prime\prime}$ for every $i\in[n]$. We have
$|s-t|$ $\leq$ ${\textstyle|s-\sum_{x\in X_{0}}h(x)|+\sum_{i=1}^{n}|\sum_{x\in Z_{i}^{\prime\prime}}h(x)-\sum_{x\in Z_{i}}h(x)|\,+}$
$+$ ${\textstyle|\sum_{i=1}^{n}\sum_{x\in Z_{i}}h(x)-t|\leq\frac{\varepsilon}{3}+\sum_{i=1}^{n}2^{-i}\cdot\frac{\varepsilon}{3}+\frac{\varepsilon}{3}\leq\varepsilon\,.}$
$\Box$
Let $X_{1}$, $X_{2}$ be sets and $X_{1}\times X_{2}$ be their Cartesian product. If $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i=1,2$, are series, then the series
$S_{1}\cdot S_{2}=\sum_{\langle x,\,y\rangle\in X_{1}\times X_{2}}h_{1}(x)h_{2}(y)$}
is the product of $S_{1}$ and $S_{2}$.
Proposition 2.5
If $S_{1}$ and $S_{2}$ absolutely converge and have sums $s_{1}$ and $s_{2}$, respectively, then $S_{1}\cdot S_{2}$ absolutely converges and has the sum $s_{1}s_{2}$.
Proof. We show that $S_{1}\cdot S_{2}$ absolutely converges. Let $c\geq 0$ be such that for any finite sets $Y_{i}\subset X_{i}$, $i=1,2$, we have $\sum_{x\in Y_{i}}|h_{i}(x)|\leq c$. Let $A\subset X_{1}\times X_{2}$ be a finite set. We take any finite sets $Y_{i}\subset X_{i}$ such that $A\subset Y_{1}\times Y_{2}$ and get the desired bound:
$\sum_{\langle x,\,y\rangle\in A}|h_{1}(x)h_{2}(y)|\leq\sum_{x\in Y_{1}}|h_{1}(x)|\cdot\sum_{y\in Y_{2}}|h_{2}(y)|\leq c\cdot c=c^{2}\,.$}
We show that the sum $s$ of $S_{1}\cdot S_{1}$ equals $s_{1}s_{2}$. It suffices to prove that for any $\varepsilon\in(0,1)$ we have $|s-s_{1}s_{2}|\leq\varepsilon$. So let $\varepsilon\in(0,1)$. Using Lemma 2.3 we take finite sets
$X^{\prime}_{1}=Y(S_{1},\,\frac{\varepsilon}{4(1+|s_{2}|)})\ \ (\subset X_{1}),\ \ X^{\prime}_{2}=Y(S_{2},\,\frac{\varepsilon}{4(1+|s_{1}|)})\ \ (\subset X_{2})\$}
and
$Z=Y(S_{1}\cdot S_{2},\frac{\varepsilon}{4})\ \ (\subset X_{1}\times X_{2})\,.$}
We take finite sets $X_{1}^{\prime\prime}$ and $X_{2}^{\prime\prime}$ such that $X_{1}^{\prime}\subset X_{1}^{\prime\prime}\subset X_{1}$, $X_{2}^{\prime}\subset X_{2}^{\prime\prime}\subset X_{2}$ and $Z\subset X_{1}^{\prime\prime}\times X_{2}^{\prime\prime}$. Then $|s-s_{1}s_{2}|$ is at most
${\textstyle\big|s-\sum_{\langle x,\,y\rangle\in X_{1}^{\prime\prime}\times X_{2}^{\prime\prime}}h_{1}(x)h_{2}(y)\big|+\big|\sum_{x\in X_{1}^{\prime\prime}}h_{1}(x)\sum_{y\in X_{2}^{\prime\prime}}h_{2}(y)-s_{1}s_{2}\big|}$
${\textstyle\leq\frac{\varepsilon}{4}+|(s_{1}+t_{1})(s_{2}+t_{2})-s_{1}s_{2}|\,,}$
where $|t_{1}|\leq\frac{\varepsilon}{4(1+|s_{2}|)}$ and $|t_{2}|\leq\frac{\varepsilon}{4(1+|s_{1}|)}$. Hence $|s-s_{1}s_{2}|\leq\frac{\varepsilon}{4}+|s_{1}t_{2}|+|s_{2}t_{1}|+|t_{1}t_{2}|\leq\frac{\varepsilon}{4}+\frac{\varepsilon}{4}+\frac{\varepsilon}{4}+\frac{\varepsilon}{4}=\varepsilon$. $\Box$
Let $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i\in[k]$, be $k\in\mathbb{N}$ series. We define their product as the series
$\prod_{i=1}^{k}S_{i}=\sum_{\overline{x}\in\overline{X}}h_{1}(x_{1})h_{2}(x_{2})\dots h_{k}(x_{k})\,$,}
where $\overline{x}=\langle x_{1},\dots,x_{k}\rangle\in X_{1}\times\dots\times X_{k}=\overline{X}$. Iterating Proposition 2.5 we obtain the following result.
Proposition 2.6
If series $S_{1}$, $\dots$, $S_{k}$ absolutely converge and have sums $s_{1}$, $\dots$, $s_{k}$, respectively, then $\prod_{i=1}^{k}S_{i}$ absolutely converges and has sum $s_{1}s_{2}\dots s_{k}$.
Interestingly, a reversal of Proposition 2.5 holds.
Corollary 2.7
If nonempty series $S_{1}$ and $S_{2}$ are not identically zero and if their product $S_{1}\cdot S_{2}$ absolutely converges and has sum $s$, then $S_{1}$ and $S_{2}$ absolutely converge and their respective sums $s_{1}$ and $s_{2}$ satisfy $s_{1}s_{2}=s$.
Proof. This follows from Proposition 2.5 if we show that $S_{1}$ and $S_{2}$ absolutely converge. Let $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i=1,2$, and let $c>0$ be such that for every finite set $X\subset X_{1}\times X_{2}$ we have
$\sum_{\langle x,y\rangle\in X}|h_{1}(x)h_{2}(y)|\leq c\,.$}
We prove absolute convergence of $S_{1}$. We take any $y\in X_{2}$ with $h_{2}(y)\neq 0$. Then for every finite set $X\subset X_{1}$ we have
$\sum_{x\in X}|h_{1}(x)|=\frac{1}{|h_{2}(y)|}\sum_{\langle x,y\rangle\in X\times\{y\}}|h_{1}(x)h_{2}(y)|\leq\frac{c}{|h_{2}(y)|}\,.$}
Thus $S_{1}$ absolutely converges. For $S_{2}$ we argue similarly. $\Box$
It is clear that the assumptions on $S_{1}$ and $S_{2}$ cannot be omitted.
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and $n\in\mathbb{N}_{0}$. We obtain a sufficient condition for lightness of the weight $h$ in terms of its values. We denote by $V(h,n)$ the set of vertices in $K_{\mathbb{N}}$ reachable from $1$ by a walk with nonzero weight and length at most $n$.
Definition 2.8 (slimness)
A weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is slim if for every $n\in\mathbb{N}_{0}$ there is a constant $c=c(n)>0$ such that for every vertex $u\in V(h,n)$ and every finite set $X\subset\mathbb{N}\setminus\{u\}$ we have
$\sum_{v\in X}|h(\{u,\,v\})|\leq c\,.$}
Proposition 2.9
Every slim weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is light.
Proof. Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be slim. We show by induction on $n\in\mathbb{N}_{0}$ that the series
$\sum_{w\in W(n)}h(w)$}
absolutely converges. For $n=0$ it is trivial because $W(0)=\{\langle 1\rangle\}$. Let $n>0$ and let $c>0$ be a constant such that
$\sum_{w\in X}|h(w)|\leq c\,\text{ and }\,\sum_{v\in X^{\prime}}|h(\{u,\,v\})|\leq c$}
for every finite set $X\subset W(n-1)$, every $u\in V(h,n-1)$ and every finite set $X^{\prime}\subset\mathbb{N}\setminus\{u\}$. Let $X\subset W(n)$ be finite. We may assume that all walks in $X$ have nonzero weights. We decompose every walk $w\in X$ as $w=w^{\prime}\,\ell(w)$, where $w^{\prime}\in W(n-1)$ and $\ell(w)$ is the last vertex. Let $Y$ be the set of walks $w^{\prime}$. For $w^{\prime}\in Y$ let
$X(w^{\prime})=\{v\in\mathbb{N}\setminus\{\ell(w^{\prime})\}\colon\;w^{\prime}v\in X\}\,.$
We get the bound
${\textstyle\sum_{w\in X}|h(w)|}$ $=$ ${\textstyle\sum_{w^{\prime}\in Y}|h(w^{\prime})|\cdot\sum_{v\in X(w^{\prime})}|h(\{\ell(w^{\prime}),\,v\})|}$
$\leq$ ${\textstyle\sum_{w^{\prime}\in Y}|h(w^{\prime})|\cdot c\leq c^{2}\,.}$
$\Box$
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$. We define $V(h)=\bigcup_{n\geq 0}V(h,n)$. It is the set of vertices in $K_{\mathbb{N}}$ reachable from $1$ by walks with nonzero weights. We define the graph
$G(h)=\langle V(h),\,E(h)\rangle$
as the connected component of $1$ in the subgraph of $K_{\mathbb{N}}$ formed by edges with nonzero weight. It follows that if $G(h)$ has finite degrees then $h$ is slim and hence light.
Definition 2.10 ($\xi$-convexity)
Let $\xi\in\mathbb{C}$. We call a light weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ $\xi$-convex if for every vertex $u\in V(h)$ the sum
$\sum_{e\in E(h),\,e\ni u}h(e)=\sum_{v\in\mathbb{N}\setminus\{u\}}h(\{u,\,v\})=\xi\,.$}
We say that $h$ is convex if it is $1$-convex.
We show that the sums in the definition exist.
Proposition 2.11
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be light and $u\in V(h)$. Then the series
$\sum_{v\in\mathbb{N}\setminus\{u\}}h(\{u,v\})$}
absolutely converges.
Proof. There is a walk $w_{0}=\langle u_{0},\dots,u_{n}\rangle$ in $G(h)$ such that $u_{0}=1$ and $u_{n}=u$. Let $c=h(w_{0})$ ($\neq 0$) and $W$ be the set of walks $\langle u_{0},\dots,u_{n},u_{n+1}\rangle$ in $K_{\mathbb{N}}$. Thus vertices $u_{0}$, $\dots$, $u_{n}$ are fixed and $u_{n+1}$ runs in $\mathbb{N}\setminus\{u_{n}\}=\mathbb{N}\setminus\{u\}$. The series $\sum_{w\in W}h(w)$ is a subseries of $\sum_{w\in W(n+1)}h(w)$ and therefore absolutely converges. By Proposition 2.2,
$\sum_{v\in\mathbb{N}\setminus\{u\}}h(\{u,v\})=\frac{1}{c}\sum_{w\in W}h(w)$}
absolutely converges. $\Box$
We survey results on generating functions. We use univariate formal power series $U=U(x)=\sum_{n\geq 0}u_{n}x^{n}$ with complex coefficients. They form an integral domain denoted by $\mathbb{C}[[x]]$. For $U$ and $V$ in $\mathbb{C}[[x]]$,
$U+V=\sum_{n\geq 0}(u_{n}+v_{n})x^{n}\,\text{ and }\,U\cdot V=\sum_{n\geq 0}\big(\sum_{j=0}^{n}u_{j}v_{n-j}\big)x^{n}\,.$}
Neutral elements are $0=0+0x+0x^{2}+\dots$ and $1=1+0x+0x^{2}+\cdots$. Units in $\mathbb{C}[[x]]$ are exactly the formal power series $\sum_{n\geq 0}u_{n}x^{n}$ with $u_{0}\neq 0$.
Let $\mathbb{C}_{1}=\{x\in\mathbb{C}\colon\;|x|=1\}$ and $\xi\in\mathbb{C}_{1}$. We define
$\xi\mathbb{R}_{\geq 0}=\{\xi x\colon\;x\in\mathbb{R}_{\geq 0}\}\ \ (\subset\mathbb{C})\,.$
It is the ray in $\mathbb{C}$ going from the origin in the direction $\xi$. Clearly, $\xi\mathbb{R}_{\geq 0}$ is closed under addition. If $\alpha_{j}\in\xi_{j}\mathbb{R}_{\geq 0}$, $j=1,2$, then $\alpha_{1}\alpha_{2}\in(\xi_{1}\xi_{2})\mathbb{R}_{\geq 0}$. If
$h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$
is a light weight and $n\in\mathbb{N}_{0}$ then $h(W(n))\in\xi^{n}\mathbb{R}_{\geq 0}$. The lines $\xi\mathbb{R}$ in $\mathbb{C}$ going through the origin in the direction $\xi$ have the same properties. For $\alpha\in\mathbb{C}$ and $U=U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{C}[[x]]$ we define
$(U)_{\alpha}=(U)_{\alpha}(x):=U(\alpha x)=\sum_{n\geq 0}u_{n}\alpha^{n}x^{n}\ \ (\in\mathbb{C}[[x]])\,.$}
Note that $(U+V)_{\alpha}=(U)_{\alpha}+(V)_{\alpha}$ and $(U\cdot V)_{\alpha}=(U)_{\alpha}\cdot(V)_{\alpha}$.
Definition 2.12 ($U(1)_{1}$ and $U(1)_{2}$)
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$ be in $\mathbb{C}[[x]]$. We introduce the following terminology and objects.
  1. 1.
    If the (proper) limit
    $U(1)_{1}:=\lim_{n\to\infty}{\textstyle\sum_{j=0}^{n}u_{j}\ \ (\in\mathbb{C})}$
    exists, we say that $U(1)_{1}$ exists.
  2. 2.
    If the series
    $\sum_{n\geq 0}u_{n}1^{n}=\sum_{n\geq 0}u_{n}$}
    absolutely converges, we denote the sum by $U(1)_{2}$ and say that $U(1)_{2}$ exists.
If $U(1)_{2}$ exists, then $U(1)_{1}$ exists and $U(1)_{1}=U(1)_{2}$. The opposite implication is, in general, not true. For example, if
$U(x)=\sum_{n\geq 1}(-1)^{n-1}\frac{1}{n}x^{n}$}
then $U(1)_{1}=\log 2$, but $U(1)_{2}$ does not exist. But if
$U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{R}_{\geq 0}[[x]]$}
and $U(1)_{1}$ exists, then $U(1)_{2}$ exists and $U(1)_{1}=U(1)_{2}$. The following is a corollary of the well known Hadamard’s formula for the radius of convergence of a formal power series.
Proposition 2.13
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{C}[[x]]$. Then $U(x)\in H_{1}$ iff
$\limsup_{n\to\infty}|u_{n}|^{1/n}\leq 1\,.$
Thus, for example, the existence of $U(1)_{1}$ is a sufficient (but not necessary) condition for $U(x)\in H_{1}$.
Note that if $U\in\mathbb{C}[[x]]$ then $(U)_{\xi}\in H_{1}$ either for all or no $\xi\in\mathbb{C}_{1}$.
Definition 2.14 ($U(1)_{1}=\pm\infty$)
If $U(x)=\sum_{n\geq 0}u_{n}x^{n}$ is in $\mathbb{R}[[x]]$ and
$\lim_{n\to\infty}\sum_{j=0}^{n}u_{j}=+\infty\,$,}
we say that $U(1)_{1}=+\infty$. We similarly define that $U(1)_{1}=$-$\infty$.
Proposition 2.15
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$ and $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ be in $\mathbb{C}[[x]]$, and let $\xi\in\mathbb{C}^{*}$. If $(V)_{1/\xi}(1)_{1}$ exists and $U(x)=V(x)\frac{1}{1-\xi x}$, then
$\lim_{n\to\infty}\xi^{-n}u_{n}=(V)_{1/\xi}(1)_{1}\,.$
Proof. Let $n\in\mathbb{N}_{0}$. Since $\frac{1}{1-\xi x}=1+\xi x+\xi^{2}x^{2}+\dots$, we have
$u_{n}=v_{0}\xi^{n}+v_{1}\xi^{n-1}+\dots+v_{n}\xi^{0}=v_{0}\cdot\xi^{n}+v_{1}\xi^{-1}\cdot\xi^{n}+\dots+v_{n}\xi^{-n}\cdot\xi^{n}\,.$
Dividing by $\xi^{n}$ and taking the limit for $n\to\infty$, we get the stated formula. $\Box$
We manipulate the quantities $U(1)_{1}$ by means of three Abelian theorems, which we state only for the radius of convergence $R\geq 1$. The first theorem is classical, see [1, 2] and [29, Chapter II.7].
Theorem 2.16 (Abel 1)
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{C}[[x]]$. If $U(1)_{1}$ exists, so that $U(x)\in H_{1}$ by Proposition 2.13, then
$\lim_{x\nearrow 1}F_{U}(x)=U(1)_{1}\ \ (\in\mathbb{C})\,.$
The second Abelian theorem is rarely mentioned, but see [2] for a simple proof. It is given only for the case $U(1)_{1}=+\infty$, but it works for the sum and limit -$\infty$ as well.
Theorem 2.17 (Abel 2)
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{R}[[x]]\cap H_{1}$. If $U(1)_{1}=\pm\infty$ then
$\lim_{x\nearrow 1}F_{U}(x)=\pm\infty\,$,
with equal signs.
The third Abelian theorem allows only nonnegative real coefficients. It has the advantage over the previous two theorems that it is an equivalence. Now one can go not only from the sum to the functional limit, but also in the reverse. We omit the easy proof.
Theorem 2.18 (Abel 3)
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{R}_{\geq 0}[[x]]\cap H_{1}$. Then the following limit and sum always exist (and are possibly $\infty$) and are equal,
$\lim_{x\nearrow 1}F_{U}(x)=U(1)_{1}\ \ (\in[0,\,+\infty)\cup\{+\infty\})\,.$
Proposition 2.19
Let $\alpha,\beta\in\mathbb{C}$ and $U(x),V(x),W(x)\in\mathbb{C}[[x]]$. If $U(x)=\alpha V(x)+\beta W(x)$ and if $V(1)_{1}$ and $W(1)_{1}$ exist, then
$U(1)_{1}=\alpha V(1)_{1}+\beta W(1)_{1}\,.$
Proof. Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$, $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ and $W(x)=\sum_{n\geq 0}w_{n}x^{n}$. Then
$\sum_{j=0}^{n}u_{j}=\sum_{j=0}^{n}(\alpha v_{j}+\beta w_{j})=\alpha\sum_{j=0}^{n}v_{j}+\beta\sum_{j=0}^{n}w_{j}\,.$}
Taking the limit for $n\to\infty$ we get that that $U(1)_{1}=\alpha V(1)_{1}+\beta W(1)_{1}$. $\Box$
Here the formulation “then $U(1)_{1}=\alpha V(1)_{1}+\beta W(1)_{1}$” abbreviates “then $U(1)_{1}$ exists and $U(1)_{1}=\alpha V(1)_{1}+\beta W(1)_{1}$”. We will use such abbreviated formulations. This result can be easily generalized to linear combinations with more than two terms. Next we consider products of formal power series.
Proposition 2.20
Let $U(x),V_{1}(x),\dots,V_{n}(x)\in\mathbb{C}[[x]]$ be $n+1$, $n\in\mathbb{N}$, formal power series. If $U(x)=\prod_{i=1}^{n}V_{i}(x)$ and $U(1)_{1}$, $V_{1}(1)_{1}$, $\dots$, $V_{n}(1)_{1}$ exist, then
$U(1)_{1}=\prod_{i=1}^{n}V_{i}(1)_{1}\,.$}
Proof. By Propositions 2.4, 2.6 and 2.13, and by the assumptions we have
$F_{U}(x)=\prod_{i=1}^{n}F_{V_{i}}(x)$ for every $x\in\mathbb{C}_{<1}$
(see proofs of Propositions 2.24 and 2.27 for details of similar arguments). Using Theorem 2.16 and properties of limits of functions we get
$U(1)_{1}$ $=$ $\lim_{x\nearrow 1}F_{U}(x)=\lim_{x\nearrow 1}{\textstyle\prod_{i=1}^{n}F_{V_{i}}(x)}$
$=$ ${\textstyle\prod_{i=1}^{n}\lim_{x\nearrow 1}F_{V_{i}}(x)=\prod_{i=1}^{n}V_{i}(1)_{1}}\,.$
$\Box$
Since we have to assume that $U(1)_{1}$ exists, the case $n>2$ does not follow by iterating the case $n=2$. The example in [7] shows that the assumption of existence of $U(1)_{1}$ cannot be omitted.
Proposition 2.21
Let $U(x),V(x)\in\mathbb{C}[[x]]$. If $U(x)=\frac{1}{V(x)}$, and if $U(1)_{1}$ and $V(1)_{1}$ exist, then $V(1)_{1}\neq 0$ and
$U(1)_{1}=\frac{1}{V(1)_{1}}\,.$
Proof. In $\mathbb{C}[[x]]$ we have the identity $U(x)V(x)=1$. Proposition 2.20 gives that $U(1)_{1}V(1)_{1}=1$. Thus $V(1)_{1}\neq 0$ and dividing by $V(1)_{1}$ gives the stated formula. $\Box$
Proposition 2.22
Let $U(x),V(x),W(x)\in\mathbb{C}[[x]]$. If $U(x)^{2}=1-\frac{V(x)}{W(x)}$, if $U(1)_{1}$, $V(1)_{1}$ and $W(1)_{1}$ exist, and if $W(1)_{1}\neq 0$, then
$\big(U(1)_{1}\big)^{2}=1-\frac{V(1)_{1}}{W(1)_{1}}\,.$
Proof. In $\mathbb{C}[[x]]$ we have the identity $U(x)^{2}\cdot W(x)=W(x)-V(x)$. Let T(x)=W(x)V(x)T(x)=W(x)-V(x). By Proposition 2.19, $T(1)_{1}$ exists and $T(1)_{1}=W(1)_{1}-V(1)_{1}$. By Proposition 2.20 we have $(U(1)_{1})^{2}\cdot W(1)_{1}=T(1)_{1}=W(1)_{1}-V(1)_{1}$. Dividing by $W(1)_{1}$ we get the stated formula. $\Box$
We obtain analogs of the previous results in absolute convergence.
Proposition 2.23
Let $\alpha,\beta\in\mathbb{C}$ and $U(x),V(x),W(x)\in\mathbb{C}[[x]]$. If $U(x)=\alpha V(x)+\beta W(x)$ and if $V(1)_{2}$ and $W(1)_{2}$ exist, then
$U(1)_{2}=\alpha V(1)_{2}+\beta W(1)_{2}\,.$
Proof. Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$, $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ and $W(x)=\sum_{n\geq 0}w_{n}x^{n}$. Since $V(1)_{2}$ and $W(1)_{2}$ are sums of absolutely convergent series $\sum_{n\in\mathbb{N}_{0}}v_{n}$ and $\sum_{n\in\mathbb{N}_{0}}w_{n}$, respectively, and $u_{n}=\alpha v_{n}+\beta w_{n}$, the result follows from Proposition 2.2. $\Box$
Proposition 2.24
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$, $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ and $W(x)=\sum_{n\geq 0}w_{n}x^{n}$ be in $\mathbb{C}[[x]]$. If $U(x)=V(x)W(x)$ and if $V(1)_{2}$ and $W(1)_{2}$ exist, then the following holds.
  1. 1.
    $U(1)_{2}$ exists and $U(1)_{2}=V(1)_{2}W(1)_{2}$.
  2. 2.
    The sum $\sum_{n\geq 0}|u_{n}|\leq\sum_{n\geq 0}|v_{n}|\cdot\sum_{n\geq 0}|w_{n}|$.
Proof. 1. We consider the series
$S=\sum_{\langle j,\,k\rangle\in\mathbb{N}_{0}^{2}}v_{j}w_{k}=\sum_{j\in\mathbb{N}_{0}}v_{j}\cdot\sum_{k\in\mathbb{N}_{0}}w_{k}\,.$}
By Proposition 2.5 it absolutely converges and has sum $V(1)_{2}W(1)_{2}$. On the other hand, $S$ has the grouping
$\sum_{n\geq 0}\sum_{j=0}^{n}v_{j}w_{n-j}\,$,}
where the inner $\sum$ means sum, the outer $\sum$ means series, and $\sum_{j=0}^{n}v_{j}w_{n-j}=u_{n}$. By Proposition 2.4, $S$ has the sum $u_{0}+u_{1}+\dots=U(1)_{2}$. Thus $U(1)_{2}=V(1)_{2}W(1)_{2}$.
2. This follows from the inequality ($n\in\mathbb{N}_{0}$)
$\sum_{m=0}^{n}|u_{m}|=\sum_{m=0}^{n}\big|\sum_{j=0}^{m}v_{j}w_{m-j}\big|\leq\sum_{j=0}^{n}|v_{j}|\cdot\sum_{j=0}^{n}|w_{j}|\,.$}
$\Box$
For nonnegative real coefficients we can go in the reverse, from the product to the factors.
Proposition 2.25
Let $U(x)$, $V(x)$ and $W(x)$ be in $\mathbb{R}_{\geq 0}[[x]]$. Suppose that $U(x)=V(x)W(x)$, that $U(x)$, $V(x)$ and $W(x)$ are in $H_{1}$, $V(x)W(x)\neq 0$ and that $U(1)_{2}$ exists. Then $V(1)_{2}$ and $W(1)_{2}$ exist, and
$U(1)_{2}=V(1)_{2}W(1)_{2}\,.$
Proof. By Theorem 2.18 and the assumption $V(x)W(x)\neq 0$ we have $V(1)_{2}=V(1)_{1}=\lim_{x\nearrow 1}F_{V}(x)>0$ and the same for $W(x)$. Since $F_{U}=F_{V}F_{W}$ on $\mathbb{C}_{<1}$, we have
$U(1)_{2}$ $=$ $U(1)_{1}=\lim_{x\nearrow 1}F_{U}(x)$
$=$ $\lim_{x\nearrow 1}F_{V}(x)\cdot\lim_{x\nearrow 1}F_{W}(x)=V(1)_{1}W(1)_{1}=V(1)_{2}W(1)_{2}\,.$
In the second equality we used Theorem 2.18. The third equality holds because the product of limits is neither $0\cdot(+\infty)$ nor $(+\infty)\cdot 0$. $\Box$
Proposition 2.20 for $n=2$ and item 1 of Proposition 2.24 are incomparable. The former assumes only conditional convergence, but for all three series. The latter assumes stronger absolute convergence, but only for two series. Another variant of the proposition is due to F. Mertens.
Theorem 2.26 (Mertens’s)
Suppose that $U(x)$, $V(x)$ and $W(x)$ are in $\mathbb{C}[[x]]$. If $U(x)=V(x)W(x)$, and if $V(1)_{1}$ and $W(1)_{2}$ exist, then
$U(1)_{1}=V(1)_{1}W(1)_{2}\,.$
See [7] for a proof.
Proposition 2.27
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$ and $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ be in $\mathbb{C}[[x]]$ and $v_{0}=1$. If $U(x)=\frac{1}{V(x)}$ and $c=|v_{1}|+|v_{2}|+\dots<1$, then the following holds.
  1. 1.
    $V(1)_{2}$ and $U(1)_{2}$ exist, $V(1)_{2}\neq 0$ and $U(1)_{2}=\frac{1}{V(1)_{2}}$.
  2. 2.
    The sum $\sum_{n\geq 0}|u_{n}|\leq\frac{1}{1-c}$.
Proof. 1. From the assumption on $c$ it follows that $V(1)_{2}$ exists. From
$|V(1)_{2}-1|\leq c<1$
we get $V(1)_{2}\neq 0$. We consider the series
$S=1+\sum_{\begin{subarray}{c}m\in\mathbb{N}\\ j_{m,1},\,j_{m,2},\,\dots,\,j_{m,m}\in\mathbb{N}\end{subarray}}(-1)^{m}v_{j_{m,1}}v_{j_{m,2}}\dots v_{j_{m,m}}\,.$}
We show that $S$ absolutely converges. This follows from the bound that the sum of absolute values of summands in any finite subseries of $S$ is at most
$1+\sum_{m\geq 1}(|v_{1}|+|v_{2}|+\dots)^{m}=\frac{1}{1-c}\,.$}
Let $m\in\mathbb{N}$. We consider the series
$S_{m}=\prod_{i=1}^{m}(v_{1}+v_{2}+\dots)=\sum_{\langle j_{1},\,\dots,\,j_{m}\rangle\in\mathbb{N}^{m}}v_{j_{1}}v_{j_{2}}\dots v_{j_{m}}\,.$}
By Proposition 2.6, it absolutely converges and has sum $(V(1)_{2}-1)^{m}$. Thus by Propositions 2.2 and 2.4, $S$ has sum
$1+\sum_{m\geq 1}(-1)^{m}(V(1)_{2}-1)^{m}=\frac{1}{V(1)_{2}}\,.$}
On the other hand, in $\mathbb{C}[[x]]$ we have the identity
$U(x)=1+\sum_{m\geq 1}(-1)^{m}(v_{1}x+v_{2}x^{2}+\dots)^{m}\,.$}
Thus $u_{0}=1$ and for every $n\in\mathbb{N}$,
$u_{n}=\sum_{m\geq 1}(-1)^{m}\sum_{\begin{subarray}{c}j_{m,1},\,j_{m,2},\,\dots,\,j_{m,m}\in\mathbb{N}\\ j_{m,1}+j_{m,2}+\dots+j_{m,m}=n\end{subarray}}v_{j_{m,1}}v_{j_{m,2}}\dots v_{j_{m,m}}\,.$}
Proposition 2.4 shows that $S$ also has sum $u_{0}+u_{1}+u_{2}+\cdots=U(1)_{2}$. The stated formula follows.
2. Indeed, the sum
${\textstyle\sum_{k=0}^{n}|u_{k}|=}$
${\textstyle=1+\sum_{k=1}^{n}\big|\sum_{m\geq 1}(-1)^{m}\sum_{\begin{subarray}{c}j_{m,1},\,j_{m,2},\,\dots,\,j_{m,m}\in\mathbb{N}\\ j_{m,1}+j_{m,2}+\dots+j_{m,m}=k\end{subarray}}v_{j_{m,1}}v_{j_{m,2}}\dots v_{j_{m,m}}\big|}$
${\textstyle\leq\sum_{m\geq 0}(|v_{1}|+|v_{2}|+\dots)^{m}=\frac{1}{1-c}\,.}$
$\Box$
Again, Proposition 2.21 and item 1 of Proposition 2.27 are incomparable.
The analog of Proposition 2.22 for absolute convergence requires two lemmas.
Lemma 2.28
Let $V(x)\in\mathbb{C}[[x]]$ and
$V(x)^{2}=tx^{2l}(1+t_{1}x+t_{2}x^{2}+\dots)$
where $l\in\mathbb{N}_{0}$ and $t,t_{j}\in\mathbb{C}$ — every square in $\mathbb{C}[[x]]$ has this form. Then there exists a number $v\in\mathbb{C}$ such that $v^{2}=t$ and
$V(x)=vx^{l}\sum_{n\geq 0}\binom{1/2}{n}(t_{1}x+t_{2}x^{2}+\dots)^{n}\,.$}
Proof. Let $n\in\mathbb{N}_{0}$ and a,ba,\,b be formal variables. By Vandermonde’s identity ([30]),
$\sum_{j=0}^{n}\binom{a}{j}\binom{b}{n-j}=\binom{a+b}{n}\ \ (\in\mathbb{Q}[a,\,b])\,.$}
Let $V_{0}(x)=\sum_{n\geq 0}\binom{1/2}{n}(t_{1}x+t_{2}x^{2}+\dots)^{n}$. For any $v_{0}\in\mathbb{C}$ such that $(v_{0})^{2}=t$ the instance $a=b=\frac{1}{2}$ of the identity gives
$\big(v_{0}x^{l}V_{0}(x)\big)^{2}$ $=$ ${\textstyle tx^{2l}\sum_{n\geq 0}\big(\sum_{j=0}^{n}\binom{1/2}{j}\binom{1/2}{n-j}\big)(t_{1}x+t_{2}x^{2}+\dots)^{n}}$
$=$ ${\textstyle tx^{2l}\sum_{n\geq 0}\binom{1}{n}(t_{1}x+t_{2}x^{2}+\dots)^{n}}$
$=$ $tx^{2l}(1+t_{1}x+t_{2}x^{2}+\dots)=V(x)^{2}\,.$
Hence $V(x)=\pm v_{0}x^{l}V_{0}(x)$. $\Box$
Lemma 2.29
Let $U(x),V(x)\in\mathbb{C}[[x]]$ and
$U(x)=V(x)^{2}=tx^{2l}(1+t_{1}x+t_{2}x^{2}+\dots)$
where $l\in\mathbb{N}_{0}$ and $t,t_{j}\in\mathbb{C}$. If $c=|t_{1}|+|t_{2}|+\dots<1$, then $U(1)_{2}$ and $V(1)_{2}$ exist, and
$U(1)_{2}=\big(V(1)_{2}\big)^{2}\,.$
Proof. It follows from the assumption on $c$ that $U(1)_{2}$ exists. By Lemma 2.28 there is a number $v\in\mathbb{C}$ such that $v^{2}=t$ and
$V(x)=\sum_{n\geq 0}v_{n}x^{n}=vx^{l}\sum_{n\geq 0}\binom{1/2}{n}(t_{1}x+t_{2}x^{2}+\dots)^{n}\,.$}
Thus $v_{n}=0$ for $n<l$, $v_{l}=v$ and for $n>l$,
$v_{n}=v\sum_{m\geq 1}\binom{1/2}{m}\sum_{\begin{subarray}{c}j_{m,1},\,j_{m,2},\,\dots,\,j_{m,m}\in\mathbb{N}\\ j_{m,1}+j_{m,2}+\dots+j_{m,m}=n-l\end{subarray}}t_{j_{m,1}}t_{j_{m,2}}\dots t_{j_{m,m}}\,.$}
We consider the series
$S=0+0+\dots+0+v+\sum_{\begin{subarray}{c}m\in\mathbb{N}\\ j_{m,1},\,j_{m,2},\,\dots,\,j_{m,m}\in\mathbb{N}\end{subarray}}v\binom{1/2}{m}t_{j_{m,1}}t_{j_{m,2}}\dots t_{j_{m,m}}$}
beginning with $l$ summands 0. We show that it absolutely converges. Let $T$ be a finite subseries of $S$. Since $\big|\binom{1/2}{m}\big|\leq 1$ for every $m\in\mathbb{N}$, the sum of absolute values of summands in $T$ is at most
$|v|+|v|\sum_{m\geq 1}(|t_{1}|+|t_{2}|+\dots)^{m}=\frac{|v|}{1-c}\,.$}
Thus $S$ absolutely converges. By Proposition 2.4, the series $S$ has the sum $v_{0}+v_{1}+\dots=V(1)_{2}$. By item 1 of Proposition 2.24, $U(1)_{2}=\big(V(1)_{2}\big)^{2}$. $\Box$
Proposition 2.30
Let $U(x),V(x)=\sum_{n\geq 0}v_{n}x^{n}$ and $W(x)=\sum_{n\geq 0}w_{n}x^{n}$ be in $\mathbb{C}[[x]]$. Suppose that $V(1)_{2}$ exists, $w_{0}=1$, $U(x)^{2}=1-\frac{V(x)}{W(x)}$ and $c=|w_{1}|+|w_{2}|+\dots<1$. If $V(x)=W(x)$, then $U(x)=0$. Else, if $k\in\mathbb{N}_{0}$ is minimum such that $v_{k}\neq w_{k}$ and
$\big(|w_{k}-v_{k}|\cdot(1-c)\big)^{-1}\sum_{n>k}|w_{n}-v_{n}|<1\,$,}
then $U(1)_{2}$ and $W(1)_{2}$ exist, $W(1)_{2}\neq 0$ and
$\big(U(1)_{2}\big)^{2}=1-\frac{V(1)_{2}}{W(1)_{2}}\,.$}
Proof. If $V(x)=W(x)$ then $U(x)^{2}=0$ and $U(x)=0$. Let $V(x)\neq W(x)$ and $k\in\mathbb{N}_{0}$ be as stated. We write
$U(x)^{2}=\frac{W(x)-V(x)}{W(x)}=tx^{2l}(1+t_{1}x+t_{2}x^{2}+\dots)$}
where $l\in\mathbb{N}_{0}$, $t=w_{k}-v_{k}\in\mathbb{C}^{*}$ and $t_{j}\in\mathbb{C}$. Let
$U_{0}(x)=\frac{1}{W(x)}=\sum_{n\geq 0}u_{0,n}x^{n},\ U_{1}(x)=W(x)-V(x)$}
and $U_{2}(x)=U_{0}(x)U_{1}(x)=\sum_{n\geq 0}u_{2,n}x^{n}$. By item 1 of Proposition 2.27, $U_{0}(1)_{2}$ and $W(1)_{2}$ exist, $W(1)_{2}\neq 0$ and $U_{0}(1)_{2}=\frac{1}{W(1)_{2}}$. By Proposition 2.23, $U_{1}(1)_{2}$ exists and $U_{1}(1)_{2}=W(1)_{2}-V(1)_{2}$. By item 1 of Proposition 2.24, $U_{2}(1)_{2}$ exists and $U_{2}(1)_{2}=U_{0}(1)_{2}\cdot U_{1}(1)_{2}$. By item 2 of Proposition 2.27, $\sum_{n\geq 0}|u_{0,n}|\leq\frac{1}{1-c}$. By item 2 of Proposition 2.24,
$\sum_{n\geq 0}|u_{2,n}|\leq\sum_{n\geq 0}|u_{0,n}|\cdot\sum_{n>k}|w_{n}-v_{n}|\leq\frac{1}{1-c}\sum_{n>k}|w_{n}-v_{n}|\,.$}
Since $U(x)^{2}=U_{2}(x)$, we have bound
$|t_{1}|+|t_{2}|+\dots\leq\frac{1}{|t|(1-c)}\sum_{n>k}|w_{n}-v_{n}|<1\,.$}
By Lemma 2.29, $U(1)_{2}$ exists and
$\big(U(1)_{2}\big)^{2}=U_{2}(1)_{2}=U_{0}(1)_{2}\cdot U_{1}(1)_{2}=1-\frac{V(1)_{2}}{W(1)_{2}}\,.$}
$\Box$
In conclusion, we consider $\xi$-convex light weights.
Proposition 2.31
Let $\xi\in\mathbb{C}$. For every $\xi$-convex light weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ we have
$\sum_{n=0}^{\infty}h(W(n))x^{n}=\frac{1}{1-\xi x}\,.$}
In particular, for $\xi=0$ we have $\sum_{n=0}^{\infty}h(W(n))x^{n}=1$.
Proof. Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be light and $\xi$-convex. We prove by induction on $n\in\mathbb{N}_{0}$ that $h(W(n))=\xi^{n}$. For $n=0$ this holds as $W(0)=\{\langle 1\rangle\}$. Let $n>0$. Recall that $\ell(w)$ denotes the last vertex of a walk $w$. Using Propositions 2.2 and 2.4, the $\xi$-convexity assumption and induction, we get
$h(W(n))$ $=$ ${\textstyle\sum_{w\in W(n)}h(w)}$
$=$ ${\textstyle\sum_{\begin{subarray}{c}w^{\prime}\in W(n-1)\\ \ell(w^{\prime})\in V(h)\end{subarray}}h(w^{\prime})\sum_{v\in\mathbb{N}\setminus\{\ell(w^{\prime})\}}h(\{\ell(w^{\prime}),\,v\})}$
$=$ ${\textstyle\sum_{w^{\prime}\in W(n-1)}h(w^{\prime})\cdot\xi=\xi^{n-1}\cdot\xi=\xi^{n}\,.}$
We can add and remove the condition $\ell(w^{\prime})\in V(h)$ without affecting the sum because walks $w^{\prime}$ not satisfying it have zero weight. $\Box$
3 Extensions of Pólya’s theorem with $v=1$
We begin with the case $v=1$ when the vertex to be visited by walks coincides with their starting vertex. In Theorems 3.23.7 we establish the first group of generalizations of Pólya’s theorem.
For that end, we need some generating functions. Let $n\in\mathbb{N}_{0}$, $v\in\mathbb{N}$, and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. Recall that $W(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, and that $W(v,n)\subset W(n)$ is the subset of walks visiting $v$ in some step $i>0$. We are interested in the generating function
$A_{h}(x)=\sum_{n\geq 0}a_{n}^{h}x^{n}=\sum_{n\geq 0}h(W(1,\,n))x^{n}$}
and define three more generating functions
$B_{h}(x)$ $=$ ${\textstyle\sum_{n\geq 0}b_{n}^{h}x^{n}=\sum_{n\geq 0}h(W_{b}(n))x^{n}\,,}$
$C_{h}(x)$ $=$ ${\textstyle\sum_{n\geq 0}c_{n}^{h}x^{n}=\sum_{n\geq 0}h(W_{c}(n))x^{n}\,\text{ and}}$
$D_{h}(x)$ $=$ ${\textstyle\sum_{n\geq 0}d_{n}^{h}x^{n}=\sum_{n\geq 0}h(W(n))x^{n}\,,}$
all in $\mathbb{C}[[x]]$. Here $W_{b}(n)$ are the walks of length $n$ in $K_{\mathbb{N}}$ that start and end at $1$, and $W_{c}(n)$ are the walks of length $n$ in $K_{\mathbb{N}}$ that start and end at $1$, and avoid $1$ between. The coefficients $a_{n}^{h}$, $\dots$, $d_{n}^{h}$ are correctly defined because they are sums of subseries of the absolutely convergent series
$d_{n}^{h}=\sum_{w\in W(n)}h(w)\,.$}
The constant coefficients are $a_{0}^{h}=c_{0}^{h}=0$ and $b_{0}^{h}=d_{0}^{h}=1$.
Proposition 3.1
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight and $A_{h}(x)$, $B_{h}(x)$, $C_{h}(x)$ and $D_{h}(x)$ be the above generating functions. The following relations hold.
  1. 1.
    $A_{h}(x)=C_{h}(x)D_{h}(x)$.
  2. 2.
    $B_{h}(x)=\frac{1}{1-C_{h}(x)}$, equivalently, $C_{h}(x)=1-\frac{1}{B_{h}(x)}$.
Proof. 1. We show that $a_{n}^{h}=\sum_{j=0}^{n}c_{j}^{h}\cdot d_{n-j}^{h}$ for every $n\in\mathbb{N}_{0}$. For $n\leq 1$ it holds, then $a_{n}^{h}=c_{n}^{h}=0$. For $n\geq 2$ the set $W(1,n)$ is countable and we split every walk in it at the first revisit of $1$. We get a map
$F\colon W(1,\,n)\to\bigcup_{j=0}^{n}W_{c}(j)\times W(n-j)$}
that is defined as follows. For any walk
$w=\langle u_{0},\,u_{1},\,\dots,\,u_{j},\,\dots,\,u_{n}\rangle\in W(1,\,n)\,$,
where $u_{0}=u_{j}=1$, $j>0$ and $u_{i}\neq 1$ for $i\in[j-1]$, we set
$F(w)=\langle w_{1},\,w_{2}\rangle\ \ (\in W_{c}(j)\times W(n-j))\,$,
with $w_{1}=\langle u_{0},u_{1},\dots,u_{j}\rangle$ and $w_{2}=\langle u_{j},u_{j+1},\dots,u_{n}\rangle$. It is easy to see that $F$ is a bijection. $F$ is also weight-preserving (WP): for every weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and value $F(w)=\langle w_{1},w_{2}\rangle$, we have $h(w)=h(w_{1})h(w_{2})$. If $X$ is a set of pairs of walks in $K_{\mathbb{N}}$, we write $h(X)$ for the sum of the series
$\sum_{\langle w,\,w^{\prime}\rangle\in X}h(w)h(w^{\prime})\,$,}
if it absolutely converges.
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight and let $n\geq 2$. Using the map $F$, we get
$a_{n}^{h}$ $=$ $h(W(1,\,n))\stackrel{{\scriptstyle\mathrm{WP}}}{{=}}h(F[W(1,\,n)])$
[…]
2. We show that
$B_{h}(x)=\frac{1}{1-C_{h}(x)}=1+\sum_{j\geq 1}(C_{h}(x))^{j}\,.$}
We recall that $b_{0}^{h}=1$ and $c_{0}^{h}=0$ and show that for every $n\in\mathbb{N}$,
$b_{n}^{h}=\sum_{j=1}^{\infty}\sum_{\begin{subarray}{c}n_{1},\,\dots,\,n_{j}\in\mathbb{N}\\ n_{1}+\dots+n_{j}=n\end{subarray}}c_{n_{1}}^{h}c_{n_{2}}^{h}\dots c_{n_{j}}^{h}\,.$}
For odd $n$ it is true: for odd $n$ we have $b^{h}_{n}=c^{h}_{n}=0$, and in every inner summand at least one index $n_{i}$ is odd. For even $n\geq 2$ the set $W_{b}(n)$ is countable. We split every walk in it at the visits of $1$ and get a map
$F\colon W_{b}(n)\to\bigcup_{j=1}^{\infty}\bigcup_{\begin{subarray}{c}n_{1},\,\dots,\,n_{j}\in\mathbb{N}\\ n_{1}+\dots+n_{j}=n\end{subarray}}W_{c}(n_{1})\times\cdots\times W_{c}(n_{j})$}
that is defined as follows. For any walk
$w=\langle u_{0},\,\dots,\,u_{m_{1}},\,\dots,\,\dots,\,u_{m_{j-1}},\,\dots,\,u_{n}\rangle\in W_{b}(n)\,$,
where $0=m_{0}<m_{1}<\dots<m_{j}=n$ and $u_{k}=1$ iff $k=0$ or $k=m_{i}$ for some $i\in[j]$, we set
$F(w)=\langle w_{1},\,w_{2},\,\dots,\,w_{j}\rangle\ \ (\in W_{c}(n_{1})\times W_{c}(n_{2})\times\dots\times W_{c}(n_{j}),\,n_{i}=m_{i}-m_{i-1})\,$,
with $w_{i}=\langle u_{m_{i-1}},u_{m_{i-1}+1},\dots,u_{m_{i}}\rangle$, $i\in[j]$. It is easy to see that $F$ is a bijection and that it is weight-preserving, as in part 1. If $X$ is a set of tuples of walks, we write $h(X)$ for the sum of the series
$\sum_{\langle w_{1},\,\dots,\,w_{k})\in X}h(w_{1})\dots h(w_{k})\,$,}
if it absolutely converges.
Let $h$ be a light weight and $n\geq 2$ be even. Using the map $F$, we get
$b_{n}^{h}$ $=$ $h(W_{b}(n))\stackrel{{\scriptstyle\mathrm{WP}}}{{=}}h(F[W_{b}(n)])$
[…]
$\Box$
We label our fourteen generalizations of Pólya’s theorem for better orientation by the triples
$\langle x,\,y,\,z\rangle\in\{v=1,\,v\neq 1\}\times\{\mathrm{gen},\,\mathrm{con}\}\times\{U(1)_{1},\,U(1)_{2},\,U(1)_{1}=\infty\}\,$,
where “gen” refers to general edge weights $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and “con” to $\xi$-convex ones. For $U(1)_{1}$ and $U(1)_{2}$ see Definition 2.12, and for $U(1)_{1}=\infty$ see Definition 2.14. The generating functions $A_{h}(x)$, $B_{h}(x)$, $C_{h}(x)$, and $D_{h}(x)$ are defined at the beginning of this section.
Theorem 3.2 (1: $v=1$, gen, $U(1)_{1}$)
If $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a light weight such that $A_{h}(1)_{1}$, $B_{h}(1)_{1}$, $C_{h}(1)_{1}$ and $D_{h}(1)_{1}$ exist, then $B_{h}(1)_{1}\neq 0$ and
$A_{h}(1)_{1}=\big(1-\frac{1}{B_{h}(1)_{1}}\big)D_{h}(1)_{1}\,.$}
Proof. By item 1 of Proposition 3.1 we have $A_{h}(x)=C_{h}(x)D_{h}(x)$. By Proposition 2.20, $A_{h}(1)_{1}=C_{h}(1)_{1}D_{h}(1)_{1}$. By item 2 of Proposition 3.1 we have $1-C_{h}(x)=\frac{1}{B_{h}(x)}$. By Proposition 2.21 we have $B_{h}(1)_{1}\neq 0$ and
$(1-C_{h})(1)_{1}=\frac{1}{B_{h}(1)_{1}},\,\text{ so that }\,C_{h}(1)_{1}=1-\frac{1}{B_{h}(1)_{1}}\,.$}
The stated formula follows. $\Box$
Theorem 3.3 (2: $v=1$, gen, $U(1)_{2}$)
If $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a light weight such that $\sum_{n\geq 1}|b_{n}^{h}|<1$ and $D_{h}(1)_{2}$ exists, then $B_{h}(1)_{2}\neq 0$, and
$A_{h}(1)_{2}=\big(1-\frac{1}{B_{h}(1)_{2}}\big)D_{h}(1)_{2}\,.$}
Proof. By item 2 of Proposition 3.1 we have $1-C_{h}(x)=\frac{1}{B_{h}(x)}$. By item 1 of Proposition 2.27, $B_{h}(1)_{2}\neq 0$ and
$(1-C_{h}(1))_{2}=\frac{1}{B_{h}(1)_{2}},\,\text{ so that }\,C_{h}(1)_{2}=1-\frac{1}{B_{h}(1)_{2}}\,.$}
By item 1 of Proposition 3.1 we have $A_{h}(x)=C_{h}(x)D_{h}(x)$. By item 1 of Proposition 2.24, $A_{h}(1)_{2}=C_{h}(1)_{2}D_{h}(1)_{2}$ . The stated formula follows. $\Box$
Theorem 3.4 (3: $v=1$, gen, $U(1)_{1}=\infty$)
If $\xi\in\mathbb{C}_{1}$ and $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ is a light weight such that $(D_{h})_{1/\xi}(1)_{1}=+\infty$, then
$(A_{h})_{1/\xi}(1)_{1}=+\infty\,.$
Proof. From $(D_{h})_{1/\xi}(1)_{1}=+\infty$ it follows that $\xi^{-n}d_{n}^{h}>0$ for some $n\in\mathbb{N}$. Hence there is a vertex $u\in\mathbb{N}\setminus\{1\}$ with $\xi^{-1}h(\{1,u\})>0$. The walk $\langle 1,u,1\rangle$ shows by the subseries bound principle that $\xi^{-2}c_{2}^{h}>0$. Since by item 1 of Proposition 3.1 we have $a_{n}^{h}=\sum_{j=0}^{n}c_{j}^{h}d_{n-j}^{h}$, we get that
$\xi^{-2-n}a_{n+2}^{h}\geq\xi^{-2}c_{2}\cdot\xi^{-n}d_{n}^{h}$
for every $n\in\mathbb{N}_{0}$. Hence
$\sum_{j=0}^{n}\xi^{-2-j}a_{j+2}^{h}\geq\xi^{-2}c_{2}^{h}\cdot\sum_{j=0}^{n}\xi^{-j}d_{j}^{h}\to+\infty$}
for $n\to\infty$, and $(A_{h})_{1/\xi}(1)_{1}=+\infty$. $\Box$
We proceed to $\xi$-convex weights.
Theorem 3.5 (4: $v=1$, con, $U(1)_{1}$)
1. If $\xi\in\mathbb{C}_{1}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $\xi$-convex light weight and if $(B_{h})_{1/\xi}(1)_{1}$ and $(C_{h})_{1/\xi}(1)_{1}$ exist, then $(B_{h})_{1/\xi}(1)_{1}\neq 0$ and
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h}=1-\frac{1}{(B_{h})_{1/\xi}(1)_{1}}\,.$
2. If $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a 0-convex light weight, and if $B_{h}(1)_{1}$ and $C_{h}(1)_{1}$ exist, then $B_{h}(1)_{1}\neq 0$ and
$A_{h}(1)_{1}=1-\frac{1}{B_{h}(1)_{1}}\,.$
Proof. 1. By item 1 of Proposition 3.1 and by Proposition 2.31 we have $A_{h}(x)=C_{h}(x)D_{h}(x)=C_{h}(x)\frac{1}{1-\xi x}$. By Proposition 2.15,
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h}=(C_{h})_{1/\xi}(1)_{1}\,.$
By item 2 of Proposition 3.1 we have $1-C_{h}(x)=\frac{1}{B_{h}(x)}$. By Proposition 2.21 we have $(B_{h})_{1/\xi}(1)_{1}\neq 0$ and
[…]
The stated formula follows.
2. Now, by item 1 of Proposition 3.1 and by Proposition 2.31, we have
$A_{h}(x)=C_{h}(x)$
and argue as in the previous case for $\xi=1$. $\Box$
This theorem generalizes the case $d\geq 3$ of Pólya’s theorem with $\overline{v}=\overline{0}$. The case $d\leq 2$ is generalized in Theorem 3.7.
Theorem 3.6 (5: $v=1$, con, $U(1)_{2}$)
1. If $\xi\in\mathbb{C}_{1}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $\xi$-convex light weight and $\sum_{n\geq 1}\big|b_{n}^{h}\big|<1$, then $(B_{h})_{1/\xi}(1)_{2}\neq 0$ and
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h}=1-\frac{1}{(B_{h})_{1/\xi}(1)_{2}}\,.$
2. If $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a 0-convex light weight and $\sum_{n\geq 1}\big|b_{n}^{h}\big|<1$, then $B_{h}(1)_{2}\neq 0$ and
$A_{h}(1)_{2}=1-\frac{1}{B_{h}(1)_{2}}\,.$
Proof. 1. By item 2 of Proposition 3.1 we have $1-C_{h}(x)=\frac{1}{B_{h}(x)}$. By item 1 of Proposition 2.27, $(B_{h})_{1/\xi}(1)_{2}\neq 0$ and
[…]
By item 1 of Proposition 3.1 and Proposition 2.31, $A_{h}(x)=C_{h}(x)D_{h}(x)=C_{h}(x)\frac{1}{1-\xi x}$. By Proposition 2.15
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h}=(C_{h})_{1/\xi}(1)_{1}=(C_{h})_{1/\xi}(1)_{2}=1-\frac{1}{(B_{h})_{1/\xi}(1)_{2}}\,.$}
2. In the previous case we have for $\xi=1$ that $C_{h}(1)_{2}=1-\frac{1}{B_{h}(1)_{2}}$. By item 1 of Proposition 3.1 and by Proposition 2.31, $A_{h}(x)=C_{h}(x)$. The stated formula follows. $\Box$
Theorem 3.7 (6: $v=1$, con, $U(1)_{1}=\infty$)
1. If $\xi\in\mathbb{C}_{1}$, $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ is a $\xi$-convex light weight and if $(B_{h})_{1/\xi}(1)_{1}=+\infty$, then
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h}=1\,.$
2. If $\xi\in\mathbb{C}_{1}$, $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}$ is a 0-convex light weight, $B_{h},C_{h}\in H_{1}$, and if $(B_{h})_{1/\xi}(1)_{1}=\pm\infty$, then
$\lim_{x\nearrow 1}F_{(A_{h})_{1/\xi}}(x)=1\,.$
Proof. 1. By item 1 of Proposition 3.1 and by Proposition 2.31 we have
$A_{h}(x)=C_{h}(x)D_{h}(x)=C_{h}(x)\frac{1}{1-\xi x}\,.$}
By Proposition 2.15,
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h}=(C_{h})_{1/\xi}(1)_{1}\,.$
By Proposition 2.31 and the subseries bound principle,
$0\leq\xi^{-n}b_{n}^{h},\,\xi^{-n}c_{n}^{h}\leq\xi^{-n}d_{n}^{h}=1$
for every $n\in\mathbb{N}_{0}$. Hence $B_{h},C_{h}\in H_{1}$. By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$. Thus
$(C_{h})_{1/\xi}(1)_{1}=\lim_{x\nearrow 1}F_{(C_{h})_{1/\xi}}(x)={\textstyle 1-\frac{1}{\lim_{x\nearrow 1}F_{(B_{h})_{1/\xi}}(x)}=1-\frac{1}{(B_{h})_{1/\xi}(1)_{1}}=1\,,}$
where in the first and third equality we used Theorem 2.18. The stated formula follows.
2. By item 1 of Proposition 3.1 and by Proposition 2.31, we have $A_{h}(x)=C_{h}(x)$. By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$. Thus
$\lim_{x\nearrow 1}F_{(C_{h})_{1/\xi}}(x)={\textstyle 1-\frac{1}{\lim_{x\nearrow 1}F_{(B_{h})_{1/\xi}}(x)}=1-\frac{1}{(B_{h})_{1/\xi}(1)_{1}}=1-\frac{1}{\pm\infty}=1\,,}$
where in the second equality we used Theorem 2.17. The stated formula follows at once. $\Box$
4 Extensions of Pólya’s theorem with $v\neq 1$
In Theorems 4.34.10 we establish the second group of generalizations of Pólya’s theorem, in the more complicated case when the vertex to be visited by walks differs from their starting vertex. The grid graph $\mathbb{Z}^{d}$ is vertex-transitive, which has so far been irrelevant. For $\overline{v}\neq\overline{0}$, vertex-transitivity becomes relevant. We generalize it to $K_{\mathbb{N}}$ as follows. Let $v\in\mathbb{N}\setminus\{1\}$. We say that a weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is $v$-transitive if there is a bijection $f\colon\mathbb{N}\to\mathbb{N}$ such that $f(1)=v$ and $h(f[e])=h(e)$ for every $e\in\mathbb{N}_{2}$.
We again introduce some generating functions. Let $n\in\mathbb{N}_{0}$, $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. Recall that $W(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, and that $W(v,n)\subset W(n)$ is the subset of walks visiting the vertex $v$ ($\neq 1$); necessarily in a step $i>0$. We are interested in the generating function
$A_{h,v}(x)=\sum_{n\geq 0}a_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W(v,\,n))x^{n}\,.$}
In the previous section we introduced the generating functions $B_{h}(x)$, $C_{h}(x)$ and $D_{h}(x)$. Now we define
$B_{h,v}(x)$ $=$ ${\textstyle\sum_{n\geq 0}b_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W_{b}(v,\,n))x^{n}\,,}$
$C_{h,v}(x)$ $=$ ${\textstyle\sum_{n\geq 0}c_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W_{c}(v,\,n))x^{n}\,\text{ and}}$
$E_{h,v}(x)$ $=$ ${\textstyle\sum_{n\geq 0}e_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W_{e}(v,\,n))x^{n}\,,}$
all in $\mathbb{C}[[x]]$. Here $W_{b}(v,n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting and ending at $1$ and avoiding $v$. $W_{c}(v,n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, ending at $v$ and between avoiding $v$. Finally, $W_{e}(v,n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, ending at $v$ and between avoiding both $1$ and $v$. The constant coefficients are $a_{0}^{h,v}=c_{0}^{h,v}=c_{0}^{h}=e_{0}^{h,v}=0$ and $b_{0}^{h}=b_{0}^{h,v}=d_{0}^{h}=1$.
We obtain an analog of Proposition 3.1. Since the proof is similar, we omit details and argue only on the level of semiformal method.
Proposition 4.1
Let $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight and $A_{h,v}(x)$, $B_{h}(x)$, $B_{h,v}(x)$, $C_{h,v}(x)$, $D_{h}(x)$ and $E_{h,v}(x)$ be the above generating functions. The following relations hold.
  1. 1.
    $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)$.
  2. 2.
    $B_{h}(x)=B_{h,v}(x)+C_{h,v}(x)^{2}B_{h}(x)$, equivalently, $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$.
  3. 3.
    $C_{h,v}(x)=B_{h,v}(x)E_{h,v}(x)$.
Proof. 1. Walks $w$ weight-counted by $A_{h,v}(x)$ split at the first visit of $v$ as $w=w_{1}w_{2}$. Walks $w_{1}$ are weight-counted by $C_{h,v}(x)$ and $w_{2}$ are arbitrary walks starting at $v$. By the $v$-transitivity of $h$, walks $w_{2}$ are weight-counted by $D_{h}(x)$.
2. It suffices to prove the first equality. $B_{h}(x)$ weight-counts walks $w$ starting and ending at $1$. Those avoiding $v$ are weight-counted by $B_{h,v}(x)$. If $w$ visits $v$, it uniquely splits at the first and last visits of $v$ as $w=w_{1}w_{2}w_{3}$. Walks $w_{1}$ start at $1$, end at $v$ and between avoid $v$. Walks $w_{2}$ start and end at $v$. Walks $w_{3}$ start at $v$, end at $1$ and between avoid $v$. We reverse walks $w_{3}$ and see that both $w_{1}$ and $w_{3}$ are weight-counted by $C_{h,v}(x)$. Due to the $v$-transitivity of $h$, the middle walks $w_{2}$ are weight-counted by $B_{h}(x)$.
3. We consider walks $w$ weight-counted by $C_{h,v}(x)$. They start at $1$, end at $v$, and between avoid $v$. These walks uniquely split at the last visit of $1$ as $w=w_{1}w_{2}$. Walks $w_{1}$ are weight-counted by $B_{h,v}(x)$, and $w_{2}$ by $E_{h,v}(x)$. $\Box$
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and $v\in\mathbb{N}\setminus\{1\}$. If $v\not\in V(h)$ then $a_{n}^{h,v}=\sum_{w\in W(v,n)}h(w)=\sum_{w\in W(v,n)}0=0$ for every $n\in\mathbb{N}_{0}$ because every walk starting at $1$ and visiting $v$ contains an edge with zero weight. Henceforth, we therefore assume that $v\in V(h)$. For $w,z\in\mathbb{C}$ we define the set
$w\cdot\mathrm{sqrt}(z)=\{w\alpha\colon\;\alpha\in\mathbb{C},\,\alpha^{2}=z\}\ \ (\subset\mathbb{C})\,.$
For $w=0$ or $z=0$ it equals $\{0\}$. Else it has two non-zero elements differing only by sign. If $w=1$, we write just $\mathrm{sqrt}(z)$.
Before we turn to generalizations of Pólya’s theorem in the case $v\neq 1$, we utilize the square in item 2 of Proposition 4.1. If $h\colon X\to\mathbb{R}$ is an absolutely convergent real series and $Y\subset X$, then in general the sum $\sum_{x\in Y}h(x)$ of the subseries need not be smaller than the sum $\sum_{x\in X}h(x)$. However, due to the mentioned square, we obtain the next monotonicity result for the sums
$b_{n}^{h,v}=\sum_{w\in W_{b}(v,\,n)}h(w)\,\text{ and }\,b_{n}^{h}=\sum_{w\in W_{b}(n)}h(w)\,.$}
It is expressed in terms of the generating functions $B_{h,v}(x)=\sum_{n\geq 0}b_{n}^{h,v}x^{n}$ and $B_{h}(x)=\sum_{n\geq 0}b_{n}^{h}x^{n}$.
Theorem 4.2
Let $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{R}$ be a $v$-transitive light weight and let $B_{h,v},B_{h},C_{h,v}\in H_{1}$. Then for every real $c\in(-1,1)$ the following holds.
  1. 1.
    If $B_{h}(c)=0$ then also $B_{h,v}(c)=0$.
  2. 2.
    If $B_{h}(c)\neq 0$ then
    $\frac{B_{h,v}(c)}{B_{h}(c)}\leq 1\,.$
Proof. This follows from relations 2 in Proposition 4.1. $\Box$
We proceed to the remaining generalizations of Pólya’s theorem.
Theorem 4.3 (7: $v\neq 1$, gen, $U(1)_{1}$)
If $v$ is in $\mathbb{N}\setminus\{1\}$, if $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive light weight with $v\in V(h)$, and if $A_{h,v}(1)_{1}$, $B_{h}(1)_{1}\neq 0$, $B_{h,v}(1)_{1}$, $C_{h,v}(1)_{1}$ and $D_{h}(1)_{1}$ exist, then
$A_{h,\,v}(1)_{1}\in{\textstyle D_{h}(1)_{1}\cdot\mathrm{sqrt}\big(1-\frac{B_{h,\,v}(1)_{1}}{B_{h}(1)_{1}}}\big)\,.$
Proof. By item 1 of Proposition 4.1 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)$. By Proposition 2.20, $A_{h,v}(1)_{1}=C_{h,v}(1)_{1}D_{h}(1)_{1}$. By item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$. Thus by Proposition 2.22, $\big(C_{h,v}(1)_{1}\big)^{2}=1-\frac{B_{h,v}(1)_{1}}{B_{h}(1)_{1}}$. The stated formula follows. $\Box$
If we replace the assumption of non-vanishing of $B_{h}(1)_{1}$ with the existence of $E_{h,v}(1)_{1}$, we obtain the following variant of the theorem.
Theorem 4.4 (7: $v\neq 1$, gen, $U(1)_{1}$)
If $v$ is in $\mathbb{N}\setminus\{1\}$, if $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive light weight with $v\in V(h)$, and if $A_{h,v}(1)_{1}$, $B_{h}(1)_{1}$, $B_{h,v}(1)_{1}$, $C_{h,v}(1)_{1}$, $D_{h}(1)_{1}$ and $E_{h,v}(1)_{1}$ exist, then
$A_{h,\,v}(1)_{1}\in{\textstyle D_{h}(1)_{1}\cdot\mathrm{sqrt}\big(1-\frac{B_{h,\,v}(1)_{1}}{B_{h}(1)_{1}}}\big)\,$,
where for $B_{h}(1)_{1}=0$ we have $B_{h,v}(1)_{1}=0$ and interpret the expression $\frac{0}{0}$ as $1$, so that then $A_{h,v}(1)_{1}=0$.
Proof. If $B_{h}(1)_{1}\neq 0$, we use the previous proof. If $B_{h}(1)_{1}=0$ then it follows from item 2 of Proposition 4.1, and Propositions 2.20 and 2.19 that also $B_{h,v}(1)_{1}=0$. Item 3 of Proposition 4.1 and Proposition 2.20 give $C_{h,v}(1)_{1}=0$. Finally, item 1 of Proposition 4.1 and Proposition 2.20 give $A_{h,v}(1)_{1}=0$. $\Box$
Theorem 4.5 (8: $v\neq 1$, gen, $U(1)_{2}$)
Suppose that $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive light weight with $v\in V(h)$,
$c=\sum_{n\geq 1}|b_{n}^{h}|<1$}
and that $B_{h,v}(1)_{2}$ and $D_{h}(1)_{2}$ exist. If $B_{h,v}(x)=B_{h}(x)$ then $A_{h,v}(x)=0$. For $B_{h,v}(x)\neq B_{h}(x)$ let $k\in\mathbb{N}_{0}$ be minimum such that $b_{k}^{h,v}\neq b_{k}^{h}$. If
$\big(|b_{k}^{h,v}-b_{k}^{h}|\cdot(1-c)\big)^{-1}\sum_{n>k}|b_{n}^{h,v}-b_{n}^{h}|<1\,$,}
then $B_{h}(1)_{2}\neq 0$ and
$A_{h,\,v}(1)_{2}\in{\textstyle D_{h}(1)_{2}\cdot\mathrm{sqrt}\big(1-\frac{B_{h,\,v}(1)_{2}}{B_{h}(1)_{2}}}\big)\,.$
Proof. If $B_{h,v}(x)=B_{h}(x)$, by item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}=1-1=0$ and $C_{h,v}(x)=0$. Item 1 of Proposition 4.1 gives $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=0$.
Let $B_{h,v}(x)\neq B_{h}(x)$, $k$ be as stated and let the displayed bound hold. By item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$. By Proposition 2.30, $B_{h}(1)_{2}\neq 0$ and
$(C_{h,v}(1)_{2})^{2}=1-\frac{B_{h,v}(1)_{2}}{B_{h}(1)_{2}}\,.$}
By item 1 of Proposition 4.1 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)$. By item 1 of Proposition 2.24, $A_{h,v}(1)_{2}=C_{h,v}(1)_{2}D_{h}(1)_{2}$. The stated formula follows. $\Box$
Theorem 4.6 (9: $v\neq 1$, gen, $U(1)_{1}=\infty$)
If $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ is a $v$-transitive light weight with $v\in V(h)$ and if $(D_{h})_{1/\xi}(1)_{1}=+\infty$, then
$(A_{h,v})_{1/\xi}(1)_{1}=+\infty\,.$
Proof. We denote the minimum length of a walk joining $1$ and $v$ in $G(h)$ by $m$. The inner vertices of it differ from $v$. By the subseries bound principle we get $\xi^{-m}c^{h,v}_{m}>0$. By item 1 of Proposition 4.1 we have $a_{n}^{h,v}=\sum_{j=0}^{n}c_{j}^{h,v}d_{n-j}^{h}$ and we conclude that $\xi^{-n-m}a_{n+m}^{h,v}\geq\xi^{-m}c_{m}^{h,v}\cdot\xi^{-n}d_{n}^{h}$ for every $n\in\mathbb{N}_{0}$. Hence
$\sum_{j=0}^{n}\xi^{-j-m}a_{j+m}^{h,v}\geq\xi^{-m}c_{m}^{h,v}\cdot\sum_{j=0}^{n}\xi^{-j}d_{j}^{h}\to+\infty$}
for $n\to\infty$, and $(A_{h,v})_{1/\xi}(1)_{1}=+\infty$. $\Box$
We proceed to $\xi$-convex weights.
Theorem 4.7 (10: $v\neq 1$, con, $U(1)_{1}$)
1. If $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive $\xi$-convex light weight with $v\in V(h)$, and if $(B_{h,v})_{1/\xi}(1)_{1}$, $(B_{h})_{1/\xi}(1)_{1}\neq 0$ and $(C_{h,v})_{1/\xi}(1)_{1}$ exist, then
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}\in\mathrm{sqrt}\big(1-\frac{(B_{h,v})_{1/\xi}(1)_{1}}{(B_{h})_{1/\xi}(1)_{1}}\big)\,.$}
2. If $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive 0-convex light weight with $v\in V(h)$, and if $B_{h,v}(1)_{1}$, $B_{h}(1)_{1}\neq 0$ and $C_{h,v}(1)_{1}$ exist, then
$A_{h,v}(1)_{1}\in\mathrm{sqrt}\big(1-\frac{B_{h,v}(1)_{1}}{B_{h}(1)_{1}}\big)\,.$}
Proof. 1. By item 1 of Proposition 4.1 and by Proposition 2.31 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=C_{h,v}(x)\frac{1}{1-\xi x}$. By Proposition 2.15,
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}=(C_{h,v})_{1/\xi}(1)_{1}\,.$
By item 2 of Proposition 4.1 we have $(C_{h,v})_{1/\xi}(x)^{2}=1-\frac{(B_{h,v})_{1/\xi}(x)}{(B_{h})_{1/\xi}(x)}$. By Proposition 2.22, $\big((C_{h,v})_{1/\xi}(1)_{1}\big)^{2}=1-\frac{(B_{h,v})_{1/\xi}(1)_{1}}{(B_{h})_{1/\xi}(1)_{1}}$. The stated formula follows.
2. Now, by item 1 of Proposition 4.1 and by Proposition 2.31, we have
$A_{h,v}(x)=C_{h,v}(x)\,.$
The rest of the argument is the same as in the previous case for $\xi=1$. $\Box$
Again, if we replace the assumption of non-vanishing of $B_{h,v}(1)_{1}$ with the existence of $E_{h,v}(1)_{1}$, we obtain the following variant of the theorem.
Theorem 4.8 (10: $v\neq 1$, con, $U(1)_{1}$)
1. If $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive $\xi$-convex light weight with $v\in V(h)$, and if $(B_{h,v})_{1/\xi}(1)_{1}$, $(B_{h})_{1/\xi}(1)_{1}$, $(C_{h,v})_{1/\xi}(1)_{1}$ and $(E_{h,v})_{1/\xi}(1)_{1}$ exist, then
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}\in\mathrm{sqrt}\big(1-\frac{(B_{h,v})_{1/\xi}(1)_{1}}{(B_{h})_{1/\xi}(1)_{1}}\big)\,$,}
where for $(B_{h})_{1/\xi}(1)_{1}=0$ we have $(B_{h,v})_{1/\xi}(1)_{1}=0$ and interpret the expression $\frac{0}{0}$ as $1$, so that then the limit is zero.
2. If $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive 0-convex light weight with $v\in V(h)$, and if $B_{h,v}(1)_{1}$, $B_{h}(1)_{1}$, $C_{h,v}(1)_{1}$ and $E_{h,v}(1)_{1}$ exist, then
$A_{h,v}(1)_{1}\in\mathrm{sqrt}\big(1-\frac{B_{h,v}(1)_{1}}{B_{h}(1)_{1}}\big)\,$,}
where for $B_{h}(1)_{1}=0$ we have $B_{h,v}(1)_{1}=0$ and interpret the $\frac{0}{0}$ as $1$, so that then $A_{h,v}(1)_{1}=0$.
Proof. 1. If $(B_{h})_{1/\xi}(1)_{1}\neq 0$, we use the previous proof. If $(B_{h})_{1/\xi}(1)_{1}=0$ then it follows from item 2 of Proposition 4.1 and Propositions 2.20 and 2.19 that also $(B_{h,v})_{1/\xi}(1)_{1}=0$. Item 3 of Proposition 4.1 and Proposition 2.20 give $(C_{h,v})_{\xi}(1)_{1}=0$. Thus, as in the previous theorem,
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}=(C_{h,v})_{1/\xi}(1)_{1}=0\,.$
2. Now, by item 1 of Proposition 4.1 and by Proposition 2.31, we have
$A_{h,v}(x)=C_{h,v}(x)\,.$
The rest of the argument is the same as in the previous case for $\xi=1$. $\Box$
These two theorems generalize the case $d\geq 3$ of Pólya’s theorem with $\overline{v}\neq\overline{0}$. The case $d\leq 2$ is generalized in Theorem 4.10.
Theorem 4.9 (11: $v\neq 1$, con, $U(1)_{2}$)
1. Suppose that $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive $\xi$-convex light weight with $v\in V(h)$,
$c=\sum_{n\geq 1}|b_{n}^{h}|<1$}
and that $(B_{h,v})_{1/\xi}(1)_{2}$ exists. If $B_{h,v}(x)=B_{h}(x)$ then $A_{h,v}(x)=0$. For $B_{h,v}(x)\neq B_{h}(x)$ let $k\in\mathbb{N}_{0}$ be minimum such that $b_{k}^{h,v}\neq b_{k}^{h}$. If
$\big(|b_{k}^{h,v}-b_{k}^{h}|\cdot(1-c)\big)^{-1}\sum_{n>k}|b_{n}^{h,v}-b_{n}^{h}|<1\,$,}
then $(B_{h})_{1/\xi}(1)_{2}\neq 0$ and
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}\in{\textstyle\mathrm{sqrt}\big(1-\frac{(B_{h,v})_{1/\xi}(1)_{2}}{(B_{h})_{1/\xi}(1)_{2}}}\big)\,.$
2. Suppose that $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive 0-convex light weight with $v\in V(h)$,
$c=\sum_{n\geq 1}|b_{n}^{h}|<1$}
and that $B_{h,v}(1)_{2}$ exists. If $B_{h,v}(x)=B_{h}(x)$ then $A_{h,v}(x)=0$. For $B_{h,v}(x)\neq B_{h}(x)$ let $k\in\mathbb{N}_{0}$ be minimum such that $b_{k}^{h,v}\neq b_{k}^{h}$. If
$\big(|b_{k}^{h,v}-b_{k}^{h}|\cdot(1-c)\big)^{-1}\sum_{n>k}|b_{n}^{h,v}-b_{n}^{h}|<1\,$,}
then $(B_{h})(1)_{2}\neq 0$ and
$A_{h,v}(1)_{2}\in{\textstyle\mathrm{sqrt}\big(1-\frac{B_{h,v}(1)_{2}}{B_{h}(1)_{2}}}\big)\,.$
Proof. If $B_{h,v}(x)=B_{h}(x)$, by item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}=0$ and $C_{h,v}(x)=0$. Item 1 of Proposition 4.1 gives $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=0$.
1. Let $h$ be $\xi$-converx with $\xi\in\mathbb{C}_{1}$, $B_{h,v}(x)\neq B_{h}(x)$, $k$ be as stated and let the displayed bound hold. By item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$. By Proposition 2.30, $(B_{h})_{1/\xi}(1)_{2}\neq 0$ and
$\big((C_{h,v})_{1/\xi}(1)_{2}\big)^{2}=1-\frac{(B_{h,v})_{1/\xi}(1)_{2}}{(B_{h})_{1/\xi}(1)_{2}}\,.$}
By item 1 of Proposition 4.1 and by Proposition 2.31 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=C_{h,v}(x)\frac{1}{1-\xi x}$. Proposition 2.15 gives the stated formula:
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}={\textstyle(C_{h,\,v})_{1/\xi}(1)_{1}=(C_{h,\,v})_{1/\xi}(1)_{2}\in\mathrm{sqrt}\big(1-\frac{(B_{h,v})_{1/\xi}(1)_{2}}{(B_{h})_{1/\xi}(1)_{2}}\big)\,.}$
2. Let $h$ be 0-convex, $B_{h,v}(x)\neq B_{h}(x)$, $k$ be as stated and let the displayed bound hold. The beginning of the previous case gives for $\xi=1$ that $\big(C_{h,v}(1)_{2}\big)^{2}=1-\frac{B_{h,v}(1)_{2}}{B_{h}(1)_{2}}$. Since by item 1 of Proposition 4.1 and by Proposition 2.31 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=C_{h,v}(x)$, the stated formula follows. $\Box$
Theorem 4.10 (12: $v\neq 1$, con, $U(1)_{1}=\infty$)
1. If $\xi\in\mathbb{C}_{1}$, $v$ is in $\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ is a $v$-transitive $\xi$-convex light weight with $v\in V(h)$, and if $(B_{h})_{1/\xi}(1)_{1}=+\infty$, then
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}=1\,.$
2. If $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}$ is a $v$-transitive 0-convex light weight with $v\in V(h)$, $B_{h},C_{h,v}\in H_{1}$, $(B_{h,v})_{1/\xi}(1)_{1}$ exists and if $(B_{h})_{1/\xi}(1)_{1}=\pm\infty$, then
$\lim_{x\nearrow 1}F_{(A_{h,v})_{1/\xi}}(x)\in\{-1,\,1\}\,.$
Proof. 1. By Proposition 2.31, $D_{h}(x)=\frac{1}{1-\xi x}$. By the subseries bound principle we have
$0\leq\xi^{-n}b_{n}^{h},\,\xi^{-n}b_{n}^{h,v},\,\xi^{-n}c_{n}^{h,v},\,\xi^{-n}e_{n}^{h,v}\leq\xi^{-n}d_{n}^{h}=1$
for every $n\in\mathbb{N}_{0}$. Hence $B_{h}(x)$, $B_{h,v}(x)$, $C_{h,v}(x)$ and $E_{h,v}(x)$ are in $H_{1}$. Any shortest walk joining $1$ and $v$ in $G(h)$ shows that for some $m\in\mathbb{N}$ we have $\xi^{-m}e_{m}^{h,v}>0$. Thus the function $F_{(E_{h,v})_{1/\xi}}(x)$ is positive and increasing on $(0,1)$. Since $F_{(C_{h,v})_{1/\xi}}(x)$ is bounded on $[0,1)$ by item 2 of Proposition 4.1, $F_{(B_{h,v})_{1/\xi}}(x)$ is bounded by item 3. Since by item 1 of Proposition 4.1 and by Proposition 2.31 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=C_{h,v}(x)\frac{1}{1-\xi x}$, Proposition 2.15 gives
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}=(C_{h,v})_{1/\xi}(1)_{1}=(C_{h,v})_{1/\xi}(1)_{2}\,.$
By item 2 of Proposition 4.1, Propositions 2.2, 2.4 and 2.5, and by Theorem 2.18 we have
$(C_{h,v}^{2})_{1/\xi}(1)_{1}=\lim_{x\nearrow 1}F_{(C_{h,v}^{2})_{1/\xi}}(x)=1-\frac{\lim_{x\nearrow 1}F_{(B_{h,v})_{1/\xi}}(x)}{\lim_{x\nearrow 1}F_{(B_{h})_{1/\xi}}(x)}=1\,$,}
because $F_{(B_{h,v})_{1/\xi}}(x)$ is bounded on $[0,1)$ and $\lim_{x\nearrow 1}F_{(B_{h})_{1/\xi}}(x)=B_{h}(1)_{1}=+\infty$. By Proposition 2.25, $(C_{h,v})_{1/\xi}(1)_{2}$ exists and
$\big((C_{h,v})_{1/\xi}(1)_{2}\big)^{2}=\big((C_{h,v})_{1/\xi}\big)^{2}(1)_{2}=(C_{h,v}^{2})_{1/\xi}(1)_{2}=(C_{h,v}^{2})_{1/\xi}(1)_{1}=1\,.$
Since $(C_{h,v})_{1/\xi}(1)_{2}\geq 0$, we have $(C_{h,v})_{1/\xi}(1)_{2}=1$ and the stated formula follows.
2. By item 1 of Proposition 4.1 and by Proposition 2.31, $A_{h,v}(x)=C_{h,v}(x)$. By item 2 of Proposition 4.1 and by Theorems 2.16 and 2.17,
$\lim_{x\nearrow 1}F_{(C_{h,v}^{2})_{1/\xi}}(x)=1-\frac{\lim_{x\nearrow 1}F_{(B_{h,v})_{1/\xi}}(x)}{\lim_{x\nearrow 1}F_{(B_{h})_{1/\xi}}(x)}=1-\frac{(B_{h,v})_{1/\xi}(1)_{1}}{\pm\infty}=1\,.$}
By the continuity of the functions $F_{\dots}$,
$\lim_{x\nearrow 1}F_{(A_{h,v})_{1/\xi}}(x)=\lim_{x\nearrow 1}F_{(C_{h,v})_{1/\xi}}(x)\in\{-1,\,1\}\,.$
$\Box$
5 Concluding remarks
In [17] we consider the same problem of determining formulas for quantities such as $A_{h,v}(1)=\sum_{n\geq 0}h(W(v,n))$ in the case of formal weights $h$. These are maps
$h\colon\mathbb{N}_{2}\to\mathbb{C}[[x_{1},\,x_{2},\,\dots,\,x_{k}]]\,$,
where the last integral domain of formal power series in $k$ variables is endowed with the usual non-Archimedean norm.
The main novelty of our generalization is this. In equalities like
$P_{d}(\overline{v},\,n)=h(W(v,\,n))$
that was stated in the Introduction, the left-hand side involves, in the standard rigorous probabilistic treatment, an implicit probability space supporting the corresponding Markov chain. Construction of this probability space is far from trivial; see, for example, [6]. In our generalization, the right-hand side is completely self-contained.