### v8

# Extending the symbolic method in enumerative combinatorics. I

Abstract
We use our extension of the symbolic method in enumerative combinatorics (we extend finite sums determining coefficients in generating functions to infinite series) to generalize Pólya’s theorem. This theorem finds the limits of probabilities that walks in the grid graph $\mathbb{Z}^{d}$, starting at the origin, visit the given vertex. We replace the grid graph with weighted countable complete graphs $K_{\mathbb{N}}$ such that the edge weights have values in $\mathbb{C}$ and satisfy a convergence condition. Thus general Pólya’s theorems are results in enumerative combinatorics.
1 Introduction
We recall in general terms the fundamental symbolic method in enumerative combinatorics, briefly the symbolic method, and explain how we extend it. Later in the article, we exemplify our extension on Pólya’s theorem. Our selection of references on the symbolic method is [3, 8, 11, 13, 29].
1.1 The symbolic method and its extension
Using this method, we enumerate—by exact counting formulas and by approximate asymptotic formulas—an infinite (countable) set $A$ of combinatorial objects. $A$ is equipped with a size function $s\colon A\to\mathbb{N}_{0}$ ($\{0,1,\dots\}$). It is assumed that for every $n\in\mathbb{N}_{0}$ the set of size $n$ objects in $A$,
$A_{n}=\{a\in A\colon\;s(a)=n\}\,$,
is finite; see [3, p. 1], [8, p. ix], [11, Definition I.1 on p. 16] and [13, Remark 2.2.3 on p. 32, implicitly]. We associate with $A$ and $s$ the generating function
$A(x)=\sum_{a\in A}x^{s(a)}=\sum_{n\geq 0}|A_{n}|x^{n}\ \ (\in\mathbb{C}[[x]])\,$,}
where $|X|$ denotes the cardinality of a finite set $X$. It is common, and crucial in the approach here, to have a weight function $h\colon A\to R$, where $R$ is a ring. The weighted version of $A(x)$ is then
$A_{h}(x)=\sum_{n\geq 0}(\sum_{a\in A_{n}}h(a))x^{n}\ \ (\in R[[x]])\,.$}
Note that every coefficient $\sum_{a\in A_{n}}h(a)$ in $A_{h}(x)$ is a finite sum. The symbolic method works in two phases. The first phase translates combinatorial relations between objects in $A$ into algebraic, differential, or functional equations for $A_{h}(x)$ and the auxiliary generating functions. Multivariate generating functions can also be used. The second phase deduces from these equations asymptotic formulas for the coefficients of generating functions. The second analytic phase accounts for the usefulness and popularity of the method (see [11]), but the first formal phase is the primary one (the memoir [3] is an impressive automatization of this phase).
We give two examples to illustrate the symbolic method: one short now and another, detailed in the next subsection. Let $A$ be the set of all proper bracketings $b$, let $s(b)$ be the number of pairs of brackets in $b$ and let $h(b)=1$. For example,
$A_{3}=\{()()(),\,()(()),\,(())(),\,(()()),\,((()))\}\,.$
The numbers $(|A_{n}|)_{n\geq 0}=(1,1,2,5,14,42,\dots)$ are the Catalan numbers [28]. The two phases of the symbolic method yield the following.
  1. 1.
    We consider the first pair of brackets in $b$ and get the algebraic equation $A(x)=1+xA(x)^{2}$ (we omit the index h because the weight is trivial). It leads to the following exact counting formulas.
    1. (a)
      $|A_{0}|=1$ and $|A_{n+1}|=\sum_{j=0}^{n}|A_{j}|\cdot|A_{n-j}|$ for $n\in\mathbb{N}_{0}$.
    2. (b)
      $|A_{n}|=\frac{1}{n+1}\binom{2n}{n}$ for $n\in\mathbb{N}_{0}$.
    3. (c)
      $|A_{0}|=1$ and $|A_{n+1}|=\frac{4n+2}{n+2}\cdot|A_{n}|$ for $n\in\mathbb{N}_{0}$.
    For $m\in\mathbb{N}_{0}$, let $v(m)=\log(m+2)$; up to multiplicative constant, $v(m)$ is the number of binary (or $q$-ary, for any $q\geq 2$) digits in $m$. Each of (a)–(c) is an efficient closed formula for the counting function $C:\;n\mapsto|A_{n}|$ because it provides an algorithm that computes $C$ in
    $O(\max(v(n),\,v(|A_{n}|))^{c})=O((n+1)^{c}),\ n\in\mathbb{N}_{0},\,$
    ($c>0$ is a constant) bit operations. See the next phase for the asymptotics of $|A_{n}|$.
  2. 2.
    By [11, p. 38], the equation $xA(x)^{2}-A(x)+1=0$ (or (b)) produces the approximate asymptotic formula
    $|A_{n}|=(1+o(1))\pi^{-1/2}n^{-3/2}4^{n}\ \ (n\to\infty)\,.$}
    See [11, Chapter VII.7] for the theory of asymptotics of coefficients of algebraic formal power series.
At the end of phase 1 we remind the definition of an efficient closed formula for a counting function, which was given by the first author in [15, p. 10, Wilfian formulas]. This example of application of the symbolic method is standard, maybe even archetypal. Our next example in the next subsection is less standard
Our idea of extension of the symbolic method is simple and natural. We drop the assumption that the sets $A_{n}$ are finite, and replace it with the assumption that the ring $R$ is endowed with a complete norm, for example, $R=\mathbb{R}$ or $R=\mathbb{C}$, and that all possibly infinite series
$\sum_{a\in A_{n}}h(a)$}
determining coefficients in $A_{h}(x)$ absolutely converge. This extension relies on our semiformal approach to formal power series. In this approach, we allow operations with formal power series in which coefficients arise from limit transitions. In this article, $R=\mathbb{C}$ and we work out an extensive example in which we apply the extended symbolic method to generalize Pólya’s theorem, as explained below. In [17] we develop the semiformal approach in general. In [19] we obtain analogs of the results in this article for the ring $R=\mathbb{C}[[x_{1},\dots,x_{k}]]$ which is endowed with a complete non-Archimedean norm. In [16] we use the semiformal approach in number theory and present two proofs of the transcendence of the number $\mathrm{e}$ based on formal power series.
1.2 Pólya’s theorem for $\overline{v}=\overline{0}$ via the symbolic method
Let $d,n\in\mathbb{N}$ ($\{1,2,\dots\}$) and $\overline{v}\in\mathbb{Z}^{d}$, where $\mathbb{Z}=\{\dots,-1,0,1,\dots\}$ are the integers. Let $P_{d}(\overline{v},n)$ be the probability that a walk of length $n$ in the grid graph $\mathbb{Z}^{d}$, starting at the origin $\overline{0}=\langle 0,0,\dots,0\rangle$ ($\in\mathbb{Z}^{d}$), later visits the vertex $\overline{v}$. In 1921, G. Pólya proved in [24] a theorem, which we generalize in our article, that for every dimension $d\in\mathbb{N}$ and every vertex $\overline{v}$,
$\lim_{n\to\infty}P_{d}(\overline{v},\,n)\left\{\begin{array}[]{lll}=1&\dots&\text{if $d\leq 2$ and}\\ \in(0,\,1)&\dots&\text{if $d\geq 3$}\,.\end{array}\right$. (1)
Pólya’s theorem is almost always mentioned in the special case when $\overline{v}=\overline{0}$, but Pólya proved it for every vertex $\overline{v}$ (see [18, Section 1] for a quote in which he states his result). A selection of articles and books discussing Pólya’s theorem is [4, 5, 6, 9, 10, 12, 14, 20, 21, 22, 23, 25, 26, 27, 32, 33].
Now we derive (1) for $\overline{v}=\overline{0}$ by the symbolic method. The grid graph $G_{d}$ is the $2d$-regular graph $G_{d}=\langle\mathbb{Z}^{d},E_{d}\rangle$ with the edges
$\{\overline{a},\,\overline{b}\}\in E_{d}\iff\sum_{i=1}^{d}|\overline{a}_{i}-\overline{b}_{i}|=1\,.$}
We take the constant map $h\colon E_{d}\to\{\frac{1}{2d}\}$ and extend it to walks in $G_{d}$ by $h(w)=(2d)^{-n}$ where $n$ ($\in\mathbb{N}_{0}$) is the length of the walk $w$. Let $A$ be the set of (finite) walks in $G_{d}$ that start at the origin $\overline{0}$ and later revisit it. The size function $s$ is defined on every walk $w$ in $G_{d}$ and $s(w)$ is the length of $w$. Then for every $d\in\mathbb{N}$ and $n\in\mathbb{N}_{0}$,
$P_{d}(\overline{0},\,n)=\sum_{w\in A_{n}}h(w)=:h(A_{n})$}
(recall that $A_{n}=\{w\in A\colon\;s(w)=n\}$). We are interested in the main generating function
$A_{h}(x)=\sum_{n\geq 0}h(A_{n})x^{n}\ \ (\in\mathbb{C}[[x]])\,.$}
We obtain a formula for $\lim_{n\to\infty}h(A_{n})$ with the help of three auxiliary generating functions:
$B_{h}(x)=\sum_{n\geq 0}h(B_{n})x^{n},\ C_{h}(x)=\sum_{n\geq 0}h(C_{n})x^{n},\ D_{h}(x)=\sum_{n\geq 0}h(D_{n})x^{n}\,.$}
Here $B$ is the set of closed walks in $G_{d}$ that begin and end at $\overline{0}$, $C\subset B$ is the subset of closed walks with inner vertices different from $\overline{0}$ and $D$ is the set of all walks in $G_{d}$ starting at $\overline{0}$. We have $h(A_{0})=h(C_{0})=0$ and $h(B_{0})=h(D_{0})=1$.
The first formal phase of the symbolic method yields two formal identities,
$A_{h}(x)=C_{h}(x)D_{h}(x)=C_{h}(x)\frac{1}{1-x}\,\text{ and }\,C_{h}(x)=1-\frac{1}{B_{h}(x)}\ \ (\in\mathbb{C}[[x]])\,.$}
The second equality reflects the fact that $h(D_{n})=1$ for every $n\in\mathbb{N}_{0}$. The other two equalities are also easy to prove; in Proposition 3.1 we prove them in a much more general setting. The first identity implies that for every $n\in\mathbb{N}_{0}$,
$h(A_{n})=\sum_{j=0}^{n}h(C_{j})\,.$}
We proceed to the second analytic phase. We consider the sums
$C_{h}(1)=\sum_{n\geq 0}h(C_{n}):=\lim_{n\to\infty}\sum_{j=0}^{n}h(C_{j})\ \ (\in\mathbb{C})$}
and $B_{h}(1)$ which is defined similarly. Since the formal power series $B_{h}(x)$ and $C_{h}(x)$ have coefficients in $[0,1]$, the formal identity $C_{h}(x)=1-\frac{1}{B_{h}(x)}$ in $\mathbb{C}[[x]]$ converts to the identity
$C_{h}(1)=1-\frac{1}{B_{h}(1)}\ \ (\in[0,\,1])\,$,}
where $B_{h}(1)\in[1,\,+\infty)\cup\{+\infty\}$. We rigorously prove it. We formally have $B_{h}(x)C_{h}(x)=B_{h}(x)-1$. Since $0\leq h(B_{n}),h(C_{n})\leq h(D_{n})=1$ for every $n\in\mathbb{N}_{0}$, the formal power series $B_{h}(x)$ and $C_{h}(x)$ absolutely converge for every $x\in[0,1)$. Using the operations of linear combination, grouping, and product for complex series in Propositions 2.2, 2.4, and 2.5, respectively, we get that
$\text{$B_{h}(x)C_{h}(x)=B_{h}(x)-1$ for every number $x\in[0,\,1)$}\,.$
Since $B_{h}(x)\geq 1$ for every $x\in[0,1)$, we can divide by it and get that
$\text{$C_{h}(x)=1-\frac{1}{B_{h}(x)}$ for every $x\in[0,\,1)$}\,.$}
Using the arithmetic of functional limits and Theorem 2.18 (Abel’s theorem for nonnegative real coefficients), we get that indeed
$C_{h}(1)=\lim_{x\nearrow 1}C_{h}(x)=\lim_{x\nearrow 1}\big(1-\frac{1}{B_{h}(x)}\big)=1-\frac{1}{\lim_{x\nearrow 1}B_{h}(x)}=1-\frac{1}{B_{h}(1)}\,.$}
Here $\lim_{x\nearrow 1}$ denotes that $x$ approaches $1$ through the values $x\in[0,1)$.
Thus
$\lim_{n\to\infty}P_{d}(\overline{0},\,n)=\lim_{n\to\infty}h(A_{n})=\lim_{n\to\infty}{\textstyle\sum_{j=0}^{n}h(C_{j})=C_{h}(1)=1-\frac{1}{B_{h}(1)}\,.}$
Since $B_{h}(1)=+\infty$ for $d\leq 2$ and $1<B_{h}(1)<+\infty$ for $d\geq 3$ (this requires some work), Pólya’s theorem for $\overline{v}=\overline{0}$ follows. $\Box$
See Theorem 3.5 for a generalization of this result. We have not seen the formula
$\lim_{n\to\infty}P_{d}(\overline{0},\,n)=\lim_{n\to\infty}h(A_{n})=1-\frac{1}{B_{h}(1)}\,$,} (2)
which unites the cases $d\leq 2$ and $d\geq 3$ of Pólya’s theorem for $\overline{v}=\overline{0}$, in the literature.
The use of generating functions in the proof of Pólya’s theorem is well known, but the theorem itself is mostly viewed as a result in probability or random walks; see, for example, [6] for a probabilistic proof. One of our goals is to convince the reader that Pólya’s theorem can also be naturally viewed as a result in enumerative combinatorics. The latter interpretation does not require tools of probability theory, such as Markov chains.
1.3 Pólya’s theorem for $\overline{v}\neq\overline{0}$ via the symbolic method
Let $\overline{v}\neq\overline{0}$ and $A_{\overline{v}}$ be the set of walks in $G_{d}$ starting at $\overline{0}$ and visiting $\overline{v}$. In the case $d\leq 2$ when $B_{h}(1)=+\infty$ we have, as a special case of item 1 in Theorem 4.13, that
$\lim_{n\to\infty}P_{d}(\overline{v},\,n)=\lim_{n\to\infty}h(A_{\overline{v},\,n})=1\,.$ (3)
In the case $d\geq 3$ when $1<B_{h}(1)<+\infty$, we can obtain, by a more involved version of the argument in Section 1.2, the formula
$\lim_{n\to\infty}P_{d}(\overline{v},\,n)=\lim_{n\to\infty}h(A_{\overline{v},\,n})={\textstyle\sqrt{1-\frac{B_{\overline{v},\,h}(1)}{B_{h}(1)}}}\,.$ (4)
Here $B_{\overline{v}}$ ($\subset B$) is the set of walks in $G_{d}$ that begin and end at $\overline{0}$ and avoid $\overline{v}$. It is easy to see that
$1\leq B_{\overline{v},\,h}(1)<B_{h}(1)\,$,
which implies the second case in (1). This result is generalized in Theorem 4.9. We have not seen the formula (4) in the literature.
1.4 General Pólya’s theorems
The idea of generalizing Pólya’s theorem is simple. If we extend the grid graph $G_{d}$ to the countable complete graph
$K=\big\langle\mathbb{Z}^{d},\,\binom{\mathbb{Z}^{d}}{2}\big\rangle\,$,}
extend the weight $h\colon E_{d}\to\{\frac{1}{2d}\}$ to the weight $h\colon\binom{\mathbb{Z}^{d}}{2}\to\{0,\frac{1}{2d}\}$ by setting $h=0$ on $\binom{\mathbb{Z}^{d}}{2}\setminus E_{d}$, if we then extend $h$ multiplicatively to walks in $K$ (see below), and if $A_{\overline{v}}$ is the set of walks in $K$ starting at $\overline{0}$ and later visiting the given vertex $\overline{v}$, then nothing changes in the above discussion of Pólya’s theorem in Sections 1.2 and 1.3. So we generalize it by allowing any complex weight $h\colon\binom{\mathbb{Z}^{d}}{2}\to\mathbb{C}$, assuming that a convergence condition holds so that the coefficients $h(A_{\overline{v},n})$ and coefficients in auxiliary generating functions, which are now sums of (possibly) infinite series, exist.
For simplicity we replace $\mathbb{Z}^{d}$ with the set $\mathbb{N}$ of natural numbers and work with the countable complete graph
$K_{\mathbb{N}}=\langle\mathbb{N},\,\mathbb{N}_{2}\rangle,\ \mathbb{N}_{2}=\binom{\mathbb{N}}{2}\,$.}
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a weight. $\mathbb{C}$ denotes the field of complex numbers, with the usual complete Archimedean norm (absolute value) $|z|=\sqrt{z\cdot\overline{z}}$. A walk $w$ in $K_{\mathbb{N}}$ is an $n+1$-tuple of vertices
$w=\langle u_{0},\,u_{1},\,\dots,\,u_{n}\rangle,\,\text{$n\in\mathbb{N}_{0}$ and $u_{i}\in\mathbb{N}$}\,$,
such that $u_{i-1}\neq u_{i}$ for every $i=1,2,\dots,n$. Let $W$ be the set of all walks and $s\colon W\to\mathbb{N}_{0}$ be the size function given by $s(w)=n$, that is, $s(w)$ is the length of $w$. We standardly extend the weight $h$ to $W$: $h(w)=1$ if $s(w)=0$, and
$h(w)=\prod_{i=1}^{n}h(\{u_{i-1},\,u_{i}\})$}
if $s(w)=n\geq 1$. By the reversal of walks we mean the bijection (even involution) $\varphi\colon W\to W$,
$\varphi(\langle u_{0},\,u_{1},\,\dots,\,u_{n}\rangle)=\langle u_{n},\,u_{n-1},\,\dots,\,u_{0}\rangle\,.$
The map $\varphi$ is length-preserving and weight-preserving: always $s(w)=s(\varphi(w))$ and $h(w)=h(\varphi(w))$.
For $n\in\mathbb{N}_{0}$ we denote by $D_{n}$ the set of walks of length (size) $n$ in $K_{\mathbb{N}}$ starting at $1$. Thus $D_{0}=\{\langle 1\rangle\}$ and for $n\geq 1$ the set $D_{n}$ is infinite and countable. Let $v\in\mathbb{N}$. We denote by $A_{v}$ the set of walks in $K_{\mathbb{N}}$ starting at $1$ and later visiting $v$. Then
$A_{v,n}=(A_{v})_{n}=\{\langle u_{0},\,\dots,\,u_{n}\rangle\in W\colon\;\text{$u_{0}=1$ and $\exists\,i>0$ such that $u_{i}=v$}\}\,.$
The said convergence condition is Definition 2.1: every series
$\sum_{w\in D_{n}}h(w),\ n\in\mathbb{N}_{0}\,$,}
absolutely converges. Absolute convergence is defined at the beginning of the next section. Then every series $\sum_{w\in A_{v,n}}h(w)$, $n\in\mathbb{N}_{0}$, absolutely converges as well; we denote its sum by $h(A_{v,n})$. We consider the main generating function
$A_{v,\,h}(x)=(A_{v})_{h}(x)=\sum_{n\geq 0}h(A_{v,\,n})x^{n}\ \ (\in\mathbb{C}[[x]])\,.$}
The classical theorem of Pólya is the formula (1). Many other graph families besides $G(d)$ have been investigated. What should the general Pólya’s theorems be about? It is not so clear, weights in them are general, $h\colon\mathbb{N}_{2}\to\mathbb{C}$, and the original probabilistic interpretation disappears in the distance. We investigate a variety of weights: general weights $h\colon\mathbb{N}_{2}\to\mathbb{C}$, $\xi$-convex weights with $\xi\in\mathbb{C}$ and $|\xi|=1$, 0-convex weights (see Definition 2.10), weights $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ and weights $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}$. Our answer is that general Pólya’s theorems are identities of the forms (2)–(4) for the walks $A_{v}$ in terms of quantities of four kinds, which we present below in the order of decreasing precision.
  1. 1.
    The limit $\lim_{n\to\infty}h(A_{v,n})$ of total weights of walks in $A_{v,n}$.
  2. 2.
    The sum $A_{v,h}(1)_{2}$ of absolutely convergent series
    $h(A_{v,0})+h(A_{v,1})+\dots$
    of those total weights.
  3. 3.
    The sum $A_{v,h}(1)_{1}$ of conditionally convergent series
    $h(A_{v,0})+h(A_{v,1})+\dots\$
    of those total weights.
  4. 4.
    The generalized sum
    $\lim_{x\nearrow 1}A_{v,h}(x)=\lim_{x\nearrow 1}{\textstyle\sum_{n\geq 0}h(A_{v,\,n})x^{n}}\,.$
In accordance with the ranges of weights, these quantities are usually twisted by powers of $\xi$ for $\xi\in\mathbb{C}^{*}$ ($\mathbb{C}\setminus\{0\}$). In Sections 3 and 4 we obtain along these lines a total of sixteen general Pólya’s theorems.
Some questions remain. In Theorem 3.2 we obtain the formula $A_{h}(1)_{1}=(1-\frac{1}{B(1)_{1}})D_{h}(1)_{1}$ which we deduce from the (simpler!) formula $A_{h}(1)_{1}=C_{h}(1)_{1}D_{h}(1)_{1}$. Why do we prefer the former formula over the latter? First, the walks $B$ (starting and ending at $1$) have simpler definition than the walks $C$ (starting and ending at $1$ and avoiding $1$ between). Second, (1) follows from formulas of the former type, but it is not clear how to deduce it from the latter. See also the discussion after Proposition 4.3.
1.5 Plan of the article
In this section, we have hopefully explained the main ideas of our article and its context. To work with generating functions whose coefficients are sums of infinite series, we need a theory of complex series defined on arbitrary countable sets, such as the sets $D_{n}$ of walks in $K_{\mathbb{N}}$ of length $n$ starting at $1$. We could not find such a theory in the literature, at least not in the form we need, and therefore we develop it in Section 2. Here is one example. The product rule (formula) in the symbolic method, the identity that
$|A\times B|=|A|\cdot|B|$
for any two finite sets $A$ and $B$, is so simple and obvious that one encounters difficulties when asked to provide a detailed proof: first, axioms of set theory and natural numbers have to be specified, and so on. The actual product formula operates on the level of generating functions, ordinary or exponential ones, and is less immediate, but it has at its heart the displayed set-theoretic identity; see [3, pp. 22, 40], [8, Theorem on pp. 3–4], [11, Definition I.6 on p. 23] and [13, 2.2.14 Lemma on pp. 36–37]. In our extension of the symbolic method, we lift the product rule to the nontrivial Proposition 2.5 on products of absolutely convergent series. Proposition 2.4 concerns another important operation on series, the grouping. Other useful results in Section 2 are three Abel’s Theorems 2.162.18 (stated without proofs) and various conversions of formal identities to identities at $x=1$ (we saw one example in Section 1.2), such as Proposition 2.30.
In Section 3 we present generalizations of Pólya’s theorem in the simple case $v=1$ when the vertex to be visited by walks in $K_{\mathbb{N}}$ coincides with their starting vertex. In Proposition 3.1, we give detailed proofs of formal identities
$A_{h}(x)=C_{h}(x)D_{h}(x)\,\text{ and }\,C_{h}(x)=1-\frac{1}{B_{h}(x)}\ \ (\in\mathbb{C}[[x]])$}
whose instances for the graph $G_{d}$ we saw in Section 1.2. At first glance, they look like typical identities produced in the first phase of the symbolic method, but they are in fact completely new. The coefficients in the generating functions $A_{h}(x)$, $B_{h}(x)$, $C_{h}(x)$, and $D_{h}(x)$ arise as sums of infinite (absolutely convergent) series. In the symbolic method, such coefficients always arise as finite sums or finite linear combinations. The proof of Proposition 3.1 therefore transcends analogous proofs in the standard symbolic method. Our seven general Pólya’s theorems for $v=1$ are given in Theorems 3.23.8. As we already mentioned, these theorems show that formulas (2)–(4) hold for weights $h$ far more general than probabilistic ones and they explore the logical space of conditional and absolute convergence assumptions on the values at $x=1$ of generating functions. See Section 5 for the summary of these theorems.
Results in Section 4 in Theorems 4.54.13 are analogous to Section 3 but they are obtained in the more complicated case $v\neq 1$ of general Pólya’s theorems; see Section 5 for the summary of these theorems.. Proofs of the formal identities
$A_{v,\,h}(x)=C_{v,\,h}(x)D_{h}(x),\ B_{h}(x)=B_{v,\,h}(x)+C_{v,\,h}(x)^{2}B_{h}(x)\,\text{ and}$
$C_{v,\,h}(x)=B_{v,\,h}(x)E_{v,\,h}(x)$
in Proposition 4.2 are only sketched. For $v\neq 1$ one can introduce more auxiliary generating functions than for $v=1$, and in Proposition 4.3 we give another two formal identities for the generating function $F_{v,h}(x)$ of walks going from $1$ to $v$. In Theorem 4.4 we show that for any weight $h\colon\mathbb{N}_{2}\to\mathbb{R}$ and any vertex $v\neq 1$ we have, for every real $x\in(-1,1)$ and under appropriate assumptions, that if $B_{h}(x)=0$ then $B_{v,h}(x)=0$, and if $B_{h}(x)\neq 0$ then
$\frac{B_{v,\,h}(x)}{B_{h}(x)}\leq 1\,.$
This generalizes the trivial bound
$0\leq B_{v,\,h}(x)\leq B_{h}(x)$
for $x\in[0,1)$ and $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$.
1.6 Perspectives and further research
We already mentioned that in [17] we develop a semiformal theory of formal power series. Many isolated instances of such uses of formal power series can already be found in the literature, and in [17] we want to systematize them. For example, the classical identity
$1-\frac{(\pi x)^{2}}{3!}+\frac{(\pi x)^{4}}{5!}-\dots=\prod_{n\geq 1}\Big(1-\frac{x^{2}}{n^{2}}\Big)\ \ (\in\mathbb{C}[[x]])$
resulting from the representation of the function $\sin(\pi x)/(\pi x)$ by an infinite product can be understood as an identity for formal power series employing semiformal operations. In [19] we generalize Pólya’s theorem to $K_{\mathbb{N}}$ with edge weights in the ring $\mathbb{C}[[x_{1},\dots,x_{k}]]$. There the norm is non-Archimedean, and we are curious what results we obtain. Weights with values in polynomial or formal power series rings are common in enumerative combinatorics. It would be interesting to obtain general Pólya’s theorems for several (maybe infinitely many?) vertices $v\in\mathbb{N}$ to be visited by walks in $K_{\mathbb{N}}$ starting at $1$.
2 Series and generating functions in $\mathbb{C}$
In this section, we develop a theory of general complex series and obtain propositions on the conversion of formal identities for generating functions to identities at $x=1$. We use these results in the two following sections.
2.1 General complex series
series is a map $h\colon X\to\mathbb{C}$ defined on a finite or countable set $X$; we write it also as $\sum_{x\in X}h(x)$. We say that $\sum_{x\in X}h(x)$ absolutely converges if two equivalent conditions hold.
  1. 1.
    For some constant $c\geq 0$ and every finite set $Y\subset X$ we have
    $\sum_{x\in Y}|h(x)|\leq c\,.$}
  2. 2.
    If $X$ is countable, then for every bijection $f\colon\mathbb{N}\to X$, the limit
    $s=\lim_{n\to\infty}{\textstyle\sum_{i=1}^{n}h(f(i))\ \ (\in\mathbb{C})}$
    exists and does not depend on $f$.
In fact, the independence of $s$ on $f$ may be omitted from the condition. We call the number $s$ the sum of the series and denote it again by $\sum_{x\in X}h(x)$ or by $h(X)$. If the set $X$ is finite and nonempty, $X=\{x_{1},x_{2},\dots,x_{n}\}$, then every series $\sum_{x\in X}h(x)$ absolutely converges and has the sum
$h(x_{1})+h(x_{2})+\dots+h(x_{n})\,.$
We set $h(\emptyset)=0$. If
$U=U(x)=\sum_{n\geq 0}u_{n}x^{n}\ \ (\in\mathbb{C}[[x]])$}
is a formal power series, so that $x$ is a formal variable, we denote for any $c\in\mathbb{C}$ by $U(c)$ the sum of the power series $\sum_{n\geq 0}u_{n}c^{n}$, if it absolutely converges. Let
$H_{1}={\textstyle\{U(x)\in\mathbb{C}[[x]]\colon\;\text{$U(c)$ absolutely converges for every $c\in\mathbb{C}_{<1}$}\}\,,}$
where $\mathbb{C}_{<1}=\{c\in\mathbb{C}\colon\;|c|<1\}$. So for any formal power series $U(x)\in H_{1}$ the sums $U(c)$ form the function
$U\colon\mathbb{C}_{<1}\to\mathbb{C},\ c\mapsto U(c)\,.$
Next definition is basic in our approach. Recall that $D_{n}$ is the set of walks in $K_{\mathbb{N}}$ with length $n$ starting at $1$.
Definition 2.1 (light weights)
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a weight. We call $h$ light if for every $n\in\mathbb{N}_{0}$, the series
$\sum_{w\in D_{n}}h(w)$}
absolutely converges.
Let $S=\sum_{x\in X}h(x)$ be a series. If $Y\subset X$, we say that $\sum_{x\in Y}h(x)$ is a subseries of $S$. Any subseries of an absolutely convergent series absolutely converges, but we cannot compare the sizes of sums $h(Y)$ and $h(X)$; we may have $h(X)=0$ and $h(Y)\neq 0$. However, if $h\colon X\to\mathbb{R}_{\geq 0}$, $\sum_{x\in X}h(x)$ absolutely converges, and if $Y\subset X$, then obviously
$0\leq h(Y)\leq h(X)\,.$
We call it the subseries bound.
2.2 Linear combination, grouping and product of series
Suppose that $R=\sum_{x\in X}g(x)$ and $S=\sum_{x\in X}h(x)$ are series and that $\alpha,\beta\in\mathbb{C}$. Then the series
$\alpha R+\beta S=\sum_{x\in X}(\alpha g(x)+\beta h(x))$}
is the linear combination of $R$ and $S$. We omit the easy proof of the next result.
Proposition 2.2
If $R$ and $S$ absolutely converge and have sums $r$ and $s$, respectively, then $\alpha R+\beta S$ absolutely converges and has sum $\alpha r+\beta s$.
This proposition easily extends to linear combinations with more terms. The next approximation lemma for series is useful.
Lemma 2.3
If $S=\sum_{x\in X}h(x)$ absolutely converges and has sum $s$, then for every $\varepsilon>0$ there exists a finite set $Y\subset X$, denoted by $Y(S,\,\varepsilon)$, such that for every finite set $Z$ with $Y\subset Z\subset X$ we have
$|s-h(Z)|\leq\varepsilon\,.$
Proof. Let an $\varepsilon>0$ be given. For finite $X$ we set $Y=X$. For countable $X$ we take any bijection $f\colon\mathbb{N}\to X$, take an $N\in\mathbb{N}$ such that
$\big|\sum_{n=1}^{N}h(f(n))-s\big|\leq\frac{\varepsilon}{2}\,\text{ and }\,\sum_{n>N}|h(f(n))|\leq\frac{\varepsilon}{2}\,$,}
and set $Y=\{f(1),f(2),\dots,f(N)\}$. Then for every finite set $Z$ with $Y\subset Z\subset X$ we have
$|s-h(Z)|\leq|s-h(Y)|+\sum_{x\in Z\setminus Y}|h(x)|\leq\frac{\varepsilon}{2}+\frac{\varepsilon}{2}=\varepsilon\,.$}
$\Box$
Two important operations on series are grouping and product. A partition of a set $X$ is a set $P$ of nonempty and disjoint sets such that $\bigcup P=X$. If $X$ is at most countable, then so is $P$ and every set $Z\in P$. If $S=\sum_{x\in X}h(x)$ is a series and $P$ is a partition of $X$ such that for every $Z\in P$ the subseries $S_{Z}=\sum_{x\in Z}h(x)$ absolutely converges and has sum $s_{Z}$, then the series
$S_{P}=\sum_{Z\in P}s_{Z}$}
is called the grouping of $S$.
Proposition 2.4
If $S=\sum_{x\in X}h(x)$ absolutely converges and has sum $s$, and if $P$ is a partition of $X$, then the grouping series $S_{P}$ is correctly defined, absolutely converges and has the same sum $s$.
Proof. For every $Z\in P$ the series $S_{Z}=\sum_{x\in Z}h(x)$ is a subseries of $S$ and absolutely converges. Thus the series $S_{P}$ is correctly defined. We show that it absolutely converges. Let
$c=\sup(\{\sum_{x\in Z}|h(x)|\colon\;\text{$Z\subset X$ and $Z$ is finite}\})\ \ (<+\infty)$}
and $\{Z_{1},Z_{2},\dots,Z_{n}\}\subset P$ be a finite set. We use Lemma 2.3, take for $i=1,2,\dots,n$ finite sets
$Z_{i}^{\prime}=Y(S_{Z_{i}},\,2^{-i})\ \ (\subset Z_{i})$
and define the finite set $Z_{0}=Z_{1}^{\prime}\cup Z_{2}^{\prime}\cup\dots\cup Z_{n}^{\prime}$ ($\subset X$). Then
${\textstyle\sum_{i=1}^{n}|h(Z_{i})|}$ $\leq$ ${\textstyle\sum_{i=1}^{n}|h(Z_{i})-h(Z_{i}^{\prime})|+\sum_{x\in Z_{0}}|h(x)|}$
$\leq$ ${\textstyle\sum_{i=1}^{n}2^{-i}+c\leq 1+c}$
and we see that the series $S_{P}$ absolutely converges.
Let $t$ be the sum of $S_{P}$. We show that $|t-s|\leq\varepsilon$ for every $\varepsilon>0$; then $t=s$. Let an $\varepsilon>0$ be given. We use Lemma 2.3 and take finite sets
$X^{\prime}=Y(S,\,\frac{\varepsilon}{3})\ \ (\subset X)\,\text{ and }\,P^{\prime}=Y(S_{P},\,\frac{\varepsilon}{3})\ \ (\subset P)\,.$}
We take a finite set $\{Z_{1},\dots,Z_{n}\}\subset P$ such that
$P^{\prime}\subset\{Z_{1},\,\dots,\,Z_{n}\}\,\text{ and }\,X^{\prime}\subset\bigcup_{i=1}^{n}Z_{i}\,.$}
We again use Lemma 2.3 and take for $i=1,2,\dots,n$ finite sets
$Z_{i}^{\prime}=Y(S_{Z_{i}},\,2^{-i}\cdot\frac{\varepsilon}{3})\ \ (\subset Z_{i})\,.$}
For $i=1,2,\dots,n$ let
$Z_{i}^{\prime\prime}=Z_{i}^{\prime}\cup(X^{\prime}\cap Z_{i})\ \ (\subset Z_{i})\,.$
We form the disjoint union
$X_{0}=\bigcup_{i=1}^{n}Z_{i}^{\prime\prime}\ \ (\subset X)\,.$}
Then $X_{0}$ is a finite set and $X^{\prime}\subset X_{0}$. Also, $Z_{i}^{\prime}\subset Z_{i}^{\prime\prime}$ for every $i=1,2,\dots,n$. We have
$|s-t|$ $\leq$ ${\textstyle|s-h(X_{0})|+\sum_{i=1}^{n}|h(Z_{i}^{\prime\prime})-h(Z_{i})|+|\sum_{i=1}^{n}h(Z_{i})-t|}$
$\leq$ ${\textstyle\frac{\varepsilon}{3}+\sum_{i=1}^{n}2^{-i}\cdot\frac{\varepsilon}{3}+\frac{\varepsilon}{3}\leq\varepsilon\,.}$
$\Box$
If $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i=1,2$ are series, then the series
$S_{1}\times S_{2}=\sum_{\langle x,\,y\rangle\in X_{1}\times X_{2}}h_{1}(x)\cdot h_{2}(y)$}
is called the product of the series $S_{1}$ and $S_{2}$.
Proposition 2.5
If $S_{1}$ and $S_{2}$ absolutely converge and have sums $s_{1}$ and $s_{2}$, respectively, then the product series $S_{1}\times S_{2}$ absolutely converges and has the sum $s_{1}s_{2}$.
Proof. We show that the series $S_{1}\times S_{2}$ absolutely converges. Let $c\geq 0$ be such that for any finite sets $Y_{i}\subset X_{i}$, $i=1,2$ we have $|h_{i}|(Y_{i})\leq c$. Let $A\subset X_{1}\times X_{2}$ be a finite set. We take finite sets $Y_{i}\subset X_{i}$ such that $A\subset Y_{1}\times Y_{2}$ and get the desired bound:
$\sum_{\langle x,\,y\rangle\in A}|h_{1}(x)h_{2}(y)|\leq\sum_{x\in Y_{1}}|h_{1}(x)|\cdot\sum_{y\in Y_{2}}|h_{2}(y)|\leq c\cdot c=c^{2}\,.$}
We show that the sum $s$ of $S_{1}\times S_{2}$ equals $s_{1}s_{2}$. It suffices to prove that for every $\varepsilon\in(0,1)$ we have $|s-s_{1}s_{2}|\leq\varepsilon$. So let $\varepsilon\in(0,1)$. Using Lemma 2.3 we take finite sets
$X^{\prime}_{1}=Y(S_{1},\,\frac{\varepsilon}{4(1+|s_{2}|)})\ \ (\subset X_{1}),\ \ X^{\prime}_{2}=Y(S_{2},\,\frac{\varepsilon}{4(1+|s_{1}|)})\ \ (\subset X_{2})\$}
and
$Z=Y(S_{1}\times S_{2},\frac{\varepsilon}{4})\ \ (\subset X_{1}\times X_{2})\,.$}
We take finite sets $X_{1}^{\prime\prime}$ and $X_{2}^{\prime\prime}$ such that $X_{1}^{\prime}\subset X_{1}^{\prime\prime}\subset X_{1}$, $X_{2}^{\prime}\subset X_{2}^{\prime\prime}\subset X_{2}$ and $Z\subset X_{1}^{\prime\prime}\times X_{2}^{\prime\prime}$. Then
$|s-s_{1}s_{2}|$ $\leq$ ${\textstyle\big|s-\sum_{\langle x,\,y\rangle\in X_{1}^{\prime\prime}\times X_{2}^{\prime\prime}}h_{1}(x)h_{2}(y)\big|+|h_{1}(X_{1}^{\prime\prime})h_{2}(X_{2}^{\prime\prime})-s_{1}s_{2}|}$
$\leq$ ${\textstyle\frac{\varepsilon}{4}+|(s_{1}+t_{1})(s_{2}+t_{2})-s_{1}s_{2}|\,,}$
where $|t_{1}|\leq\frac{\varepsilon}{4(1+|s_{2}|)}$ and $|t_{2}|\leq\frac{\varepsilon}{4(1+|s_{1}|)}$. Hence $|s-s_{1}s_{2}|\leq\frac{\varepsilon}{4}+|s_{1}t_{2}|+|s_{2}t_{1}|+|t_{1}t_{2}|\leq\frac{\varepsilon}{4}+\frac{\varepsilon}{4}+\frac{\varepsilon}{4}+\frac{\varepsilon}{4}=\varepsilon$. $\Box$
Let $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i=1,2,\dots,k$ be series. Their product is the series
$S_{1}\times S_{2}\times\dots\times S_{k}=\sum_{\overline{x}\in\overline{X}}h_{1}(x_{1})\cdot h_{2}(x_{2})\cdot\ldots\cdot h_{k}(x_{k})\,$,}
where $\overline{x}=\langle x_{1},\dots,x_{k}\rangle$ and $\overline{X}=X_{1}\times\dots\times X_{k}$. By iterating Proposition 2.5 we get the following result.
Proposition 2.6
Let $S_{1}$, $S_{2}$, $\dots$, $S_{k}$ be absolutely convergent series with respective sums $s_{1}$, $s_{2}$, $\dots$, $s_{k}$. Then the series $S_{1}\times S_{2}\times\dots\times S_{k}$ absolutely converges and has sum $s_{1}s_{2}\dots s_{k}$.
Interestingly, a reversal of Proposition 2.5 holds.
Corollary 2.7
Let $S_{1}$ and $S_{2}$ be nonempty series that are not identically zero. If $S_{1}\times S_{2}$ absolutely converges and has sum $s$, then $S_{1}$ and $S_{2}$ absolutely converge, and their respective sums $s_{1}$ and $s_{2}$ satisfy $s_{1}s_{2}=s$.
Proof. This follows from Proposition 2.5 if we show that $S_{1}$ and $S_{2}$ absolutely converge. Let $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i=1,2$, and let $c>0$ be such that for every finite set $X\subset X_{1}\times X_{2}$ we have
$\sum_{\langle x,y\rangle\in X}|h_{1}(x)h_{2}(y)|\leq c\,.$}
We show that $S_{1}$ absolutely converges. Let $y_{0}\in X_{2}$ be such that $h_{2}(y_{0})\neq 0$. Then for every finite set $X\subset X_{1}$ we have
$|h_{1}|(X)=\frac{1}{|h_{2}(y_{0})|}\sum_{\langle x,y\rangle\in X\times\{y_{0}\}}|h_{1}(x)h_{2}(y)|\leq\frac{c}{|h_{2}(y_{0})|}\,.$}
Thus $S_{1}$ absolutely converges. For the series $S_{2}$ we argue similarly. $\Box$
It is clear that the assumptions on series $S_{1}$ and $S_{2}$ cannot be omitted.
2.3 Slim weights
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and $n\in\mathbb{N}_{0}$. We obtain a sufficient condition for lightness of the weight $h$ in terms of its values. We denote by $V(h,n)$ the set of vertices in $K_{\mathbb{N}}$ reachable from $1$ by a walk with length at most $n$ and nonzero weight.
Definition 2.8 (slim weights)
A weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is slim if for every $n$ in $\mathbb{N}_{0}$ there is a constant $c=c(n)>0$ such that for every vertex $u\in V(h,n)$ and every finite set $X\subset\mathbb{N}\setminus\{u\}$ we have
$\sum_{v\in X}|h(\{u,\,v\})|\leq c\,.$}
Proposition 2.9
Every slim weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is light.
Proof. Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be slim. We show by induction on $n\in\mathbb{N}_{0}$ that the series $\sum_{w\in D_{n}}h(w)$ absolutely converges. For $n=0$ it is trivial because $D_{0}=\{\langle 1\rangle\}$. Let $n>0$ and let $c>0$ be a constant such that
$\sum_{w\in X}|h(w)|\leq c\,\text{ and }\,\sum_{v\in X^{\prime}}|h(\{u,\,v\})|\leq c$}
for every finite set $X\subset D_{n-1}$, every $u\in V(h,n-1)$ and every finite set $X^{\prime}\subset\mathbb{N}\setminus\{u\}$.
Let $X\subset D_{n}$ be a finite set. We may assume that all walks in $X$ have nonzero weights. We decompose every walk $w\in X$ as $w=w^{\prime}\,\ell(w)$, where $w^{\prime}\in D_{n-1}$ and $\ell(w)$ is the last vertex. Let $Y$ be the set of walks $w^{\prime}$. For $w^{\prime}\in Y$ let
$X(w^{\prime})=\{v\in\mathbb{N}\setminus\{\ell(w^{\prime})\}\colon\;w^{\prime}v\in X\}\,.$
We get the bound
${\textstyle\sum_{w\in X}|h(w)|}$ $=$ ${\textstyle\sum_{w^{\prime}\in Y}|h(w^{\prime})|\cdot\sum_{v\in X(w^{\prime})}|h(\{\ell(w^{\prime}),\,v\})|}$
$\leq$ ${\textstyle\sum_{w^{\prime}\in Y}|h(w^{\prime})|\cdot c\leq c\cdot c=c^{2}\,.}$
$\Box$
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$. We define $V(h)=\bigcup_{n\geq 0}V(h,n)$. It is the set of vertices in $K_{\mathbb{N}}$ reachable from $1$ by walks with nonzero weights. We define the graph
$G(h)=\langle V(h),\,E(h)\rangle$
as the connected component of $1$ in the subgraph of $K_{\mathbb{N}}$ formed by edges with nonzero weight. If $G(h)$ has finite vertex degrees, then every series $\sum_{w\in D_{n}}h(w)$ has only finitely many nonzero summands and $h$ is therefore light.
2.4 $\xi$-convex weights
We generalize the concept of convexity.
Definition 2.10 ($\xi$-convex weights)
Suppose that $\xi\in\mathbb{C}$. We say that a light weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is $\xi$-convex if for every vertex $u\in V(h)$, the sum
$\sum_{e\in E(h),\,e\ni u}h(e)=\sum_{v\in\mathbb{N}\setminus\{u\}}h(\{u,\,v\})=\xi\,.$}
We say that a light weight is convex if it is $1$-convex.
Note that 0-convex weights are the ones that satisfy Kirchhoff’s second law. We show that the sums in the definition exist.
Proposition 2.11
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight and let $u\in V(h)$. Then the series
$\sum_{v\in\mathbb{N}\setminus\{u\}}h(\{u,v\})$}
absolutely converges.
Proof. There is a walk $w_{0}=\langle u_{0},\dots,u_{n}\rangle$ in $G(h)$ such that $u_{0}=1$ and $u_{n}=u$. Let $c=h(w_{0})$ ($\neq 0$) and $W$ be the set of walks $\langle u_{0},\dots,u_{n},u_{n+1}\rangle$ in $K_{\mathbb{N}}$. Thus, the vertices $u_{0}$, $\dots$, $u_{n}$ are fixed, and $u_{n+1}$ runs in $\mathbb{N}\setminus\{u_{n}\}=\mathbb{N}\setminus\{u\}$. The series $\sum_{w\in W}h(w)$ is a subseries of $\sum_{w\in D_{n+1}}h(w)$ and therefore absolutely converges. By Proposition 2.2, the series
$\sum_{v\in\mathbb{N}\setminus\{u\}}h(\{u,v\})=\frac{1}{c}\sum_{w\in W}h(w)$}
absolutely converges. $\Box$
We proceed to generating functions.
2.5 Sums $U(1)_{1}$ and $U(1)_{2}$ for a generating function $U(x)$
We use univariate formal power series
$U=U(x)=\sum_{n\geq 0}u_{n}x^{n}$}
with complex coefficients. They form an integral domain denoted by $\mathbb{C}[[x]]$. For $U=\sum_{n\geq 0}u_{n}x^{n}$ and $V=\sum_{n\geq 0}v_{n}x^{n}$ in $\mathbb{C}[[x]]$,
$U+V=\sum_{n\geq 0}(u_{n}+v_{n})x^{n}\,\text{ and }\,U\cdot V=UV=\sum_{n\geq 0}\big(\sum_{j=0}^{n}u_{j}v_{n-j}\big)x^{n}\,.$}
Neutral elements are $0=0x^{0}+0x^{1}+0x^{2}+\dots$ and $1=1x^{0}+0x^{1}+0x^{2}+\cdots$. Units in $\mathbb{C}[[x]]$ are exactly the formal power series $\sum_{n\geq 0}u_{n}x^{n}$ with $u_{0}\neq 0$.
Let $\mathbb{C}_{1}=\{x\in\mathbb{C}\colon\;|x|=1\}$ and $\xi\in\mathbb{C}_{1}$. We define the set
$\xi\mathbb{R}_{\geq 0}=\{\xi x\colon\;x\in\mathbb{R}_{\geq 0}\}\ \ (\subset\mathbb{C})\,.$
It is the ray in $\mathbb{C}$ emanating from the origin in the direction $\xi$. It is closed under addition. If $\alpha_{j}\in\xi_{j}\mathbb{R}_{\geq 0}$, $j=1,2$ then $\alpha_{1}\alpha_{2}\in(\xi_{1}\xi_{2})\mathbb{R}_{\geq 0}$. If
$h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$
is a light weight, $n\in\mathbb{N}_{0}$ and $X\subset D_{n}$, then $h(X)\in\xi^{n}\mathbb{R}_{\geq 0}$. The lines $\xi\mathbb{R}$ in $\mathbb{C}$ going through the origin in the direction $\xi$ have the same properties. For $\alpha\in\mathbb{C}$ and $U=U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{C}[[x]]$ we define
$(U)_{\alpha}=(U)_{\alpha}(x):=U(\alpha x)=\sum_{n\geq 0}u_{n}\alpha^{n}x^{n}\ \ (\in\mathbb{C}[[x]])\,.$}
Note that $(U+V)_{\alpha}=(U)_{\alpha}+(V)_{\alpha}$ and $(U\cdot V)_{\alpha}=(U)_{\alpha}\cdot(V)_{\alpha}$.
Definition 2.12 ($U(1)_{1}$ and $U(1)_{2}$)
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$ be in $\mathbb{C}[[x]]$. We introduce the following terminology and objects.
  1. 1.
    If the (proper) limit
    $U(1)_{1}:=\lim_{n\to\infty}{\textstyle\sum_{j=0}^{n}u_{j}\ \ (\in\mathbb{C})}$
    exists, we say that $U(1)_{1}$ exists.
  2. 2.
    If the series
    $\sum_{n\geq 0}u_{n}1^{n}=\sum_{n\geq 0}u_{n}$}
    absolutely converges, we denote its sum by $U(1)_{2}$ and say that $U(1)_{2}$ exists.
If $U(1)_{2}$ exists, then $U(1)_{1}$ exists and $U(1)_{1}=U(1)_{2}$. The opposite implication is, in general, not true. For example, if
$U(x)=\sum_{n\geq 1}(-1)^{n-1}\frac{1}{n}x^{n}$}
then $U(1)_{1}=\log 2$, but $U(1)_{2}$ does not exist. But if
$U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{R}_{\geq 0}[[x]]$}
and $U(1)_{1}$ exists, then $U(1)_{2}$ exists and $U(1)_{1}=U(1)_{2}$. The following is a corollary of Hadamard’s formula for the radius of convergence.
Proposition 2.13
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{C}[[x]]$. Then $U(x)\in H_{1}$ iff
$\limsup_{n\to\infty}|u_{n}|^{1/n}\leq 1\,.$
Thus, for example, the existence of $U(1)_{1}$ is a sufficient (but not necessary) condition for $U(x)\in H_{1}$.
Note that if $U\in\mathbb{C}[[x]]$ then $(U)_{\xi}\in H_{1}$ either for all or no $\xi\in\mathbb{C}_{1}$.
Definition 2.14 ($U(1)_{1}=\pm\infty$)
If $U(x)=\sum_{n\geq 0}u_{n}x^{n}$ is in $\mathbb{R}[[x]]$ and
$\lim_{n\to\infty}\sum_{j=0}^{n}u_{j}=+\infty\,$,}
we say that $U(1)_{1}=+\infty$. We similarly define that $U(1)_{1}=$-$\infty$.
Proposition 2.15
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$ and $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ be in $\mathbb{C}[[x]]$, and let $\xi\in\mathbb{C}^{*}$. If $(V)_{1/\xi}(1)_{1}$ exists and $U(x)=V(x)\frac{1}{1-\xi x}$, then
$\lim_{n\to\infty}\xi^{-n}u_{n}=(V)_{1/\xi}(1)_{1}\,.$
Proof. Let $n\in\mathbb{N}_{0}$. Since $\frac{1}{1-\xi x}=1+\xi x+\xi^{2}x^{2}+\dots$, we have
$u_{n}=v_{0}\xi^{n}+v_{1}\xi^{n-1}+\dots+v_{n}\xi^{0}=v_{0}\cdot\xi^{n}+v_{1}\xi^{-1}\cdot\xi^{n}+\dots+v_{n}\xi^{-n}\cdot\xi^{n}\,.$
Dividing by $\xi^{n}$ and taking the limit for $n\to\infty$, we get the stated formula. $\Box$
2.6 Three Abelian theorems
We manipulate the quantities $U(1)_{1}$ by means of three Abelian theorems. We state them only for the radius of convergence at least $1$. The first theorem is classical, see [1, 2] and [30, Chapter II.7].
Theorem 2.16 (Abel 1)
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{C}[[x]]$. If $U(1)_{1}$ exists, so that $U(x)\in H_{1}$ by Proposition 2.13, then
$\lim_{x\nearrow 1}U(x)=U(1)_{1}\ \ (\in\mathbb{C})\,.$
The second theorem is seldom mentioned, but see [2] for a simple proof. It is given there only in the case $U(1)_{1}=+\infty$, but it works for the sum and limit -$\infty$ as well.
Theorem 2.17 (Abel 2)
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{R}[[x]]\cap H_{1}$. If $U(1)_{1}=\pm\infty$ then
$\lim_{x\nearrow 1}U(x)=U(1)_{1}=\pm\infty\,.$
The third theorem admits only nonnegative real coefficients. It has two advantages over the previous Abelian theorems: it works for every $U(x)$ in $\mathbb{R}_{\geq 0}[[x]]$ and it is an equivalence, one can go not only from the sum to the limit, but also in the reverse. We omit the easy proof.
Theorem 2.18 (Abel 3)
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{R}_{\geq 0}[[x]]\cap H_{1}$. Then the following limit and sum always exist, are possibly $\infty$ and are equal,
$\lim_{x\nearrow 1}U(x)=U(1)_{1}\ \ (\in[0,\,+\infty)\cup\{+\infty\})\,.$
2.7 Conversion of identities in conditional convergence
We believe that the precise statements and proofs of conversions of formal identities to identities at $x=1$, given in this and the next subsection, is our original contribution to the subject.
Proposition 2.19
Let $\alpha,\beta\in\mathbb{C}$ and $U(x),V(x),W(x)\in\mathbb{C}[[x]]$. If $U(x)=\alpha V(x)+\beta W(x)$ and if $V(1)_{1}$ and $W(1)_{1}$ exist, then
$U(1)_{1}=\alpha V(1)_{1}+\beta W(1)_{1}\,.$
Proof. Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$, $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ and $W(x)=\sum_{n\geq 0}w_{n}x^{n}$. Then
$\sum_{j=0}^{n}u_{j}=\sum_{j=0}^{n}(\alpha v_{j}+\beta w_{j})=\alpha\sum_{j=0}^{n}v_{j}+\beta\sum_{j=0}^{n}w_{j}\,.$}
Taking the limit for $n\to\infty$ we get $U(1)_{1}=\alpha V(1)_{1}+\beta W(1)_{1}$. $\Box$
In this proposition, the formulation “then $U(1)_{1}=\alpha V(1)_{1}+\beta W(1)_{1}$” abbreviates the formulation “then $U(1)_{1}$ exists and $U(1)_{1}=\alpha V(1)_{1}+\beta W(1)_{1}$”. We freely use such abbreviations. This result easily generalizes to linear combinations with more terms.
Proposition 2.20
Let $U(x)$, $V_{1}(x)$, $\dots$, $V_{n}(x)\in\mathbb{C}[[x]]$ be $n+1\geq 2$ formal power series. If $U(x)=\prod_{i=1}^{n}V_{i}(x)$ and if $U(1)_{1}$, $V_{1}(1)_{1}$, $\dots$, $V_{n}(1)_{1}$ exist, then
$U(1)_{1}=\prod_{i=1}^{n}V_{i}(1)_{1}\,.$}
Proof. By Propositions 2.4, 2.6 and 2.13, and by the assumptions we have
$U(x)=\prod_{i=1}^{n}V_{i}(x)$ for every number $x\in\mathbb{C}_{<1}$
(see proofs of Propositions 2.24 and 2.27 for details of similar arguments). Using Theorem 2.16 and properties of limits of functions we get
$U(1)_{1}$ $=$ $\lim_{x\nearrow 1}U(x)=\lim_{x\nearrow 1}{\textstyle\prod_{i=1}^{n}V_{i}(x)}$
$=$ ${\textstyle\prod_{i=1}^{n}\lim_{x\nearrow 1}V_{i}(x)=\prod_{i=1}^{n}V_{i}(1)_{1}}\,.$
$\Box$
Since we have to assume that $U(1)_{1}$ exists, the case $n>2$ does not follow by iterating the case $n=2$. The example in [7] shows that the assumption of existence of $U(1)_{1}$ cannot be omitted.
Proposition 2.21
Let $U(x),V(x)\in\mathbb{C}[[x]]$. If $U(x)=\frac{1}{V(x)}$, and if $U(1)_{1}$ and $V(1)_{1}$ exist, then $V(1)_{1}\neq 0$ and
$U(1)_{1}=\frac{1}{V(1)_{1}}\,.$
Proof. In $\mathbb{C}[[x]]$ we have the identity $U(x)V(x)=1$. Proposition 2.20 gives that $U(1)_{1}V(1)_{1}=1$. Thus $V(1)_{1}\neq 0$, and by dividing by it we get the stated identity. $\Box$
Proposition 2.22
Let $U(x),V(x),W(x)\in\mathbb{C}[[x]]$. If $U(x)^{2}=1-\frac{V(x)}{W(x)}$, if $U(1)_{1}$, $V(1)_{1}$ and $W(1)_{1}$ exist, and if $W(1)_{1}\neq 0$, then
$\big(U(1)_{1}\big)^{2}=1-\frac{V(1)_{1}}{W(1)_{1}}\,.$
Proof. Formally $U(x)^{2}\cdot W(x)=W(x)-V(x)$. Let T(x)=W(x)V(x)T(x)=W(x)-V(x). By Proposition 2.19, $T(1)_{1}$ exists and $T(1)_{1}=W(1)_{1}-V(1)_{1}$. By Proposition 2.20 we have $(U(1)_{1})^{2}\cdot W(1)_{1}=T(1)_{1}=W(1)_{1}-V(1)_{1}$. Dividing by $W(1)_{1}$ we get the stated formula. $\Box$
2.8 Conversion of identities in absolute convergence
We obtain analogs of the results in the previous subsection for absolute convergence.
Proposition 2.23
Let $\alpha,\beta\in\mathbb{C}$ and $U(x),V(x),W(x)\in\mathbb{C}[[x]]$. If $U(x)=\alpha V(x)+\beta W(x)$ and if $V(1)_{2}$ and $W(1)_{2}$ exist, then
$U(1)_{2}=\alpha V(1)_{2}+\beta W(1)_{2}\,.$
Proof. Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$, $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ and $W(x)=\sum_{n\geq 0}w_{n}x^{n}$. Since $V(1)_{2}$ and $W(1)_{2}$ are sums of absolutely convergent series $\sum_{n\in\mathbb{N}_{0}}v_{n}$ and $\sum_{n\in\mathbb{N}_{0}}w_{n}$, respectively, and $u_{n}=\alpha v_{n}+\beta w_{n}$, the stated formula follows from Proposition 2.2. $\Box$
Proposition 2.24
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$, $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ and $W(x)=\sum_{n\geq 0}w_{n}x^{n}$ be in $\mathbb{C}[[x]]$. If $U(x)=V(x)W(x)$ and if $V(1)_{2}$ and $W(1)_{2}$ exist, then the following holds.
  1. 1.
    $U(1)_{2}$ exists and $U(1)_{2}=V(1)_{2}W(1)_{2}$.
  2. 2.
    The sum $\sum_{n\geq 0}|u_{n}|\leq\sum_{n\geq 0}|v_{n}|\cdot\sum_{n\geq 0}|w_{n}|$.
Proof. 1. We consider the series
$S=\sum_{\langle j,\,k\rangle\in\mathbb{N}_{0}^{2}}v_{j}w_{k}=\sum_{j\in\mathbb{N}_{0}}v_{j}\times\sum_{k\in\mathbb{N}_{0}}w_{k}\,.$}
By Proposition 2.5, the series $S$ absolutely converges and has the sum $V(1)_{2}W(1)_{2}$. On the other hand, $S$ has the grouping
$\sum_{n\geq 0}\sum_{j=0}^{n}v_{j}w_{n-j}=\sum_{n\geq 0}u_{n}\,.$}
Here the inner $\sum$ means sum and the outer $\sum$ means series. By Proposition 2.4, the series $S$ has the sum $u_{0}+u_{1}+\dots=U(1)_{2}=V(1)_{2}W(1)_{2}$.
2. This follows from the inequality ($n\in\mathbb{N}_{0}$)
$\sum_{m=0}^{n}|u_{m}|=\sum_{m=0}^{n}\big|\sum_{j=0}^{m}v_{j}w_{m-j}\big|\leq\sum_{j=0}^{n}|v_{j}|\cdot\sum_{j=0}^{n}|w_{j}|\,.$}
$\Box$
For nonnegative coefficients one can go from the product to the factors.
Proposition 2.25
Let $U(x)$, $V(x)$ and $W(x)$ be in $\mathbb{R}_{\geq 0}[[x]]\cap H_{1}$. If $U(x)=V(x)W(x)$, $V(x)W(x)\neq 0$ and $U(1)_{2}$ exists, then $V(1)_{2}$ and $W(1)_{2}$ exist, and
$U(1)_{2}=V(1)_{2}W(1)_{2}\,.$
Proof. By Theorem 2.18 and by the assumption $V(x)W(x)\neq 0$, we have $V(1)_{2}=V(1)_{1}=\lim_{x\nearrow 1}V(x)>0$, and the same applies for $W(x)$. Since $U(x)=V(x)W(x)$ for every $x\in\mathbb{C}_{<1}$, we have
$(+\infty>)\ U(1)_{2}$ $=$ $U(1)_{1}=\lim_{x\nearrow 1}U(x)=\lim_{x\nearrow 1}(V(x)W(x))$
$=$ $\lim_{x\nearrow 1}V(x)\cdot\lim_{x\nearrow 1}W(x)=V(1)_{1}W(1)_{1}=V(1)_{2}W(1)_{2}\,.$
In particular, $V(1)_{2},W(1)_{2}<+\infty$. In the second and fifth equalities we used Theorem 2.18. In the fourth equality we used the arithmetic of functional limits. It holds because the product of limits is neither $0\cdot(+\infty)$ nor $(+\infty)\cdot 0$. $\Box$
Proposition 2.20 for $n=2$ and item 1 of Proposition 2.24 are incomparable. The former assumes only conditional convergence, but for all three series. The latter assumes stronger absolute convergence, but only for two series. Another variant of this results is due to F. Mertens.
Theorem 2.26 (Mertens)
Suppose that $U(x)$, $V(x)$ and $W(x)$ are in $\mathbb{C}[[x]]$. If $U(x)=V(x)W(x)$, and if $V(1)_{1}$ and $W(1)_{2}$ exist, then
$U(1)_{1}=V(1)_{1}W(1)_{2}\,.$
See [7] for a proof.
Proposition 2.27
Let $U(x)=\sum_{n\geq 0}u_{n}x^{n}$, $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ be in $\mathbb{C}[[x]]$ and $v_{0}=1$. If $U(x)=\frac{1}{V(x)}$ and $c=|v_{1}|+|v_{2}|+\dots<1$, then the following holds.
  1. 1.
    $U(1)_{2}$ and $V(1)_{2}$ exist, $V(1)_{2}\neq 0$ and $U(1)_{2}=\frac{1}{V(1)_{2}}$.
  2. 2.
    The sum $\sum_{n\geq 0}|u_{n}|\leq\frac{1}{1-c}$.
Proof. 1. It follows from the assumption on $c$ that $V(1)_{2}$ exists. From
$|V(1)_{2}-1|\leq c<1$
we get $V(1)_{2}\neq 0$. We consider the series
$S=1+\sum_{\begin{subarray}{c}m\in\mathbb{N}\\ j_{m,1},\,j_{m,2},\,\dots,\,j_{m,m}\in\mathbb{N}\end{subarray}}(-1)^{m}v_{j_{m,1}}v_{j_{m,2}}\dots v_{j_{m,m}}\,.$}
We show that $S$ absolutely converges. This follows from the bound that the sum of absolute values of summands in any finite subseries of $S$ is at most
$1+\sum_{m\geq 1}(|v_{1}|+|v_{2}|+\dots)^{m}=1+\sum_{m\geq 1}c^{m}=\frac{1}{1-c}\,.$}
Let $m\in\mathbb{N}$. We consider the series
$S_{m}=\prod_{i=1}^{m}(v_{1}+v_{2}+\dots)=\sum_{\langle j_{1},\,j_{2},\,\dots,\,j_{m}\rangle\in\mathbb{N}^{m}}v_{j_{1}}v_{j_{2}}\dots v_{j_{m}}\,.$}
$S_{m}$ absolutely converges by Proposition 2.6 and has sum $(V(1)_{2}-1)^{m}$. By Propositions 2.2 and 2.4, the series $S$ has sum
$1+\sum_{m\geq 1}(-1)^{m}(V(1)_{2}-1)^{m}=\frac{1}{V(1)_{2}}\,.$}
On the other hand, in $\mathbb{C}[[x]]$ we have the identity
$U(x)=1+\sum_{m\geq 1}(-1)^{m}(v_{1}x+v_{2}x^{2}+\dots)^{m}\,.$}
Thus $u_{0}=1$ and for every $n\in\mathbb{N}$,
$u_{n}=\sum_{m\geq 1}(-1)^{m}\sum_{\begin{subarray}{c}j_{m,1},\,j_{m,2},\,\dots,\,j_{m,m}\in\mathbb{N}\\ j_{m,1}+j_{m,2}+\dots+j_{m,m}=n\end{subarray}}v_{j_{m,1}}v_{j_{m,2}}\dots v_{j_{m,m}}\,.$}
Proposition 2.4 shows that $S$ has sum $u_{0}+u_{1}+u_{2}+\cdots=U(1)_{2}$. The stated formula follows.
2. Indeed, the sum
${\textstyle\sum_{k=0}^{n}|u_{k}|=}$
${\textstyle=1+\sum_{k=1}^{n}\big|\sum_{m\geq 1}(-1)^{m}\sum_{\begin{subarray}{c}j_{m,1},\,j_{m,2},\,\dots,\,j_{m,m}\in\mathbb{N}\\ j_{m,1}+j_{m,2}+\dots+j_{m,m}=k\end{subarray}}v_{j_{m,1}}v_{j_{m,2}}\dots v_{j_{m,m}}\big|}$
${\textstyle\leq\sum_{m\geq 0}(|v_{1}|+|v_{2}|+\dots)^{m}=\frac{1}{1-c}\,.}$
$\Box$
Again, Proposition 2.21 and item 1 of Proposition 2.27 are incomparable.
The analog of Proposition 2.22 for absolute convergence requires two lemmas.
Lemma 2.28
Let $V(x)\in\mathbb{C}[[x]]$ and
$V(x)^{2}=tx^{2l}(1+t_{1}x+t_{2}x^{2}+\dots)$
where $l\in\mathbb{N}_{0}$ and $t,t_{j}\in\mathbb{C}$ — every square in $\mathbb{C}[[x]]$ has this form. Then there exists a number $v\in\mathbb{C}$ such that $v^{2}=t$ and
$V(x)=vx^{l}\sum_{n\geq 0}\binom{1/2}{n}(t_{1}x+t_{2}x^{2}+\dots)^{n}\,.$}
Proof. Let $n\in\mathbb{N}_{0}$ and a,ba,\,b be formal variables. Vandermonde’s identity [31] states that
$\sum_{j=0}^{n}\binom{a}{j}\binom{b}{n-j}=\binom{a+b}{n}\ \ (\in\mathbb{Q}[a,\,b])\,.$}
Let $V_{0}(x)=\sum_{n\geq 0}\binom{1/2}{n}(t_{1}x+t_{2}x^{2}+\dots)^{n}$. For any $v_{0}\in\mathbb{C}$ such that $(v_{0})^{2}=t$ the instance $a=b=\frac{1}{2}$ of Vandermonde’s identity gives
$\big(v_{0}x^{l}V_{0}(x)\big)^{2}$ $=$ ${\textstyle tx^{2l}\sum_{n\geq 0}\big(\sum_{j=0}^{n}\binom{1/2}{j}\binom{1/2}{n-j}\big)(t_{1}x+t_{2}x^{2}+\dots)^{n}}$
$=$ ${\textstyle tx^{2l}\sum_{n\geq 0}\binom{1}{n}(t_{1}x+t_{2}x^{2}+\dots)^{n}}$
$=$ $tx^{2l}(1+t_{1}x+t_{2}x^{2}+\dots)=V(x)^{2}\,.$
Hence $V(x)=\pm v_{0}x^{l}V_{0}(x)$. $\Box$
Lemma 2.29
Let $U(x),V(x)\in\mathbb{C}[[x]]$ and
$U(x)=V(x)^{2}=tx^{2l}(1+t_{1}x+t_{2}x^{2}+\dots)$
where $l\in\mathbb{N}_{0}$ and $t,t_{j}\in\mathbb{C}$. If $c=|t_{1}|+|t_{2}|+\dots<1$, then $U(1)_{2}$ and $V(1)_{2}$ exist, and
$U(1)_{2}=\big(V(1)_{2}\big)^{2}\,.$
Proof. It follows from the assumption on $c$ that $U(1)_{2}$ exists. By Lemma 2.28 there is a number $v\in\mathbb{C}$ such that $v^{2}=t$ and
$V(x)=\sum_{n\geq 0}v_{n}x^{n}=vx^{l}\sum_{n\geq 0}\binom{1/2}{n}(t_{1}x+t_{2}x^{2}+\dots)^{n}\,.$}
Thus $v_{n}=0$ for $n<l$, $v_{l}=v$ and for $n>l$,
$v_{n}=v\sum_{m\geq 1}\binom{1/2}{m}\sum_{\begin{subarray}{c}j_{m,1},\,j_{m,2},\,\dots,\,j_{m,m}\in\mathbb{N}\\ j_{m,1}+j_{m,2}+\dots+j_{m,m}=n-l\end{subarray}}t_{j_{m,1}}t_{j_{m,2}}\dots t_{j_{m,m}}\,.$}
We consider the series
$S=0+0+\dots+0+v+\sum_{\begin{subarray}{c}m\in\mathbb{N}\\ j_{m,1},\,j_{m,2},\,\dots,\,j_{m,m}\in\mathbb{N}\end{subarray}}v\binom{1/2}{m}t_{j_{m,1}}t_{j_{m,2}}\dots t_{j_{m,m}}\,$,}
beginning with $l$ zero summands. We show that it absolutely converges. Let $T$ be a finite subseries of $S$. Since $\big|\binom{1/2}{m}\big|\leq 1$ for every $m\in\mathbb{N}$, the sum of absolute values of summands in $T$ is at most
$|v|+|v|\sum_{m\geq 1}(|t_{1}|+|t_{2}|+\dots)^{m}=|v|+|v|\sum_{m\geq 1}c^{m}=\frac{|v|}{1-c}\,.$}
Thus $S$ absolutely converges. By Proposition 2.4, the series $S$ has sum $v_{0}+v_{1}+\dots=V(1)_{2}$. By item 1 of Proposition 2.24, $U(1)_{2}=\big(V(1)_{2}\big)^{2}$. $\Box$
Proposition 2.30
Let $U(x),V(x)=\sum_{n\geq 0}v_{n}x^{n}$ and $W(x)=\sum_{n\geq 0}w_{n}x^{n}$ be in $\mathbb{C}[[x]]$. Suppose that $V(1)_{2}$ exists, $w_{0}=1$, $U(x)^{2}=1-\frac{V(x)}{W(x)}$ and that $c=|w_{1}|+|w_{2}|+\dots<1$. If $V(x)=W(x)$, then $U(x)=0$. Else, if $k\in\mathbb{N}_{0}$ is minimum such that $v_{k}\neq w_{k}$ and if
$\big(|w_{k}-v_{k}|\cdot(1-c)\big)^{-1}\sum_{n>k}|w_{n}-v_{n}|<1\,$,}
then $U(1)_{2}$ and $W(1)_{2}$ exist, $W(1)_{2}\neq 0$ and
$\big(U(1)_{2}\big)^{2}=1-\frac{V(1)_{2}}{W(1)_{2}}\,.$}
Proof. If $V(x)=W(x)$ then $U(x)^{2}=0$ and $U(x)=0$. Let $V(x)\neq W(x)$ and $k\in\mathbb{N}_{0}$ be as stated. We write
$U(x)^{2}=\frac{W(x)-V(x)}{W(x)}=tx^{2l}(1+t_{1}x+t_{2}x^{2}+\dots)$}
where $l\in\mathbb{N}_{0}$, $t=w_{k}-v_{k}\in\mathbb{C}^{*}$ and $t_{j}\in\mathbb{C}$. Let
$U_{0}(x)=\frac{1}{W(x)}=\sum_{n\geq 0}u_{0,n}x^{n},\ U_{1}(x)=W(x)-V(x)$}
and $U_{2}(x)=U_{0}(x)U_{1}(x)=\sum_{n\geq 0}u_{2,n}x^{n}$. By item 1 of Proposition 2.27, $U_{0}(1)_{2}$ and $W(1)_{2}$ exist, $W(1)_{2}\neq 0$ and $U_{0}(1)_{2}=\frac{1}{W(1)_{2}}$. By Proposition 2.23, $U_{1}(1)_{2}$ exists and $U_{1}(1)_{2}=W(1)_{2}-V(1)_{2}$. By item 1 of Proposition 2.24, $U_{2}(1)_{2}$ exists and $U_{2}(1)_{2}=U_{0}(1)_{2}\cdot U_{1}(1)_{2}$. By item 2 of Proposition 2.27, $\sum_{n\geq 0}|u_{0,n}|\leq\frac{1}{1-c}$. By item 2 of Proposition 2.24,
$\sum_{n\geq 0}|u_{2,n}|\leq\sum_{n\geq 0}|u_{0,n}|\cdot\sum_{n>k}|w_{n}-v_{n}|\leq\frac{1}{1-c}\sum_{n>k}|w_{n}-v_{n}|\,.$}
Since $U(x)^{2}=U_{2}(x)$, we have bound
$|t_{1}|+|t_{2}|+\dots\leq\frac{1}{|t|(1-c)}\sum_{n>k}|w_{n}-v_{n}|<1\,.$}
By Lemma 2.29, $U(1)_{2}$ exists and
$\big(U(1)_{2}\big)^{2}=U_{2}(1)_{2}=U_{0}(1)_{2}\cdot U_{1}(1)_{2}=1-\frac{V(1)_{2}}{W(1)_{2}}\,.$}
$\Box$
2.9 Generating functions of walks for $\xi$-convex weights
We generalize the result in Section 1.2 that $h(D_{n})=1$ for every $n$.
Proposition 2.31
Let $\xi\in\mathbb{C}$. For every $\xi$-convex light weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ we have
$\sum_{n\geq 0}h(D_{n})x^{n}=\frac{1}{1-\xi x}\ \ (\in\mathbb{C}[[x]])\,.$}
Thus for $\xi=0$ we have $\sum_{n\geq 0}h(D_{n})x^{n}=1$ ($1x^{0}+0x^{1}+0x^{2}+\dots$).
Proof. Let $\xi\in\mathbb{C}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light and $\xi$-convex weight. We prove by induction on $n\in\mathbb{N}_{0}$ that $h(D_{n})=\xi^{n}$. For $n=0$ this holds, $h(D_{0})=1$, because $D_{0}=\{\langle 1\rangle\}$. Let $n>0$. Recall that $\ell(w)$ denotes the last vertex of a walk $w$. Using Propositions 2.2 and 2.4, the assumption of $\xi$-convexity and induction, we get
$h(D_{n})$ $=$ ${\textstyle\sum_{\begin{subarray}{c}w\in D_{n-1}\\ \ell(w)\in V(h)\end{subarray}}h(w)\cdot\sum_{v\in\mathbb{N}\setminus\{\ell(w)\}}h(\{\ell(w),\,v\})}$
$=$ ${\textstyle\sum_{w\in D_{n-1}}h(w)\cdot\xi=h(D_{n-1})\cdot\xi=\xi^{n-1}\cdot\xi=\xi^{n}\,.}$
We can add and remove the condition $\ell(w)\in V(h)$ without affecting the sum because walks $w$ not satisfying it have zero weight. $\Box$
3 General Pólya’s theorems for $v=1$
In this section, we generalize the results in Section 1.2 to weights $h\colon\mathbb{N}_{2}\to\mathbb{C}$. We first introduce three auxiliary generating functions $B_{h}(x)$, $C_{h}(x)$ and $D_{h}(x)$. In Sections 3.2 and 3.3, we obtain general Pólya’s theorems for $v=1$ with general and $\xi$-convex weights, respectively.
3.1 Auxiliary generating functions
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight (Definition 2.1). As we know from Section 1.4, we are interested in the main generating function
$A_{h}(x)=A_{1,\,h}(x)=\sum_{n\geq 0}h(A_{n})x^{n}\ \ (\in\mathbb{C}[[x]])$}
for the set $A$ of walks in $K_{\mathbb{N}}$ starting at $1$ and revisiting it later; $h(A_{n})$ is the sum of the series
$\sum_{w\in A_{n}}h(w)=\sum_{\begin{subarray}{c}w\in A\\ s(w)=n\end{subarray}}h(w)\,.$}
As in Section 1.2, we consider the set $B$ of closed walks in $K_{\mathbb{N}}$ starting and ending at $1$, the set $C\subset B$ of closed walks with inner vertices different from $1$, and the set $D$ of all walks in $K_{\mathbb{N}}$ starting at $1$. We define three auxiliary generating functions
$B_{h}(x)=\sum_{n\geq 0}h(B_{n})x^{n},\ C_{h}(x)=\sum_{n\geq 0}h(C_{n})x^{n},\ D_{h}(x)=\sum_{n\geq 0}h(D_{n})x^{n}\,.$}
The coefficients $h(A_{n})$, $h(B_{n})$ and $h(C_{n})$, $n\in\mathbb{N}_{0}$, are correctly defined because we assume that the coefficients $h(D_{n})$ are defined ($h$ is a light weight). We have $h(A_{0})=h(C_{0})=0$ and $h(B_{0})=h(D_{0})=1$. Let $F\colon X\to Y$ be a map and $Z$ be any set. In the next proof we use the notation
$F[Z]=\{F(x)\colon\;x\in X\cap Z\}\ \ (\subset Y)\,.$
Proposition 3.1
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight and $A_{h}(x)$, $B_{h}(x)$, $C_{h}(x)$ and $D_{h}(x)$ be the above generating functions. The following identities hold in $\mathbb{C}[[x]]$.
  1. 1.
    $A_{h}(x)=C_{h}(x)D_{h}(x)$.
  2. 2.
    $B_{h}(x)=\frac{1}{1-C_{h}(x)}$, equivalently, $C_{h}(x)=1-\frac{1}{B_{h}(x)}$.
Proof. 1. We show that $h(A_{n})=\sum_{j=0}^{n}h(C_{j})\cdot h(D_{n-j})$ for every $n\in\mathbb{N}_{0}$. For $n\leq 1$ it holds because then $h(A_{n})=h(C_{n})=0$. For $n\geq 2$ the set $A_{n}$ is infinite and countable. We split every walk in it at the first revisit of $1$ and get a map
$F\colon A_{n}\to\bigcup_{j=0}^{n}C_{j}\times D_{n-j}$}
that is defined as follows. For any walk
$w=\langle u_{0},\,u_{1},\,\dots,\,u_{j},\,\dots,\,u_{n}\rangle\in A_{n}\,$,
where $u_{0}=u_{j}=1$, $j>0$ and $u_{i}\neq 1$ for every $i=1,2,\dots,j-1$, we set
$F(w)=\langle w_{1},\,w_{2}\rangle\ \ (\in C_{j}\times D_{n-j})\,$,
where $w_{1}=\langle u_{0},u_{1},\dots,u_{j}\rangle$ and $w_{2}=\langle u_{j},u_{j+1},\dots,u_{n}\rangle$. It is easy to see that $F$ is a bijection. $F$ is also weight-preserving (WP): for every weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and value $F(w)=\langle w_{1},w_{2}\rangle$, we have $h(w)=h(w_{1})h(w_{2})$. If $X$ is a set of pairs of walks in $K_{\mathbb{N}}$, we denote by $h(X)$ the sum of the series
$\sum_{\langle w,\,w^{\prime}\rangle\in X}h(w)h(w^{\prime})\,$,}
if the series absolutely converges.
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight and let $n\geq 2$. Using the map $F$, we get
$h(A_{n})$ $\stackrel{{\scriptstyle\mathrm{WP}}}{{=}}$ $h(F[A_{n}])$
[…]
2. We show that
$B_{h}(x)=\frac{1}{1-C_{h}(x)}=1+\sum_{j\geq 1}(C_{h}(x))^{j}\ \ (\in\mathbb{C}[[x]])\,.$}
We recall that $h(B_{0})=1$ and show that for every $n\in\mathbb{N}$,
$h(B_{n})=\sum_{j\geq 1}\sum_{\begin{subarray}{c}n_{1},\,n_{2},\,\dots,\,n_{j}\in\mathbb{N}\\ n_{1}+n_{2}+\dots+n_{j}=n\end{subarray}}h(C_{n_{1}})h(C_{n_{2}})\dots h(C_{n_{j}})\,.$}
For $n=1$ it holds because $h(B_{1})=h(C_{1})=0$. For $n\geq 2$ the set $B_{n}$ is infinite and countable. We split every walk in it at the visits of $1$ and get a map
$F\colon B_{n}\to\bigcup_{j\geq 1}\bigcup_{\begin{subarray}{c}n_{1},\,n_{2},\,\dots,\,n_{j}\in\mathbb{N}\\ n_{1}+n_{2}+\dots+n_{j}=n\end{subarray}}C_{n_{1}}\times C_{n_{2}}\times\cdots\times C_{n_{j}}$}
that is defined as follows. For any walk
$w=\langle u_{0},\,\dots,\,u_{m_{1}},\,\dots,\,\dots,\,u_{m_{j-1}},\,\dots,\,u_{n}\rangle\in B_{n}\,$,
where $0=m_{0}<m_{1}<\dots<m_{j}=n$ and $u_{k}=1$ iff $k\in\{0,m_{1},m_{2},\dots,m_{j}\}$, we set
$F(w)=\langle w_{1},\,w_{2},\,\dots,\,w_{j}\rangle\ \ (\in C_{n_{1}}\times C_{n_{2}}\times\dots\times C_{n_{j}},\,n_{i}=m_{i}-m_{i-1})\,$,
where $w_{i}=\langle u_{m_{i-1}},u_{m_{i-1}+1},\dots,u_{m_{i}}\rangle$, $i=1,2,\dots,j$. It is easy to see that $F$ is a bijection and that it is weight-preserving (WP), as in part 1. If $X$ is a set of $k$-tuples of walks, we denote by $h(X)$ the sum of the series
$\sum_{\langle w_{1},\,w_{2},\,\dots,\,w_{k})\in X}h(w_{1})h(w_{2})\dots h(w_{k})\,$,}
if the series absolutely converges.
Let $h$ be a light weight and $n\geq 2$. Using the map $F$, we get
$h(B_{n})$ $\stackrel{{\scriptstyle\mathrm{WP}}}{{=}}$ $h(F[B_{n}])$
[…]
$\Box$
We already mentioned in Section 1.5 that the previous proof differs from all related proofs in the first phase of the symbolic method because we work with infinite sets of combinatorial objects,
3.2 Pólya’s theorems for $v=1$ and general weights
We obtain general Pólya’s theorems of the form indicated in the subsection heading. For the labels $\exists\,U(1)_{1}$ and $\exists\,U(1)_{2}$ of the theorems see Definition 2.12, and for the label $U(1)_{1}=\infty$ see Definition 2.14. The generating functions $A_{h}(x)$, $B_{h}(x)$, $C_{h}(x)$, and $D_{h}(x)$ are defined in the previous subsection.
Theorem 3.2 ($\exists\,U(1)_{1}$)
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. If $A_{h}(1)_{1}$, $B_{h}(1)_{1}$, $C_{h}(1)_{1}$ and $D_{h}(1)_{1}$ exist, then $B_{h}(1)_{1}\neq 0$ and
$A_{h}(1)_{1}=\Big(1-\frac{1}{B_{h}(1)_{1}}\Big)D_{h}(1)_{1}\,.$
Proof. By item 1 of Proposition 3.1 we have $A_{h}(x)=C_{h}(x)D_{h}(x)$. By Proposition 2.20, $A_{h}(1)_{1}=C_{h}(1)_{1}D_{h}(1)_{1}$. By item 2 of Proposition 3.1 we have $1-C_{h}(x)=\frac{1}{B_{h}(x)}$. By Proposition 2.21 we have $B_{h}(1)_{1}\neq 0$ and
$(1-C_{h})(1)_{1}=\frac{1}{B_{h}(1)_{1}},\,\text{ so that }\,C_{h}(1)_{1}=1-\frac{1}{B_{h}(1)_{1}}\,.$}
The stated formula follows. $\Box$
Theorem 3.3 ($\exists\,U(1)_{2}$)
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. If $\sum_{n\geq 1}|h(B_{n})|<1$ and $D_{h}(1)_{2}$ exists, then $B_{h}(1)_{2}\neq 0$ and
$A_{h}(1)_{2}=\Big(1-\frac{1}{B_{h}(1)_{2}}\Big)D_{h}(1)_{2}\,.$
Proof. By item 2 of Proposition 3.1 we have $1-C_{h}(x)=\frac{1}{B_{h}(x)}$. By item 1 of Proposition 2.27, $B_{h}(1)_{2}\neq 0$ and
$(1-C_{h})(1)_{2}=\frac{1}{B_{h}(1)_{2}},\,\text{ so that }\,C_{h}(1)_{2}=1-\frac{1}{B_{h}(1)_{2}}\,.$}
By item 1 of Proposition 3.1 we have $A_{h}(x)=C_{h}(x)D_{h}(x)$. By item 1 of Proposition 2.24, $A_{h}(1)_{2}=C_{h}(1)_{2}D_{h}(1)_{2}$ . The stated formula follows. $\Box$
Theorem 3.4 ($U(1)_{1}=\infty$)
Let $\xi\in\mathbb{C}_{1}$ and $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ be a light weight. If $(D_{h})_{1/\xi}(1)_{1}=+\infty$, then
$(A_{h})_{1/\xi}(1)_{1}=+\infty\,.$
Proof. From $(D_{h})_{1/\xi}(1)_{1}=+\infty$ it follows that $\xi^{-n}h(D_{n})>0$ for some $n$ in $\mathbb{N}$. Hence there is a vertex $u\in\mathbb{N}\setminus\{1\}$ such that $\xi^{-1}h(\{1,u\})>0$. The walk $\langle 1,u,1\rangle$ shows by the subseries bound that $\xi^{-2}h(C_{2})>0$. Since by item 1 of Proposition 3.1 we have $h(A_{n})=\sum_{j=0}^{n}h(C_{j})h(D_{n-j})$, we get that
$\xi^{-2-n}h(A_{n+2})\geq\xi^{-2}h(C_{2})\cdot\xi^{-n}h(D_{n})$
for every $n\in\mathbb{N}_{0}$. Hence
$\sum_{j=0}^{n}\xi^{-2-j}h(A_{j+2})\geq\xi^{-2}h(C_{2})\cdot\sum_{j=0}^{n}\xi^{-j}h(D_{j})\to+\infty$}
for $n\to\infty$, and $(A_{h})_{1/\xi}(1)_{1}=+\infty$. $\Box$
3.3 Pólya’s theorems for $v=1$ and $\xi$-convex weights
We proceed to $\xi$-convex weights and begin with twisted nonnegative ones. The next theorem generalizes Section 1.2.
Theorem 3.5 ($\exists\,U(1)_{1}$ and $h\geq 0$)
Let $\xi\in\mathbb{C}_{1}$ and $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ be a $\xi$-convex light weight. Then, with $(B_{h})_{1/\xi}(1)_{1}\in[1,+\infty)\cup\{+\infty\}$, we have
$\lim_{n\to\infty}\xi^{-n}h(A_{n})=1-\frac{1}{(B_{h})_{1/\xi}(1)_{1}}\,.$
For $(B_{h})_{1/\xi}(1)_{1}=+\infty$ the limit is $1$, and else it lies in $(0,1)$.
Proof. Proposition 2.31 gives $D_{h}(x)=\frac{1}{1-\xi x}$. Since $B_{n},C_{n}\subset D_{n}$ for every $n\in\mathbb{N}_{0}$, by the subseries bound we have $\xi^{-n}h(B_{n}),\xi^{-n}h(C_{n})\in[0,1]$. Now we argue as in Section 1.2. The last claim that $(B_{h})_{1/\xi}(1)_{1}>1$ follows from the assumption of $\xi$-convexity: there is a $u\in\mathbb{N}\setminus\{1\}$ such that $\xi^{-1}h(\{1,u\})>0$, and hence $\xi^{-2}h(B_{2})>0$. $\Box$
Theorem 3.6 ($\exists\,U(1)_{1}$)
1. Let $\xi\in\mathbb{C}_{1}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $\xi$-convex light weight. If $(B_{h})_{1/\xi}(1)_{1}$ and $(C_{h})_{1/\xi}(1)_{1}$ exist, then $(B_{h})_{1/\xi}(1)_{1}\neq 0$ and
$\lim_{n\to\infty}\xi^{-n}h(A_{n})=1-\frac{1}{(B_{h})_{1/\xi}(1)_{1}}\,.$
2. Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a 0-convex light weight. If $B_{h}(1)_{1}$ and $C_{h}(1)_{1}$ exist, then $B_{h}(1)_{1}\neq 0$ and
$A_{h}(1)_{1}=1-\frac{1}{B_{h}(1)_{1}}\,.$
Proof. 1. By item 1 of Proposition 3.1 and by Proposition 2.31 we have $A_{h}(x)=C_{h}(x)D_{h}(x)=C_{h}(x)\frac{1}{1-\xi x}$. By Proposition 2.15,
$\lim_{n\to\infty}\xi^{-n}h(A_{n})=(C_{h})_{1/\xi}(1)_{1}\,.$
By item 2 of Proposition 3.1 we have $1-C_{h}(x)=\frac{1}{B_{h}(x)}$. By Proposition 2.21 we have $(B_{h})_{1/\xi}(1)_{1}\neq 0$ and
[…]
The stated formula follows.
2. By item 1 of Proposition 3.1 and by Proposition 2.31, $A_{h}(x)=C_{h}(x)$. We use the displayed formula for $(C_{h})_{1/\xi}(1)_{1}$ with $\xi=1$. $\Box$
This theorem generalizes the case $d\geq 3$ of the formula (1) for $\overline{v}=\overline{0}$. The case $d\leq 2$ is generalized in Theorem 3.8.
Theorem 3.7 ($\exists\,U(1)_{2}$)
1. Let $\xi\in\mathbb{C}_{1}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $\xi$-convex light weight. If $\sum_{n\geq 1}\big|h(B_{n})\big|<1$, then $(B_{h})_{1/\xi}(1)_{2}\neq 0$ and
$\lim_{n\to\infty}\xi^{-n}h(A_{n})=1-\frac{1}{(B_{h})_{1/\xi}(1)_{2}}\,.$
2. Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a 0-convex light weight. If $\sum_{n\geq 1}|h(B_{n})|<1$, then $B_{h}(1)_{2}\neq 0$ and
$A_{h}(1)_{2}=1-\frac{1}{B_{h}(1)_{2}}\,.$
Proof. 1. By item 2 of Proposition 3.1 we have $1-C_{h}(x)=\frac{1}{B_{h}(x)}$. By item 1 of Proposition 2.27, $(B_{h})_{1/\xi}(1)_{2}\neq 0$ and
[…]
By item 1 of Proposition 3.1 and Proposition 2.31, $A_{h}(x)=C_{h}(x)D_{h}(x)=C_{h}(x)\frac{1}{1-\xi x}$. By Proposition 2.15
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h}=(C_{h})_{1/\xi}(1)_{1}=(C_{h})_{1/\xi}(1)_{2}=1-\frac{1}{(B_{h})_{1/\xi}(1)_{2}}\,.$}
2. By item 1 of Proposition 3.1 and by Proposition 2.31, $A_{h}(x)=C_{h}(x)$. We use the displayed formula for $(C_{h})_{1/\xi}(1)_{2}$ with $\xi=1$. $\Box$
Theorem 3.8 ($U(1)_{1}=\infty$)
1. Let $\xi\in\mathbb{C}_{1}$ and $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ be a $\xi$-convex light weight. If $(B_{h})_{1/\xi}(1)_{1}=+\infty$, then
$\lim_{n\to\infty}\xi^{-n}h(A_{n})=1\,.$
2. Let $\xi\in\mathbb{C}_{1}$ and $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}$ be a 0-convex light weight. If $B_{h}(x),C_{h}(x)\in H_{1}$ and if $(B_{h})_{1/\xi}(1)_{1}=\pm\infty$, then
$\lim_{x\nearrow 1}(A_{h})_{1/\xi}(x)=1\,.$
Proof. 1. We proved this in Theorem 3.5.
2. By item 1 of Proposition 3.1 and by Proposition 2.31, we have $A_{h}(x)=C_{h}(x)$. By item 2 of Proposition 3.1 we have $C_{h}(x)=1-\frac{1}{B_{h}(x)}$. By Theorem 2.17,
$\lim_{x\nearrow 1}(B_{h})_{1/\xi}(x)=(B_{h})_{1/\xi}(1)_{1}=\pm\infty\,.$
It follows from this and the assumptions that
$\text{$(C_{h})_{1/\xi}(x)=1-\frac{1}{(B_{h})_{1/\xi}(x)}$ for every $x\in(1-\delta,\,1)$}\,.$}
We compute that
$\lim_{x\nearrow 1}(C_{h})_{1/\xi}(x)={\textstyle 1-\frac{1}{\lim_{x\nearrow 1}(B_{h})_{1/\xi}(x)}=1-\frac{1}{(B_{h})_{1/\xi}(1)_{1}}=1-\frac{1}{\pm\infty}=1\,.}$
In the second equality we used Theorem 2.17. The stated formula follows. $\Box$
4 General Pólya’s theorems for $v\neq 1$
In this last section, we generalize the results in Section 1.3 to weights $h\colon\mathbb{N}_{2}\to\mathbb{C}$. We first introduce another three auxiliary generating functions $B_{v,h}(x)$, $C_{v,h}(x)$, and $E_{v,h}(x)$. The second identity in Proposition 4.2 contains a square, and in Theorem 4.4 we deduce a result on monotonicity from this. In Sections 4.2 and 4.3, we obtain general Pólya’s theorems for $v\neq 1$ with general and $\xi$-convex weights, respectively.
4.1 Auxiliary generating functions
Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. As we know from Section 1.4, we are interested in the main generating function
$A_{v,\,h}(x)=\sum_{n\geq 0}h(A_{v,\,n})x^{n}\ \ (\in\mathbb{C}[[x]])$}
for the set $A_{v}$ of walks in $K_{\mathbb{N}}$ starting at $1$ and visiting $v$; $h(A_{v,n})$ is the sum of the series
$\sum_{w\in A_{v,\,n}}h(w)=\sum_{\begin{subarray}{c}w\in A_{v}\\ s(w)=n\end{subarray}}h(w)\,.$}
Besides the sets $A_{v}$, $B$, and $D$ of walks in $K_{\mathbb{N}}$, we use the set $B_{v}$ of walks in $K_{\mathbb{N}}$ starting and ending at $1$ and avoiding $v$, the set $C_{v}$ of walks in $K_{\mathbb{N}}$ starting at $1$, ending at $v$ and with the inner vertices different from $v$, and the set $E_{v}$ of walks in $K_{\mathbb{N}}$ starting at $1$, ending at $v$ and with the inner vertices different from both $1$ and $v$. We define three additional auxiliary generating functions
$B_{v,\,h}(x)$ $=$ ${\textstyle\sum_{n\geq 0}h(B_{v,\,n})x^{n},\ C_{v,\,h}(x)=\sum_{n\geq 0}h(C_{v,\,n})x^{n}\,\text{ and}}$
$E_{v,\,h(x)}$ $=$ ${\textstyle\sum_{n\geq 0}h(E_{v,\,n})x^{n}\,.}$
We have $h(A_{v,0})=h(C_{v,0})=h(E_{v,0})=0$ and $h(B_{0})=h(B_{v,0})=h(D_{0})=1$.
In the case $\overline{v}=\overline{0}$ the symmetry of the grid graph $G_{d}=\langle\mathbb{Z}^{d},E_{d}\rangle$, namely its vertex-transitivity, played no role, but in the case $\overline{v}\neq\overline{0}$ this symmetry is crucial; it ensures that the enumeration of walks does not depend on the starting vertex. The same holds for the general weighted case, and we introduce the weighted version of vertex transitivity.
Definition 4.1 ($v$-transitivity)
Let $v\in\mathbb{N}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$. We say that the weight $h$ is $v$-transitive if there exists a bijection $f\colon\mathbb{N}\to\mathbb{N}$ such that
$f(1)=v\,\text{ and }\,\forall e\in\mathbb{N}_{2}:\;h(e)=h(f[e])\,.$
We obtain an analog of Proposition 3.1. The proofs are similar, and we only sketch them.
Proposition 4.2
Let $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight, and $A_{v,h}(x)$, $B_{h}(x)$, $B_{v,h}(x)$, $C_{v,h}(x)$, $D_{h}(x)$, and $E_{v,h}(x)$ be the above generating functions. The following identities hold in $\mathbb{C}[[x]]$.
  1. 1.
    $A_{v,h}(x)=C_{v,h}(x)D_{h}(x)$.
  2. 2.
    $B_{h}(x)=B_{v,h}(x)+C_{v,h}(x)^{2}B_{h}(x)$, equivalently, $C_{v,h}(x)^{2}=1-\frac{B_{v,h}(x)}{B_{h}(x)}$.
  3. 3.
    $C_{v,h}(x)=B_{v,h}(x)E_{v,h}(x)$.
Proof. 1. We split every walk $w\in A_{v}$ at the first visit of $v$ as $w=w_{1}w_{2}$. This splitting yields, when the $v$-transitivity of $h$ is applied to $w_{2}$, the stated identity.
2. It suffices to prove the first identity. We have the partition
$B=B_{v}\cup B_{v}^{\prime}$
where the set $B_{v}^{\prime}$ consists of the walks in $K_{\mathbb{N}}$ that start and end at $1$ and visit $v$. We split every walk $w\in B_{v}^{\prime}$ at the first and last visits of $v$ as $w=w_{1}w_{2}w_{3}$. The above partition, the reversal of walks (applied to $w_{3}$), and the $v$-transitivity of $h$ (applied to $w_{2}$) yield the stated identity.
3. We split every walk $w\in C_{v}$ at the last visit of $1$ as $w=w_{1}w_{2}$. This splitting yields, without help of the $v$-transitivity of $h$, the stated identity. $\Box$
One would expect to see for $v\neq 1$ the generating function of the set $F_{v}$ of walks in $K_{\mathbb{N}}$ going from $1$ to $v$. We managed for a long time without it, but at last we include two identities involving $F_{v}$. So we define another auxiliary generating function
$F_{v,\,h}(x)=\sum_{n\geq 0}h(F_{v,\,n})x^{n}\,.$}
Proposition 4.3
Let $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight, and $B_{h}(x)$, $B_{v,h}(x)$, $C_{v,h}(x)$, and $F_{v,h}(x)$ be the above generating functions. The following identities hold in $\mathbb{C}[[x]]$.
  1. 1.
    $F_{v,h}(x)=C_{v,h}(x)B_{h}(x)$.
  2. 2.
    $B_{v,h}(x)=B_{h}(x)-\frac{F_{v,h}(x)^{2}}{B_{h}(x)}$.
Proof. 1. We split every walk $w\in F_{v}$ at the first visit of $v$ as $w=w_{1}w_{2}$. This splitting yields, when the $v$-transitivity of $h$ is applied to $w_{2}$, the identity.
$F_{v,\,h}(x)=C_{v,\,h}(x)+C_{v,\,h}(x)(B_{h}(x)-1)=C_{v,\,h}(x)B_{h}(x)\,.$
2. Using the reversal of walks and the $v$-transitivity of $h$, we get
$B_{h}(x)-\frac{F_{v,\,h}(x)^{2}}{B_{h}(x)}=B_{h}(x)-F_{v,\,h}(x)^{2}\sum_{j\geq 1}(-1)^{j-1}(B_{h}(x)-1)^{j-1}=B_{v,\,h}(x)\,.$}
Indeed, if $w$ is a walk in $K_{\mathbb{N}}$ from $1$ to $1$ with $k\in\mathbb{N}_{0}$ visits of the vertex $v$, then $w$ is counted exactly
$1-\sum_{j=1}^{k}\binom{k}{j}(-1)^{j-1}=\sum_{j=0}^{k}\binom{k}{j}(-1)^{j}=(1-1)^{k}\ \ (\in\{0,\,1\})$}
times in the left-hand side $\cdots$ of the displayed equality $\dots=B_{v,h}(x)$, which agrees with the count of $w$ in the right-hand side $B_{v,h}(x)$. $\Box$
Item 2 in the previous proposition and item 1 in this one provide two expressions for $C_{v,h}(x)$:
$C_{v,h}(x)^{2}=1-\frac{B_{v,\,h}(x)}{B_{h}(x)}\,\text{ and }\,C_{v,h}(x)=\frac{F_{v,\,h}(x)}{B_{h}(x)}\,.$}
Although the latter expression is simpler than the former, both in appearance and factually ($F_{v}$ has simpler definition than $B_{v}$), we use the former: due to inclusions $B_{v,n}\subset B_{n}$ we easily show the membership in $(0,1)$ in Theorem 4.9. In the simple latter expression we have disjoint sets $F_{v,n}$ and $B_{n}$, and it is not clear how to deduce that $(F_{v,h})_{1/\xi}(1)_{1}<(B_{h})_{1/\xi}(1)_{1}$.
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and $v\in\mathbb{N}\setminus\{1\}$. If $v\not\in V(h)$ then $h(w)=0$ for every walk $w\in A_{v}$ because every walk starting at $1$ and visiting $v$ contains an edge with zero weight. Henceforth, we therefore always assume that $v\in V(h)$. For $w,z\in\mathbb{C}$, we define the set
$w\cdot\mathrm{sqrt}(z)=\{w\alpha\colon\;\alpha\in\mathbb{C},\,\alpha^{2}=z\}\ \ (\subset\mathbb{C})\,.$
For $w=0$ or $z=0$ it equals $\{0\}$. Else it has two non-zero elements differing only by sign. If $w=1$, we write just $\mathrm{sqrt}(z)$.
Before we proceed to general Pólya’s theorems for $v\neq 1$, we utilize the square in item 2 of Proposition 4.2. If $h\colon X\to\mathbb{R}$ is an absolutely convergent real series and $Y\subset X$, then the sum $h(Y)$ of the subseries need not be smaller than the sum $h(X)$. However, in the case of inclusions $B_{v,n}\subset B_{n}$, sums $h(B_{v,n})$ are, in a sense, indeed smaller than sums $h(B_{n})$, as shown in the following theorem.
Theorem 4.4
Let $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{R}$ be a $v$-transitive light weight and let $B_{v,h}(x),B_{h}(x),C_{v,h}(x)\in H_{1}$. Then for every number $x\in(-1,1)$ the following holds.
  1. 1.
    If $B_{h}(x)=0$ then also $B_{v,h}(x)=0$.
  2. 2.
    If $B_{h}(x)\neq 0$ then
    $\frac{B_{v,\,h}(x)}{B_{h}(x)}\leq 1\,.$
Proof. By item 2 of Proposition 4.2, the assumptions, and Propositions 2.2, 2.4 and 2.6 we have for every number $x\in(-1,1)$ that
$B_{h}(x)=B_{v,h}(x)+C_{v,h}(x)^{2}B_{h}(x)\,.$
The claim 1 is clear. If $B_{h}(x)\neq 0$ then
$\frac{B_{v,\,h}(x)}{B_{h}(x)}=1-C_{v,\,h}(x)^{2}\leq 1\,.$}
$\Box$
4.2 Pólya’s theorems for $v\neq 1$ and general weights
We obtain general Pólya’s theorems of the form indicated in the subsection heading.
Theorem 4.5 ($\exists\,U(1)_{1}$)
Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight such that $v\in V(h)$. If $A_{v,h}(1)_{1}$, $B_{h}(1)_{1}$, $B_{v,h}(1)_{1}$, $C_{v,h}(1)_{1}$ and $D_{h}(1)_{1}$ exist and $B_{h}(1)_{1}\neq 0$, then
$A_{v,\,h}(1)_{1}\in D_{h}(1)_{1}\cdot\mathrm{sqrt}\Big(1-\frac{B_{v,\,h}(1)_{1}}{B_{h}(1)_{1}}\Big)\,.$
Proof. By item 1 of Proposition 4.2 we have $A_{v,h}(x)=C_{v,h}(x)D_{h}(x)$. By Proposition 2.20, $A_{v,h}(1)_{1}=C_{v,h}(1)_{1}D_{h}(1)_{1}$. By item 2 of Proposition 4.2 we have $C_{v,h}(x)^{2}=1-\frac{B_{v,h}(x)}{B_{h}(x)}$. Thus by Proposition 2.22, $\big(C_{v,h}(1)_{1}\big)^{2}=1-\frac{B_{v,h}(1)_{1}}{B_{h}(1)_{1}}$. The stated formula follows. $\Box$
We trade the non-vanishing of $B_{h}(1)_{1}$ for the existence of $E_{v,h}(1)_{1}$ and obtain the following variant of the last theorem.
Theorem 4.6 ($\exists\,U(1)_{1}$)
Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight such that $v\in V(h)$. If $A_{v,h}(1)_{1}$, $B_{h}(1)_{1}$, $B_{v,h}(1)_{1}$, $C_{v,h}(1)_{1}$, $D_{h}(1)_{1}$ and $E_{v,h}(1)_{1}$ exist, then
$A_{v,\,h}(1)_{1}\in D_{h}(1)_{1}\cdot\mathrm{sqrt}\Big(1-\frac{B_{v,\,h}(1)_{1}}{B_{h}(1)_{1}}\Big)\,$,
where for $B_{h}(1)_{1}=0$ we have $B_{v,h}(1)_{1}=0$ and interpret the expression $\frac{0}{0}$ as $1$, so that then $A_{v,h}(1)_{1}=0$.
Proof. If $B_{h}(1)_{1}\neq 0$, we use the previous proof. If $B_{h}(1)_{1}=0$ then it follows from item 2 of Proposition 4.2, and Propositions 2.20 and 2.19 that also $B_{v,h}(1)_{1}=0$. Item 3 of Proposition 4.2 and Proposition 2.20 give $C_{v,h}(1)_{1}=0$. Finally, item 1 of Proposition 4.2 and Proposition 2.20 give $A_{v,h}(1)_{1}=0$. $\Box$
Theorem 4.7 ($\exists\,U(1)_{2}$)
Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight such that $v\in V(h)$. If
$c=\sum_{n\geq 1}|h(B_{n})|<1$
and $B_{v,h}(1)_{2}$ and $D_{h}(1)_{2}$ exist, then the following holds. If $B_{v,h}(x)=B_{h}(x)$ then $A_{v,h}(x)=0$. If $B_{v,h}(x)\neq B_{h}(x)$, $k\in\mathbb{N}_{0}$ is minimum such that $h(B_{v,k})\neq h(B_{k})$ and if
$\big(|h(B_{v,\,k})-h(B_{k})|\cdot(1-c)\big)^{-1}\sum_{n>k}|h(B_{v,\,n})-h(B_{n})|<1\,$,
then $B_{h}(1)_{2}\neq 0$ and
$A_{v,\,h}(1)_{2}\in D_{h}(1)_{2}\cdot\mathrm{sqrt}\Big(1-\frac{B_{v,\,h}(1)_{2}}{B_{h}(1)_{2}}\Big)\,.$
Proof. If $B_{v,h}(x)=B_{h}(x)$, then by item 2 of Proposition 4.2 we have $C_{v,h}(x)^{2}=1-\frac{B_{v,h}(x)}{B_{h}(x)}=1-1=0$ and $C_{v,h}(x)=0$. Item 1 of Proposition 4.2 gives $A_{v,h}(x)=C_{v,h}(x)D_{h}(x)=0$.
Let $B_{v,h}(x)\neq B_{h}(x)$, $k$ be as stated, and let the latter displayed bound hold. By item 2 of Proposition 4.2, we have $C_{v,h}(x)^{2}=1-\frac{B_{v,h}(x)}{B_{h}(x)}$. By Proposition 2.30, $B_{h}(1)_{2}\neq 0$ and
$\big(C_{v,\,h}(1)_{2}\big)^{2}=1-\frac{B_{v,\,h}(1)_{2}}{B_{h}(1)_{2}}\,.$}
By item 1 of Proposition 4.2 we have $A_{v,h}(x)=C_{v,h}(x)D_{h}(x)$. By item 1 of Proposition 2.24, $A_{v,h}(1)_{2}=C_{v,h}(1)_{2}D_{h}(1)_{2}$. The stated formula follows. $\Box$
Theorem 4.8 ($U(1)_{1}=\infty$)
Let $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ be a $v$-transitive light weight such that $v\in V(h)$. If $(D_{h})_{1/\xi}(1)_{1}=+\infty$, then
$(A_{v,\,h})_{1/\xi}(1)_{1}=+\infty\,.$
Proof. Let $m$ be the minimum length of a walk $w$ joining $1$ and $v$ in $G(h)$. Then $\xi^{-m}h(w)>0$ and $w\in C_{v,m}$. Using the subseries bound we get $\xi^{-m}h(C_{v,m})>0$. By item 1 of Proposition 4.2 we have $h(A_{v,n})=\sum_{j=0}^{n}h(C_{v,j})h(D_{n-j})$ and we conclude that
$\xi^{-n-m}h(A_{v,\,n+m})\geq\xi^{-m}h(C_{v,\,m})\cdot\xi^{-n}h(D_{n})$
for every $n\in\mathbb{N}_{0}$. Hence
$\sum_{j=0}^{n}\xi^{-j-m}h(A_{v,\,j+m})\geq\xi^{-m}h(C_{v,\,m})\cdot\sum_{j=0}^{n}\xi^{-j}h(D_{j})\to+\infty$}
for $n\to\infty$. So $(A_{v,h})_{1/\xi}(1)_{1}=+\infty$. $\Box$
4.3 Pólya’s theorems for $v\neq 1$ and $\xi$-convex weights
We proceed to $\xi$-convex weights and begin with twisted nonnegative ones. The next theorem generalizes Section 1.3.
Theorem 4.9 ($\exists\,U(1)_{1}$ and $h\geq 0$)
Let $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ be a $v$-transitive $\xi$-convex light weight such that $v\in V(h)$. If $(B_{h})_{1/\xi}(1)_{1}$ exists, then
$\lim_{n\to\infty}\xi^{-n}h(A_{v,\,n})=\sqrt{1-\frac{(B_{v,\,h})_{1/\xi}(1)_{1}}{(B_{h})_{1/\xi}(1)_{1}}}\in(0,\,1)\,.$
Proof. By Proposition 2.31, $D_{h}(x)=\frac{1}{1-\xi x}$. Since $B_{n},B_{v,n},C_{v,n}\subset D_{n}$ for every $n\in\mathbb{N}_{0}$, we see by the subseries bound that $(B_{h})_{1/\xi}(x)$, $(B_{v,h})_{1/\xi}(x)$ and $(C_{v,h})_{1/\xi}(x)$ have all coefficients in $[0,1]$ and are in $H_{1}$. Since $B_{v,n}\subset B_{n}$ for every $n\in\mathbb{N}_{0}$, we see by the subseries bound that
$0\leq(B_{v,\,h})_{1/\xi}(1)_{1}\leq(B_{h})_{1/\xi}(1)_{1}<+\infty\,.$
By item 2 of Proposition 4.2 we have $(C_{v,h})_{1/\xi}(x)^{2}=1-\frac{(B_{v,h})_{1/\xi}(x)}{(B_{h})_{1/\xi}(x)}$. Using the arithmetic of functional limits, Proposition 2.20 (more precisely, the version for nonnegative coefficients), the bound $(B_{h})_{1/\xi}(1)_{1}\geq 1>0$ and Theorem 2.18, we compute that
$\big((C_{v,\,h})_{1/\xi}(1)_{1}\big)^{2}=((C_{v,\,h})_{1/\xi})^{2}(1)_{1}$ $=$ $\lim_{x\nearrow 1}((C_{v,\,h})_{1/\xi})^{2}(x)$
$=$ ${\textstyle 1-\frac{\lim_{x\nearrow 1}(B_{v,\,h})_{1/\xi}(x)}{\lim_{x\nearrow 1}(B_{h})_{1/\xi}(x)}}$
$=$ ${\textstyle 1-\frac{(B_{v,h})_{1/\xi}(1)_{1}}{(B_{h})_{1/\xi}(1)_{1}}\ \ (\in[0,\,1])\,.}$
We used the third Abelian theorem on the first and third lines. Thus, due to nonnegativity, $(C_{v,\,h})_{1/\xi}(1)_{1}$ equals to the displayed root. By item 1 of Proposition 4.2 and by Proposition 2.31 we have $A_{v,h}(x)=C_{v,h}(x)D_{h}(x)=C_{v,h}(x)\frac{1}{1-\xi x}$. Proposition 2.15 gives that
$\lim_{n\to\infty}\xi^{-n}h(A_{v,\,n})=(C_{v,\,h})_{1/\xi}(1)_{1}$
and the stated formula follows. The membership to $(0,1)$ follows from the inequality
$1\leq(B_{v,\,h})_{1/\xi}(1)_{1}<(B_{h})_{1/\xi}(1)_{1}\,.$
The strict inequality is witnessed by any walk in $G(h)$ from $1$ to $v$, which yields (by going back via the reversal) a walk in $G(h)$ from $1$ to $1$ visiting $v$. $\Box$
Theorem 4.9 covers only the finite case $(B_{h})_{1/\xi}(1)_{1}<+\infty$; we move the case $(B_{h})_{1/\xi}(1)_{1}=+\infty$ to part 1 of Theorem 4.13.
Theorem 4.10 ($\exists\,U(1)_{1}$)
1. Let $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive $\xi$-convex light weight such that $v\in V(h)$. If $(B_{v,h})_{1/\xi}(1)_{1}$, $(B_{h})_{1/\xi}(1)_{1}$ and $(C_{v,h})_{1/\xi}(1)_{1}$ exist and $(B_{h})_{1/\xi}(1)_{1}\neq 0$, then
$\lim_{n\to\infty}\xi^{-n}h(A_{v,\,n})\in\mathrm{sqrt}\Big(1-\frac{(B_{v,\,h})_{1/\xi}(1)_{1}}{(B_{h})_{1/\xi}(1)_{1}}\Big)\,.$
2. Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive 0-convex light weight such that $v\in V(h)$. If $B_{v,h}(1)_{1}$, $B_{h}(1)_{1}$ and $C_{v,h}(1)_{1}$ exist and $B_{h}(1)_{1}\neq 0$, then
$A_{v,\,h}(1)_{1}\in\mathrm{sqrt}\Big(1-\frac{B_{v,\,h}(1)_{1}}{B_{h}(1)_{1}}\Big)\,.$
Proof. 1. By item 1 of Proposition 4.2 and by Proposition 2.31 we have $A_{v,h}(x)=C_{v,h}(x)D_{h}(x)=C_{v,h}(x)\frac{1}{1-\xi x}$. By Proposition 2.15,
$\lim_{n\to\infty}\xi^{-n}h(A_{v,\,n})=(C_{v,\,h})_{1/\xi}(1)_{1}\,.$
By item 2 of Proposition 4.2 we have $(C_{v,h})_{1/\xi}(x)^{2}=1-\frac{(B_{v,h})_{1/\xi}(x)}{(B_{h})_{1/\xi}(x)}$. By Proposition 2.22,
$\big((C_{v,\,h})_{1/\xi}(1)_{1}\big)^{2}=1-\frac{(B_{v,\,h})_{1/\xi}(1)_{1}}{(B_{h})_{1/\xi}(1)_{1}}\,.$}
The stated formula follows.
2. By item 1 of Proposition 4.2 and by Proposition 2.31, $A_{v,h}(x)=C_{v,h}(x)$. We use the displayed formula for $(C_{v,h})_{1/\xi}(1)_{1}$ with $\xi=1$. $\Box$
We again trade the non-vanishing of $B_{h}(1)_{1}$ for the existence of $E_{v,h}(1)_{1}$ and obtain the following variant of the last theorem.
Theorem 4.11 ($\exists\,U(1)_{1}$)
1. Let $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive $\xi$-convex light weight such that $v\in V(h)$. If $(B_{v,h})_{1/\xi}(1)_{1}$, $(B_{h})_{1/\xi}(1)_{1}$, $(C_{v,h})_{1/\xi}(1)_{1}$ and $(E_{v,h})_{1/\xi}(1)_{1}$ exist, then
$\lim_{n\to\infty}\xi^{-n}h(A_{v,\,n})\in\mathrm{sqrt}\Big(1-\frac{(B_{v,\,h})_{1/\xi}(1)_{1}}{(B_{h})_{1/\xi}(1)_{1}}\Big)\,$,
where for $(B_{h})_{1/\xi}(1)_{1}=0$ we have $(B_{v,h})_{1/\xi}(1)_{1}=0$ and interpret the expression $\frac{0}{0}$ as $1$, so that then the limit is zero.
2. Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive 0-convex light weight such that $v\in V(h)$. If $B_{v,h}(1)_{1}$, $B_{h}(1)_{1}$, $C_{v,h}(1)_{1}$ and $E_{v,h}(1)_{1}$ exist, then
$A_{v,\,h}(1)_{1}\in\mathrm{sqrt}\Big(1-\frac{B_{v,\,h}(1)_{1}}{B_{h}(1)_{1}}\Big)\,$,
where for $B_{h}(1)_{1}=0$ we have $B_{v,h}(1)_{1}=0$ and interpret the expression $\frac{0}{0}$ as $1$, so that then $A_{v,h}(1)_{1}=0$.
Proof. If $(B_{h})_{1/\xi}(1)_{1}\neq 0$, resp. $B_{h}(1)_{1}\neq 0$, we use the previous proofs. We therefore assume that $(B_{h})_{1/\xi}(1)_{1}=0$, resp. $B_{h}(1)_{1}=0$.
1. From the proof of part 1 in Theorem 4.10 we know that
$\lim_{n\to\infty}\xi^{-n}h(A_{v,\,n})=(C_{v,\,h})_{1/\xi}(1)_{1}\,.$
Using item 2 of Proposition 4.2 and Propositions 2.20 and 2.19, it follows from $(B_{h})_{1/\xi}(1)_{1}=0$ that $(B_{v,h})_{1/\xi}(1)_{1}=0$. Item 3 of Proposition 4.2 and Proposition 2.20 give $(C_{v,h})_{1/\xi}(1)_{1}=0$. So the displayed limit is 0.
2. Since $\xi=0$, item 1 of Proposition 4.2 and Proposition 2.31 give that $A_{v,h}(x)=C_{v,h}(x)$. As we know, $B_{h}(1)_{1}=0$ implies that $B_{v,h}(1)_{1}=0$ and that $C_{v,h}(1)_{1}=0$. So $A_{v,h}(1)_{1}=0$. $\Box$
The last two theorems generalize the case $d\geq 3$ of the formula (1) for $\overline{v}\neq\overline{0}$. The case $d\leq 2$ is generalized in Theorem 4.13.
Theorem 4.12 ($\exists\,U(1)_{2}$)
1. Let $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive $\xi$-convex light weight such that $v\in V(h)$. If
$c=\sum_{n\geq 1}|h(B_{n})|<1$
and if $(B_{v,h})_{1/\xi}(1)_{2}$ exists, then the following holds. If $B_{v,h}(x)=B_{h}(x)$ then $A_{v,h}(x)=0$. If $B_{v,h}(x)\neq B_{h}(x)$, $k\in\mathbb{N}_{0}$ is minimum such that $h(B_{v,k})\neq h(B_{k})$ and if
$\big(|h(B_{v,\,k})-h(B_{k})|\cdot(1-c)\big)^{-1}\sum_{n>k}|h(B_{v,\,n})-h(B_{n})|<1\,$,
then $(B_{h})_{1/\xi}(1)_{2}\neq 0$ and
$\lim_{n\to\infty}\xi^{-n}h(A_{v,\,n})\in\mathrm{sqrt}\Big(1-\frac{(B_{v,\,h})_{1/\xi}(1)_{2}}{(B_{h})_{1/\xi}(1)_{2}}\Big)\,.$
2. Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive 0-convex light weight such that $v\in V(h)$. If
$c=\sum_{n\geq 1}|h(B_{n})|<1$
and if $B_{v,h}(1)_{2}$ exists, then the following holds. If $B_{v,h}(x)=B_{h}(x)$ then $A_{v,h}(x)=0$. If $B_{v,h}(x)\neq B_{h}(x)$, $k\in\mathbb{N}_{0}$ is minimum such that $h(B_{v,k})\neq h(B_{k})$ and if
$\big(|h(B_{v,\,k})-h(B_{k})|\cdot(1-c)\big)^{-1}\sum_{n>k}|h(B_{v,\,n})-h(B_{n})|<1\,$,
then $B_{h}(1)_{2}\neq 0$ and
$A_{v,\,h}(1)_{2}\in\mathrm{sqrt}\Big(1-\frac{B_{v,\,h}(1)_{2}}{B_{h}(1)_{2}}\Big)\,.$
Proof. Let $B_{v,h}(x)=B_{h}(x)$. By item 2 of Proposition 4.2 we have $C_{v,h}(x)^{2}=1-\frac{B_{v,h}(x)}{B_{h}(x)}=1-1=0$ and $C_{v,h}(x)=0$. Item 1 of Proposition 4.2 gives $A_{v,h}(x)=C_{v,h}(x)D_{h}(x)=0$.
1. Let $h$ be as stated, $B_{v,h}(x)\neq B_{h}(x)$, $k$ be as stated, and let the latter displayed bound hold. By item 2 of Proposition 4.2, we have $C_{v,h}(x)^{2}=1-\frac{B_{v,h}(x)}{B_{h}(x)}$. By Proposition 2.30, $(B_{h})_{1/\xi}(1)_{2}\neq 0$ and
$\big((C_{v,h})_{1/\xi}(1)_{2}\big)^{2}=1-\frac{(B_{v,h})_{1/\xi}(1)_{2}}{(B_{h})_{1/\xi}(1)_{2}}\,.$}
By item 1 of Proposition 4.2 and by Proposition 2.31 we have $A_{v,h}(x)=C_{v,h}(x)D_{h}(x)=C_{v,h}(x)\frac{1}{1-\xi x}$. Proposition 2.15 gives the stated formula:
$\lim_{n\to\infty}\xi^{-n}h(A_{v,n})={\textstyle(C_{v,\,h})_{1/\xi}(1)_{1}=(C_{v,\,h})_{1/\xi}(1)_{2}\in\mathrm{sqrt}\big(1-\frac{(B_{v,h})_{1/\xi}(1)_{2}}{(B_{h})_{1/\xi}(1)_{2}}\big)\,.}$
2. Let $h$ be as stated, $B_{v,h}(x)\neq B_{h}(x)$, $k$ be as stated, and let the latter displayed bound hold. By part 1 for $\xi=1$ we have $\big(C_{v,h}(1)_{2}\big)^{2}=1-\frac{B_{v,h}(1)_{2}}{B_{h}(1)_{2}}$. By item 1 of Proposition 4.2 and by Proposition 2.31 we have $A_{v,h}(x)=C_{v,h}(x)D_{h}(x)=C_{v,h}(x)$. The stated formula follows. $\Box$
Theorem 4.13 ($U(1)_{1}=\infty$)
1. Let $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ be a $v$-transitive $\xi$-convex light weight such that $v\in V(h)$. If $(B_{h})_{1/\xi}(1)_{1}=+\infty$, then
$\lim_{n\to\infty}\xi^{-n}h(A_{v,\,n})=1\,.$
2. Let $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}$ be a $v$-transitive 0-convex light weight such that $v\in V(h)$. If $B_{h}(x),C_{v,h}(x)\in H_{1}$, $(B_{v,h})_{1/\xi}(1)_{1}$ exists and if $(B_{h})_{1/\xi}(1)_{1}=\pm\infty$, then
$\lim_{x\nearrow 1}(A_{v,\,h})_{1/\xi}(x)\in\{-1,\,1\}\,.$
Proof. 1. By Proposition 2.31, $D_{h}(x)=\frac{1}{1-\xi x}$. By the subseries bound,
$0\leq\xi^{-n}h(B_{n}),\,\xi^{-n}h(B_{v,\,n}),\,\xi^{-n}h(C_{v,\,n}),\,\xi^{-n}h(E_{v,\,n})\leq\xi^{-n}h(D_{n})=1$
for every $n\in\mathbb{N}_{0}$. Hence $B_{h}(x)$, $B_{v,h}(x)$, $C_{v,h}(x)$, and $E_{v,h}(x)$ are in $H_{1}$. Any shortest walk joining $1$ and $v$ in $G(h)$ shows that for some $m\in\mathbb{N}$ we have $\xi^{-m}h(E_{v,m})>0$. Thus the function $(E_{v,h})_{1/\xi}(x)$ is positive and increasing on $(0,1)$. Since $(C_{v,h})_{1/\xi}(x)$ is bounded on $[0,1)$ by item 2 of Proposition 4.2, $(B_{v,h})_{1/\xi}(x)$ is bounded by item 3. Since by item 1 of Proposition 4.2 and by Proposition 2.31 we have $A_{v,h}(x)=C_{v,h}(x)D_{h}(x)=C_{v,h}(x)\frac{1}{1-\xi x}$, Proposition 2.15 gives
$\lim_{n\to\infty}\xi^{-n}h(A_{v,\,n})=(C_{v,\,h})_{1/\xi}(1)_{1}=(C_{v,\,h})_{1/\xi}(1)_{2}\,.$
By item 2 of Proposition 4.2, by the arithmetic of functional limits, and by Theorem 2.18, we have
$((C_{v,\,h})_{1/\xi})^{2}(1)_{1}=\lim_{x\nearrow 1}(C_{v,\,h})_{1/\xi}(x)^{2}=1-\frac{\lim_{x\nearrow 1}(B_{v,\,h})_{1/\xi}(x)}{\lim_{x\nearrow 1}(B_{h})_{1/\xi}(x)}=1\,$,}
because the function $(B_{v,h})_{1/\xi}(x)$ is bounded on $[0,1)$ and $\lim_{x\nearrow 1}(B_{h})_{1/\xi}(x)=B_{h}(1)_{1}=+\infty$. Hence, by nonnegativity, $(C_{v,h})_{1/\xi}(1)_{1}=(C_{v,h})_{1/\xi}(1)_{2}=1$ because
$\big((C_{v,\,h})_{1/\xi}(1)_{2}\big)^{2}=\big((C_{v,\,h})_{1/\xi}\big)^{2}(1)_{2}$
by (the version for nonnegative coefficients of) Proposition 2.20. The stated formula follows.
2. By item 1 of Proposition 4.2 and by Proposition 2.31, $A_{v,h}(x)=C_{v,h}(x)$. By item 2 of Proposition 4.2 and by the arithmetic of functional limits,
$\lim_{x\nearrow 1}(C_{v,\,h})_{1/\xi}(x)^{2}=1-\frac{\lim_{x\nearrow 1}(B_{v,\,h})_{1/\xi}(x)}{\lim_{x\nearrow 1}(B_{h})_{1/\xi}(x)}=1-\frac{(B_{v,\,h})_{1/\xi}(1)_{1}}{\pm\infty}=1-0=1\,.$}
In the second equality we used Theorems 2.16 (in the numerator) and 2.17 (in the denominator). Since the involved real functions are continuous on $[0,1)$, the limit $\lim_{x\nearrow 1}(A_{v,h})_{1/\xi}(x)$ exists and
$\lim_{x\nearrow 1}(A_{v,\,h})_{1/\xi}(x)=\lim_{x\nearrow 1}(C_{v,\,h})_{1/\xi}(x)\in\{-1,\,1\}\,.$
$\Box$
We feel that part 2 of the theorem is not completely satisfactory. Perhaps the assumption that $(B_{v,h})_{1/\xi}(1)_{1}$ exists may be dropped, and it is true that if $(B_{h})_{1/\xi}(1)_{1}=\pm\infty$, then always
$\lim_{x\nearrow 1}\frac{(B_{v,\,h})_{1/\xi}(x)}{(B_{h})_{1/\xi}(x)}=0\,.$}
5 Overview of general Pólya’s theorems
Thm type assumption conclusion
3.2 $v=1,{}_{1}$ $A(1),B(1),C(1),D(1)$ $A(1)=(1-\frac{1}{B(1)})D(1)$
3.3 $v=1,{}_{2}$ $\sum|h(B_{n})|<1,D(1)$ $A(1)=(1-\frac{1}{B(1)})D(1)$
3.4 $v=1,\infty$ $\xi\mathbb{R}_{\geq 0},D(1)=+\infty$ $A(1)=+\infty$
3.5 $v=1,{}_{1}$, $\xi$-c. $\xi\mathbb{R}_{\geq 0}$
$\lim h(A_{n})=1-\frac{1}{B(1)}\in$
$\in(0,1]$
3.6.1 $v=1,{}_{1}$, $\xi$-c. $B(1),C(1)$ $\lim h(A_{n})=1-\frac{1}{B(1)}$
3.6.2 $v=1,{}_{1}$, 0-c. $B(1),C(1)$ $A(1)=1-\frac{1}{B(1)}$
3.7.1 $v=1,{}_{2}$, $\xi$-c. $\sum|h(B_{n})|<1$ $\lim h(A_{n})=1-\frac{1}{B(1)}$
3.7.2 $v=1,{}_{2}$, 0-c. $\sum|h(B_{n})|<1$ $A(1)=1-\frac{1}{B(1)}$
3.8.1 $v=1,\infty$, $\xi$-c. $\xi\mathbb{R}_{\geq 0},B(1)=+\infty$ $\lim h(A_{n})=1$
3.8.2 $v=1,\infty$, 0-c.
$\xi\mathbb{R},B(x),C(x)\in H_{1}$,
$B(1)=+\infty$
$\lim A(x)=1$
4.5 $v\neq 1,{}_{1}$
$A(1),B(1)\neq 0,B_{v}(1)$,
$C_{v}(1),D(1)$
$A(1)\in(1-\frac{B_{v}(1)}{B(1)})^{\frac{1}{2}}D(1)$
4.6 $v\neq 1,{}_{1}$
$A(1),B(1),B_{v}(1)$,
$C_{v}(1),D(1),E(1)$
$A(1)\in(1-\frac{B_{v}(1)}{B(1)})^{\frac{1}{2}}D(1)$
4.7 $v\neq 1,{}_{2}$
$B_{v}(1),D(1)$,
$\sum|h(B_{n})|,\,\sum<1$
$A(1)\in(1-\frac{B_{v}(1)}{B(1)})^{\frac{1}{2}}D(1)$
4.8 $v\neq 1,\infty$ $\xi\mathbb{R}_{\geq 0},D(1)=+\infty$ $A(1)=+\infty$
4.9 $v\neq 1,{}_{1}$, $\xi$-c. $\xi\mathbb{R}_{\geq 0},B(1)$
$\lim h(A_{n})=$
$(1-\frac{B_{v}(1)}{B(1)})^{\frac{1}{2}}\in(0,1)$
4.10.1 $v\neq 1,{}_{1}$, $\xi$-c. $B(1)\neq 0,B_{v}(1),C_{v}(1)$ $\lim h(A_{n})\in(1-\frac{B_{v}(1)}{B(1)})^{\frac{1}{2}}$
4.10.2 $v\neq 1,{}_{1}$, 0-c. $B(1)\neq 0,B_{v}(1),C_{v}(1)$ $A(1)\in(1-\frac{B_{v}(1)}{B(1)})^{\frac{1}{2}}$
4.11.1 $v\neq 1,{}_{1}$, $\xi$-c.
$B(1),B_{v}(1),C_{v}(1)$,
$E(1)$
$\lim h(A_{n})\in(1-\frac{B_{v}(1)}{B(1)})^{\frac{1}{2}}$
4.11.2 $v\neq 1,{}_{1}$, 0-c.
$B(1),B_{v}(1),C_{v}(1)$,
$E(1)$
$A(1)\in(1-\frac{B_{v}(1)}{B(1)})^{\frac{1}{2}}$
4.12.1 $v\neq 1,{}_{2}$, $\xi$-c.
$B_{v}(1),\sum h(B_{n})<1$,
$\sum<1$
$\lim h(A_{n})\in(1-\frac{B_{v}(1)}{B(1)})^{\frac{1}{2}}$
4.12.2 $v\neq 1,{}_{2}$, 0-c.
$B_{v}(1),\sum h(B_{n})<1$,
$\sum<1$
$A(1)\in(1-\frac{B_{v}(1)}{B(1)})^{\frac{1}{2}}$
4.13.1 $v\neq 1,\infty$, $\xi$-c. $\xi\mathbb{R}_{\geq 0},B(1)=+\infty$ $\lim h(A_{n})=1$
4.13.2 $v\neq 1,\infty$, 0-c.
$\xi\mathbb{R},B(x),C_{v}(x)\in H_{1}$,
$B_{v}(1),B(1)=\pm\infty$
$\lim A(x)\in\{-1,1\}$
The first column lists the theorems. The second column lists the theorem types, with 1 and 2 abbreviating $\exists\,U(1)_{1}$ and $\exists\,U(1)_{2}$, respectively. By $\infty$ we abbreviate that $U(1)_{1}=\infty$. The third and fourth columns list the assumptions and conclusions of the theorems. We always omit indices h, 1/ξ, n→∞ and x↗1, and factor $\xi^{-n}$. We omit indices 1, 2, and v whenever they can be deduced from the context. By writing $A(1)$ in the third column and first row, we mean the assumption that $A_{h}(1)_{1}$ exists, and similarly in other cases. By writing $\xi\mathbb{R}_{\geq 0}$ and $\xi\mathbb{R}$ we mean the assumptions $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ and $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}$, respectively; else $h\colon\mathbb{N}_{2}\to\mathbb{C}$. $\sum|h(B_{n})|$ abbreviates $\sum_{n\geq 1}|h(B_{n})|$, and $\sum<1$ abbreviates the complicated bound assumed in the respective theorem.